### v3

# PatchAD: Patch-based MLP-Mixer for Time Series Anomaly Detection

Abstract
Anomaly detection stands as a crucial aspect of time series analysis, aiming to identify abnormal events in time series samples. The central challenge of this task lies in effectively learning the representations of normal and abnormal patterns in a label-lacking scenario. Previous research mostly relied on reconstruction-based approaches, restricting the representational abilities of the models. In addition, most of the current deep learning-based methods are not lightweight enough, which prompts us to design a more efficient framework for anomaly detection. In this study, we introduce PatchAD, a novel multi-scale patch-based MLP-Mixer architecture that leverages contrastive learning for representational extraction and anomaly detection. Specifically, PatchAD is composed of four distinct MLP Mixers, exclusively utilizing the MLP architecture for high efficiency and lightweight architecture. Additionally, we also innovatively crafted a dual project constraint module to mitigate potential model degradation. Comprehensive experiments demonstrate that PatchAD achieves state-of-the-art results across multiple real-world multivariate time series datasets. Our code is publicly available.  [^1] https://github.com/EmorZz1G/PatchAD
1 Introduction
Time series anomaly detection is a pivotal aspect of data analysis, focused on identifying abnormal patterns within time-series data that significantly deviate from normal behavior Belay et al. (2023); Choi et al. (2021). This process becomes more complex in multivariate temporal anomaly detection, where it involves multiple channel dimensions. Driven by the rapid progression of large-scale sensing technologies and enhanced data storage capabilities, temporal anomaly detection has found widespread application in various sectors. For instance, in the field of the Internet of Things (IoT), it serves to monitor abnormal events in sensor data Yang et al. (2021). In healthcare sector, it plays a vital role in monitoring a person’s physiological data, thereby assisting in the early detection of health problems Cook et al. (2020).
However, extracting anomalies from extensive and complex temporal data presents significant challenges. First, accurately defining the representations of anomalies is difficult due to the varied and evolving nature of real-world anomalies. For instance, anomalies in machinery could manifest as either high or fluctuating temperatures under different scenarios. Second, labels for anomalous events in time-series data are probably absent or hard to acquire. Thirdly, the high dimensionality and temporal dependencies inherent in time-series data are intricate. While the typical approach involves classifying different time points as normal or anomalous, relying solely on a single time point is often not feasible.
Due to the powerful representational capacity of neural networks, numerous deep learning-based techniques Munir et al. (2019); Zhang et al. (2022) have been proposed for time series anomaly detection, demonstrating remarkable performance and attracting significant interest from the research community. While supervised and semi-supervised methods grapple with the scarcity of labeled data, unsupervised techniques Zhou et al. (2019); Su et al. (2019) have gained prominence due to their ability to operate without strict labeling requirements. While Xu et al. (2022) designed six different anomaly injection methods to train the model in a self-supervised approach, Carmona et al. (2021) utilized outlier exposure and mixup to construct a self-supervised task.
Reconstruction-based approaches, as typical and common unsupervised methods, involve learning representations of historical data and identifying samples that cannot be accurately reconstructed during testing as anomalies Park et al. (2018). This approach has experienced rapid development due to its clear concept and its proficiency in managing complex data representations. However, due to their robust reconstruction capabilities, these methods excel at recovering anomalous data, posing a challenge for reconstruction-based approaches. The core issue is the ineffectiveness of these methods in distinctly modeling normal and abnormal data, complicating their ability to accurately identify anomalies.
Recently, contrastive learning has been captivated by its straightforward, yet exceptional performance, proving effective in areas such as computer vision Chen et al. (2020). In time series anomaly detection, several methods have begun incorporating contrastive learning to enhance model representational capabilities. Nonetheless, these approaches often overlook the design of network structures tailored for contrastive learning. They either do not explicitly model inter-channel dependencies Yang et al. (2023) or rely solely on basic MLPs or Transformers that focus on temporal information Zhou et al. (2022), mainly considering point-level features Audibert et al. (2020). Nevertheless, poor training and inference speeds, architectural complexity, and a lack of lightweight features are some of the drawbacks of high-performing techniques. It was recently shown by Zeng et al. (2022) that MLP structures can also result in outstanding performance on time series tasks. Additionally, they neglected the model degradation issue in contrastive learning, meaning models might degrade or collapse during the training process.
In this paper, we propose PatchAD, a novel Patch-based MLP architecture designed for time series Anomaly Detection. PatchAD leverages contrastive learning to tackle the complexities of time series anomaly detection. Inspired by Yang et al. (2023), we hypothesize that normal time series points share similar latent patterns, indicating a strong correlation among normal time points. On the contrary, anomalies struggle to establish connections with other anomaly points or normal points. Hence, normal data exhibits consistent representations across different views, while anomalies struggle to maintain consistent representations. This underlying principle forms the core of our approach to anomaly detection.
Specifically, we introduce a novel Multi-scale Patching and Embedding approach. This method decomposes the input data along the temporal dimension into non-overlapping patches, enabling the model to learn from features that contain richer semantic information, moving beyond simple point-level features. To address various time scales, different submodels are utilized, each focusing on different patch sizes. To capture diverse views of temporal data representation, we propose different Patch Mixer Encoders to learn the mixed relationships among inter-patches, intra-patches, and inter-channel connections. We utilize a shared MixRep Mixer to uniformly represent the feature space. To prevent the degradation, we not only construct diverse encoders but also introduce a Dual Project Constraint mechanism to prevent the model from settling on overly simplistic solutions, thus substantially boosting its representational capabilities.
The main contributions of our proposed PatchAD can be summarized as follows:
  • The novel PatchAD model, innovatively structured on contrastive learning framework, is adept at managing diverse anomaly types. This is achieved by creating Interpatch Encoders and Intra-patch Encoders, which generate both inter-patch and intra-patch perspectives.
  • Employing a strategy named Multi-scale Patching and Embedding, PatchAD integrates multi-scale subsequence modules constructed exclusively with MLPs. This design is intended to facilitate the learning of semantically richer features.
  • Dual Project Constraint mechanism is specifically designed to impose constraints on the model, reducing its susceptibility to simple, trivial solutions and thus increasing its expressive power.
  • PatchAD’s efficacy is underscored by comprehensive experimental results on eight datasets. These results highlight its superior performance compared to the SOTA time series anomaly detection methods.
Figure 1: The workflow of the proposed PatchAD framework.
2 Preliminaries
In this section, we provide an initial exposition on the formal representation of time series anomaly detection, related methods, and contrastive learning. These are the foundational elements of the research problem and serve as the motivation for our work.
2.1 Problem Formulation
In this paper, we consider a multivariate time series of length $T$ in an unsupervised setting:
$\mathcal{X}=(x_{1},x_{2},\cdots,x_{T}),$ (1)
where $x_{t}\in\mathbb{R}^{C}$ represents a C-dimensional channel feature, such as data from C sensors or machines. The objective is to train a model using $\mathcal{X}_{train}$ and provide predictions $\mathcal{Y}_{test}=(y_{1},y_{2},\cdots,y_{T^{\prime}})$ on another sequence $\mathcal{X}_{test}$ of length $T^{\prime}$. Here, $y_{t}\in\{0,1\}$, where 1 signifies an anomaly at that time point and 0 indicates a normal point.
2.2 Time Series Anomaly Detection
There exist various methods proposed for time series anomaly detection, encompassing statistical approaches, classical machine learning methods, and deep learning techniques. Statistical methods involve moving averages and the autoregressive integrated moving average (ARIMA) model. Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) Schölkopf et al. (2001) and Support Vector Data Description (SVDD) Tax and Duin (2004). Deep learning methods include DAGMM, OmniAnomaly Su et al. (2019), USAD Audibert et al. (2020), LSTM-VAE Park et al. (2018), and DeepAnT, which are trained on normal data and detect anomalies by inference using reconstruction errors. Some techniques Kim et al. (2023),Chen et al. (2022) have leveraged self-attention mechanisms and have demonstrated promising detection results. Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like Li et al. (2021a) introduced sequence variational autoencoder model based on smoothness constraints, whereas Xu et al. (2021) proposed prior and sequence associations for temporal modeling, both achieving enhanced performance.
2.3 MLP Mixer for Time Series
The MLP Mixer, initially proposed as a novel architecture for computer vision Tolstikhin et al. (2021), has gained significant attention in recent years. Recently, there has been analysis indicating that MLP Mixers can effectively handle sequential data and find applications in other domains, such as time series prediction. Specifically, works like Nie et al. (2023); Gong et al. (2023) have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results. Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time-series forecasting (LSTF). To the best of our knowledge, there is currently an absence of MLP Mixer application in the time series anomaly detection.
2.4 Contrastive Learning
Contrastive learning has garnered considerable attention across various domains for its capacity to train models by contrasting positive pairs against negative ones. The objective is to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs. Additionally, methods like BYOL Grill et al. (2020) and SimSiam Chen and He (2021) have demonstrated advanced performance without explicitly constructing negative pairs. Recently, studies such as Eldele et al. (2021) and Pranavan et al. (2022) have explored the integration of contrastive learning into time series analysis. Yang et al. (2023), on the other hand, combines self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection.
3 Proposed Method
3.1 Overall Architecture
Figure 1 illustrates the overall structure of PatchAD. PatchAD adopts patch sizes at multiple scales. We exemplify its architecture using a single scale. PatchAD is an $L$-layer network, where each layer comprises distinct modules: MLP Mixers, Project head, and ReWeight, forming its structure. We have devised four distinct types of MLP Mixers: Channel Mixer, Inter Mixer, Intra Mixer, and MixRep Mixer, each dedicated to learning various channels, inter-patch, intra-patch relationships, and a unified representation space. The Project head serves to prevent the model from converging to trivial solutions, while the ReWeight module manages the weights across different layers.
The focal point of our design lies in the four distinct MLP Mixers. Inspired by Yang et al. (2023), we utilize various MLP Mixers to learn different representations from the input. Normal points share a common latent space across different views, while anomalies, being scarce and lacking specific patterns, struggle to share a coherent representation with normal points. Consequently, normal points exhibit minor discrepancies across different views, whereas anomalies showcase larger variations. PatchAD leverages inter-intra discrepancy to model this relationship.
3.2 Multi-scale Patching and Embedding
To enhance the temporal contextual representation in time series, we incorporated Transformer positional encoding. Acknowledging the significance of different sequence lengths in the time series analysis, we introduced Multi-scale Patching in PatchAD. This enables PatchAD to focus on features of varying sequence lengths. Moreover, incorporating multi-scale information helps compensate for information loss during the patching process. The input data with $C$ channels and length $T$ is transformed into patches of size $P$. It can be seen as dividing the time series data into $N$ non-overlapping blocks of patches, which is called $\operatorname{Patching}(\cdot)$. Thus, the original data of dimensions $C\times T$ is transformed into $C\times N\times P$.
$\mathcal{X}=\operatorname{Patching}(\operatorname{PE}(\mathcal{X})),\ \mathcal{X}^{C\times T}\rightarrow\mathcal{X}^{C\times N\times P},$ (2)
where $\operatorname{PE}(\cdot)$ is positional embedding.
PatchAD employs two value embeddings, i.e., $\operatorname{VE}(\cdot)$, facilitating the features from two distinct perspectives within the framework. And $\operatorname{VE}(\cdot)$ is a single linear layer.
$\mathcal{X}_{inter}$ $=\operatorname{VE}(\mathcal{X}),\ \mathcal{X}^{C\times N\times P}\rightarrow\mathcal{X}^{C\times N\times D},$ (3)
$\mathcal{X}_{intra}$ $=\operatorname{VE}(\mathcal{X}),\ \mathcal{X}^{C\times N\times P}\rightarrow\mathcal{X}^{C\times P\times D}.$
3.3 Patch Mixer Encoder
PatchAD comprises $L$ layers of Patch Mixer layers. This encoder incorporates four distinct MLP Mixers to extract features across different dimensions. Figure 1(c) illustrates the MLP block we designed. It consists of two fully connected layers, a normalization layer, a GELU non-linear activation layer, and utilizes a residual connection from input to output. This can be represented as:
$\mathcal{X}=\operatorname{FC}(\operatorname{GELU}(\operatorname{FC}(\operatorname{Norm}(\mathcal{X}))))+\mathcal{X}.$ (4)
The following are the different types of MLP Mixers utilized in PatchAD:
  1. 1.
    The Channel Mixer enables PatchAD to capture correlations among different channels.
  2. 2.
    The Inter Mixer empowers PatchAD to capture inter-patch representations, facilitating the learning of global expressions.
  3. 3.
    The Intra Mixer enables PatchAD to capture intra-patch representations, facilitating the learning of local expressions.
  4. 4.
    The MixRep Mixer allows PatchAD to embed two distinct views into the same representation space.
The diagram depicts PatchAD comprised of an inter-patch encoder and an intra-patch encoder, generating representations denoted as $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$, representing the inter-patch and intra-patch views, respectively. It’s important to note that The Channel Mixer and The MixRep Mixer share weights, while The Inter Mixer and The Intra Mixer have their own weights. This design facilitates PatchAD in learning differences between views and prevents model degradation.
3.4 Dual Project Constraint
To prevent the model from converging to trivial solutions, each layer’s output requires to be constrained by Dual Project Head. Prior research suggests that employing a similar structure effectively prevents models from converging to trivial solutions. For instance, SimCLR Chen et al. (2020) uses a projection head that maps augmented views of an image into a latent space and promotes smoother convergence during training. The projection head comprises two layers of MLP connections, excluding non-linear and normalization layers. Hence, for both the inter-patch and intra-patch encoders, we can derive the projected representations $\mathcal{N}^{\prime}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}^{\prime}\in\mathbb{R}^{C\times P\times D}$. This computation can be represented as:
$\mathcal{N}^{\prime}=\operatorname{FC}(\operatorname{FC(\mathcal{N})}),\ \mathcal{P}^{\prime}=\operatorname{FC}(\operatorname{FC(\mathcal{P})}).$ (5)
To amalgamate outputs from different layers and prevent model degradation, we employ a simple ReWeight module to assign distinct weights to various layers. Considering representations from $L$ layers, $\{\mathcal{N}_{1},\mathcal{N}_{2},\cdots,\mathcal{N}_{L}\}$ and $\{\mathcal{P}_{1},\mathcal{P}_{2},\cdots,\mathcal{P}_{L}\}$, we can derive the final result:
$\mathcal{N}_{l}$ $=\alpha_{l}\cdot\mathcal{N}_{l},\ \alpha_{l}=\operatorname{Softmax}(\alpha),$ (6)
$\mathcal{P}_{l}$ $=\beta_{l}\cdot\mathcal{P}_{l},\ \beta_{l}=\operatorname{Softmax}(\beta).$
The introduction of ReWeight allows for a better balance between outputs from different layers, resulting in improved performance.
3.5 Objective Function
The inter-patch encoder and intra-patch encoder can represent two distinct views, denoted as $\mathcal{N}$ and $\mathcal{P}$. To quantify the dissimilarity between these representations, we utilize a comparative loss function based on Kullback–Leibler divergence (KL divergence), termed as Inter-Intra Discrepancy. Given the scarcity of anomalies in the data and the abundance of normal samples sharing hidden patterns, similar inputs should yield similar representations for both.
The loss function for $\mathcal{N}$ and $\mathcal{P}$ can be defined as follows:
$\mathcal{L}_{\mathcal{N}}\{\mathcal{P},\mathcal{N}\}$ $=\sum\operatorname{KL}(\mathcal{N},\operatorname{StopGrad}(\mathcal{P}))$ (7)
$+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{P}),\mathcal{N}),$
$\mathcal{L}_{\mathcal{P}}\{\mathcal{P},\mathcal{N}\}$ $=\sum\operatorname{KL}(\mathcal{P},\operatorname{StopGrad}(\mathcal{N}))$ (8)
$+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{N}),\mathcal{P}),$
where $\operatorname{KL}(\cdot||\cdot)$ represents the KL divergence distance, and $\operatorname{StopGrad}$ denotes gradient stoppage, employed for asynchronously optimizing the two branches. The combined loss is defined as:
$\mathcal{L}_{cont}=\frac{\mathcal{L}_{\mathcal{N}}-\mathcal{L}_{\mathcal{P}}}{len(\mathcal{N})}.$ (9)
Due to the mismatch in dimensions between $\mathcal{N}$ and $\mathcal{P}$, an initial upsampling of both is required for comparability. For the inter-patch view, where only differences between patches exist, we replicate the final $\mathcal{N}$ within patches. Conversely, for the intra-patch view, exhibiting anomalies between points within patches, we replicate it multiple times to obtain the final $\mathcal{P}$. This also addresses the necessity for introducing multi-scale, as it compensates for the information loss during upsampling.
To prevent model degradation, we augment the original loss function with an additional constraint from the projection head, defined as:
$\mathcal{L}_{proj}=\frac{\mathcal{L}_{\mathcal{N}^{\prime}}-\mathcal{L}_{\mathcal{P}}}{len(\mathcal{N})}+\frac{\mathcal{L}_{\mathcal{N}}-\mathcal{L}_{\mathcal{P}^{\prime}}}{len(\mathcal{N})}.$ (10)
The final PatchAD loss consists of the aforementioned two components, with a constraint coefficient used to control the strength of the constraint:
$\mathcal{L}=(1-c)\cdot\mathcal{L}_{cont}+c\cdot\mathcal{L}_{proj}.$ (11)
3.6 Anomaly Score
Based on the assumption that normal points across different views share hidden patterns, while anomalies exhibit larger discrepancies, the difference between the two views can be employed as an anomaly score. The final anomaly score can be expressed as:
$\operatorname{AnomalyScore}(\mathcal{X})=\sum\operatorname{KL}(\mathcal{N},\mathcal{P})+\operatorname{KL}(\mathcal{P},\mathcal{N}).$ (12)
Based on the aforementioned anomaly score, we set a hyperparameter $\sigma$ to determine whether a point is considered an anomaly. If it surpasses the threshold, it is identified as an outlier.
${y}_{i}=\left\{\begin{aligned} &1\text{, }\operatorname{AnomalyScore}(\mathcal{X}_{i})\geq\sigma\\ &0\text{, }\operatorname{AnomalyScore}(\mathcal{X}_{i})<\sigma.\end{aligned}\right.$ (13)
Table 1: Comprehensive comparative results on four datasets. The P, R and F1 are the precision, recall and F1-score. All results are in %, the best ones are in Bold, and the second ones are underlined.
Dataset MSL SMAP PSM SWaT Average F1
Metric P R F1 P R F1 P R F1 P R F1
OCSVM Schölkopf et al. (2001) 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
IForest Liu et al. (2008) 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
LOF Breunig et al. (2000) 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
U-Time Perslev et al. (2019) 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
ITAD Shin et al. (2020) 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
DAGMM Zong et al. (2018) 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
VAR Anderson (1976) 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
MMPCACD Yairi et al. (2017) 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
CL-MPPCA Tariq et al. (2019) 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
TS-CP2 Deldari et al. (2021) 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
BeatGAN Zhou et al. (2019) 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
LSTM-VAE Park et al. (2018) 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
BOCPD Adams and MacKay (2007) 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
Deep-SVDD Ruff et al. (2018) 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
LSTM Hundman et al. (2018) 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
OmniAnomaly Su et al. (2019) 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
InterFusion Li et al. (2021b) 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
THOC Shen et al. (2020) 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
AnomalyTrans Xu et al. (2021) 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
DCdetector Yang et al. (2023) 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
PatchAD (Ours) 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 93.28 100.0 96.52 96.60
Table 2: Overall results on NIPS-TS datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
Dataset NIPS-TS-SWAN NIPS-TS-GECCO Average F1
Metric P R F1 P R F1
MatrixProfile Yeh et al. (2016) 16.70 17.50 17.10 4.60 18.50 7.40 12.25
GBRT Elsayed et al. (2021) 44.70 37.50 40.80 17.50 14.00 15.60 28.20
LSTM-RNN Bontemps et al. (2016) 52.70 22.10 31.20 34.30 27.50 30.50 30.85
Autoregression Rousseeuw and Leroy (2005) 42.10 35.40 38.50 39.20 31.40 34.90 36.70
OCSVM Schölkopf et al. (2001) 47.40 49.80 48.50 18.50 74.30 29.60 39.05
IForest* Liu et al. (2008) 40.60 42.50 41.60 39.20 31.50 39.00 40.30
AutoEncoder Sakurada and Yairi (2014) 49.70 52.20 50.90 42.40 34.00 37.70 44.30
AnomalyTrans Xu et al. (2021) 90.70 47.40 62.30 25.70 28.50 27.00 44.65
IForest Liu et al. (2008) 56.90 59.80 58.30 43.90 35.30 39.10 48.70
AnomalyTrans Xu et al. (2021) 90.71 47.43 62.29 29.96 48.63 37.08 49.69
DCdetector Yang et al. (2023) 96.57 59.08 73.31 38.41 59.73 46.76 60.04
PatchAD (Ours) 96.75 59.15 73.41 40.62 62.88 49.35 61.38
4 Experiments
4.1 Setups and Evaluation Metrics
We applied eight datasets for evaluation: (1) MSL (2) SMAP  [^2] MSL & SMAP: https://github.com/ML4ITS/mtad-gat-pytorch(3) PSM (4) SMD  [^3] https://github.com/NetManAIOps/OmniAnomaly (5) SWaT  [^4] https://itrust.sutd.edu.sg/itrust-labs_datasets (6) WADI  [^5] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi (7) NIPS-TS-SWAN (8) NIPSTS-GECCO  [^6] Others: github.com/DAMO-DI-ML/KDD2023-DCdetector. Details about these datasets are outlined in Table 4 in Appendix. Also, implementation and pseudo-code Algorithm 1 details in Appendix.
We compared our model with 27 benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) Classic methods; (6) Change point detection and time-series segmentation methods; (7) Transformer-based methods; See details in Table 1 and Table 2, or in Appendix Baslines.
Additionally, we employed various evaluation metrics for assessing, encompassing commonly used metrics in time-series anomaly detection such as accuracy, precision, recall, and F1-score Xu et al. (2021). Moreover, we adopted state-of-the-art evaluation techniques like affiliation precision and affiliation recall Huet et al. (2022), along with Volume under the surface (VUS) Paparrizos et al. (2022). Although the F1-score is a widely used metric, it does not effectively evaluate anomalous events. Therefore, we introduce the following two new metrics. Affiliation precision and recall are evaluation metrics based on the proximity of predicted events to the true labels. VUS considers the inclusion of anomalous events in its evaluation by computing the receiver operator characteristic (ROC) curve. Employing various methodologies allows for a more comprehensive evaluation of our model. To fairly compare with other methods, we utilized the point adjustment technique, identifying an anomalous segment if the model detects any anomaly point within the segment.
(a) Project constraint
(b) Anomaly ratio
(c) $d_{model}$ size
(d) Encoder layer
(e) Window size
(f) Multi-scale size
Figure 2: Parameter sensitivity studies of main hyper-parameters in PatchAD.
4.2 Comparison Results
Data Source Clarification: Due to substantial device variations, to ensure fair comparison with anomaly transformer (AnomalyTrans) and DCdetector, all results mentioned in the paper concerning them are obtained through our reimplementation of their codes. The results of AnomalyTrans and DCdetector in Table 1 and Table 2 are our reproductions.
From Table 1, PatchAD achieves the best average-level F1 performance. As far as we know, there is ongoing controversy regarding the evaluation of time series anomaly detection. However, precision, recall, and F1-score remain the most widely used evaluation metrics. To comprehensively assess the performance of our approach, additional metrics such as affiliation precision/recall and VUS have been introduced.
In Table 2, the NIPS-TS-SWAN and NIPS-TS-GECCO datasets represent more challenging datasets, encompassing a wider array of anomaly types. PatchAD consistently achieves SOTA results on these two datasets compared to other baseline methods. Particularly noteworthy is the 9.3% increase in F1-score over the DCdetector and a 24.9% increase over the Anomaly Transformer on the NIPS-TS-GECCO dataset.
Given the remarkable performance of AnomalyTrans and DCdetector surpassing other baselines significantly, we conducted a comprehensive comparison of PatchAD with them across all datasets using multiple metrics, as depicted in Table 5 in Appendix. Our proposed method outperforms or at least remains competitive compared to these two SOTA methods. PatchAD attains superior results across multiple datasets and multiple metrics. Despite PatchAD’s relatively simple network structure, its performance remains competitive.
(a) Global point
(b) Contextual point
(c) Seasonal
(d) Group point
(e) Trend
Figure 3: Visualization of generated synthetic data, PatchAD anomaly scores and ground truth.
(a) PatchAD’s loss analysis
(b) Time cost comparison
(c) Time cost in different $d_{model}$
(d) Time cost in different window
Figure 4: The analysis of PatchAD and DCdetector.
4.3 Model Analysis
4.3.1 Ablation Studies
To study the impact of different components in our model, we conducted ablation experiments across three datasets, as presented in Table 3. The ’wo_ch_sharing’ method employs two independent Channel Mixer modules to model channel features, while ’wo_ch_mixer’ denotes the removal of the Channel Mixer module. These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modeling inter-channel information and sharing channel information are beneficial for model learning. Besides, ’wo_pos_emb’ represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design. Furthermore, ’wo_constraint’ showcases the impact of removing the constraint on the projection head. In conjunction with Figure 3(a), we observed that a certain level of constraint facilitates easier model optimization, leading to improved results. This constraint prevents the model from converging to trivial solutions. In subsequent experiments, the default setting for the projection head constraint is 0.2.
Table 3: Ablation experiments on MSL, SMAP, WADI datasets.
Dataset MSL SMAP WADI
Methods Acc P R F1 Acc P R F1 Acc P R F1
wo_ch_sharing 98.78 91.93 96.90 94.35 98.84 93.50 97.75 95.58 98.71 84.99 94.25 89.38
wo_ch_mixer 98.78 92.03 96.82 94.36 98.77 93.50 97.10 95.27 98.92 85.40 97.93 91.24
wo_pos_emb 97.85 91.51 87.73 89.58 96.47 92.18 79.13 85.16 98.23 83.96 85.53 84.74
wo_constraint 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 85.26 94.25 89.53
Ours 98.92 92.05 98.20 95.02 99.10 94.45 98.75 96.55 99.06 85.92 100.00 92.43
4.3.2 Parameter Sensitivity
We conducted a sensitivity analysis on PatchAD’s parameters. The default parameter settings remain consistent with the basic configuration. Figure 1(a) demonstrates the impact of the projection head constraint on PatchAD. Overall, it holds more significance for smaller-scale datasets like MSL. However, excessively strong constraints hinder the model’s learning process. Therefore, setting this constraint between 0.1 to 0.4 is more advisable. Figure 1(b) analyzes PatchAD’s final model performance when the anomaly threshold $\sigma$ ranges from 0.5 to 1.2. The selection of the anomaly threshold appears more robust on the PSM and SMAP datasets compared to the MSL dataset. For most datasets, setting this value between 0.8 and 1 yields optimal results. Figure 1(c) illustrates how the model’s performance is influenced by the network dimension $d_{model}$. PatchAD demonstrates better results with smaller $d_{model}$ sizes. Therefore, to strike a balance between model performance and complexity, we set $d_{model}=40$ to ensure PatchAD operates effectively. Figure 1(d) indicates that PatchAD’s performance is influenced by the number of encoder layers, and given a compromise, we set it to 3 or 4. Figure 1(e) demonstrates PatchAD’s robustness concerning the window size, which ranges from 60 to 210. The time window is a critical parameter, typically set to an initial value of 105 across most datasets. Lastly, Figure 1(f) displays PatchAD’s performance under different multi-scale patch size combinations. The outcomes suggest that combining various patch sizes can enhance PatchAD’s performance.
4.3.3 Visual Analysis
We visualized PatchAD’s response to various anomalies in Figure 3. Synthetic univariate time series were generated by Lai et al. (2021), incorporating anomalies such as global point anomalies, contextual point anomalies, and pattern anomalies such as seasonal, group, and trend-based anomalies. The first row depicts the original data, while the second row displays PatchAD’s output of anomaly scores against the anomaly threshold. In particular, PatchAD has the ability to detect various types of anomalies.
4.3.4 Time Cost
We conducted a time consumption analysis on the SMD dataset under identical experimental conditions. Figure 3(b) demonstrates that under different sliding step conditions, our method outperforms DCdetector. Specifically, when the step is set to 1, PatchAD’s speed is 1.45x faster than DCdetector. Additionally, Figure 3(d) illustrates the time required for model training with a fixed window size (window size=105). This indicates that PatchAD exhibits higher efficiency in training compared to DCdetector. Figure 3(c) shows the time required for model training with a fixed sliding step (step=35) while varying the window size from 35 to 175 for different $d_{model}$ sizes. Both graphs show that the time consumption of PatchAD exhibits a linear growth with respect to the independent variables. Thus, increasing the window size or enlarging $d_{model}$ has a relatively minor impact on PatchAD. However, the time step noticeably affects the duration as this parameter leads to increased data volume.
5 Conclusion
We presents PatchAD, an innovative algorithm tailored for time series anomaly detection. PatchAD developed a multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning. This architecture is unique in its use of a purely MLP-based approach to discern both inter-patch and intra-patch relationships in time series sequences. By amplifying the differences between two data views, PatchAD enhance its capability to differentiate between normal and anomalous patterns. Furthermore, PatchAD adeptly learns the relationships between different data channels and incorporates a Dual Project Constraint to avert model degradation. Moreover, the efficacy of PatchAD, including its various components, is validated through extensive experimentation. The results demonstrate that proposed PatchAD outperforms other methods across eight benchmark datasets.
Appendix
This is the appendix of PatchAD: Patch-based MLP-Mixer for Time Series Anomaly Detection.
Datasets
Many publicly available real-world datasets are extensively used in existing methods. We utilize eight datasets to evaluate model performance: (1) MSL (Mars Science Laboratory dataset), (2) SMAP (Soil Moisture Active Passive dataset), (3) PSM (Pooled Server Metrics), (4) SMD (Server Machine Dataset), (5) SWaT (Secure Water Treatment Testbed), (6) WADI (Water Distribution Testbed), (7) NIPS-TS-SWAN, and (8) NIPSTS-GECCO. Table 4 shows the details.
Table 4: Details of benchmark datasets.
Dataset #Training #Test (Labeled) Dimension Anomaly ratio (%)
SML 58317 73729 55 10.5
SMAP 135183 427617 25 12.8
PSM 132481 87841 25 27.8
SMD 708377 708393 38 4.2
SWaT 99000 89984 26 12.2
WADI 1048571 172801 123 5.99
NIPS-TS-SWAN 60000 60000 38 32.6
NIPS-TS-GECCO 69260 69261 9 1.1
Baslines
We compared our model against 27 benchmark models, including: (1) Reconstruction-based: AutoEncoder Sakurada and Yairi [2014], LSTM-VAE Park et al. [2018], OmniAnomaly Su et al. [2019], BeatGAN Zhou et al. [2019], InterFusion Li et al. [2021b]; (2) Autoregression-based: VAR Anderson [1976], Autoregression Rousseeuw and Leroy [2005], LSTM-RNN Bontemps et al. [2016], LSTM Hundman et al. [2018], CL-MPPCA Tariq et al. [2019]; (3) Density-based: LOF Breunig et al. [2000], MMPCACD Yairi et al. [2017], DAGMM Zong et al. [2018]; (4) Clustering-based: DeepSVDD Ruff et al. [2018], THOC Shen et al. [2020], ITAD Shin et al. [2020]; Matrix Profile Yeh et al. [2016]; (5) The classic methods: OCSVM Schölkopf et al. [2001], OCSVM-based subsequence clustering (OCSVM*), IForest Liu et al. [2008], IForest-based subsequence clustering (IForest*), Gradient boosting regression (GBRT) Elsayed et al. [2021]; (6) Change point detection and time series segmentation methods: BOCPD Adams and MacKay [2007], U-Time Perslev et al. [2019], TS-CP2 Deldari et al. [2021]; (7) Transformer-based: Anomaly Transformer  [^7] https://github.com/thuml/Anomaly-Transformer Xu et al. [2021]; DCdetector  [^8] https://github.com/DAMO-DI-ML/KDD2023-DCdetector Yang et al. [2023]; These two are current state-of-the-art methods.
Implementation Details
All experiments were conducted using PyTorch implementation on a single NVIDIA GTX 1080TI GPU. We summarize the default hyperparameters of PatchAD as follows: PatchAD consists of 3 encoder layers ($L$=3). The hidden layer dimension $d_{{model}}$ is set to 40. The default projector constraint coefficient is $0.2$. The Adam optimizer’s default learning rate is $10^{-4}$, with a batch size of 128, trained for 3 epochs across all datasets. Different patch sizes and window sizes were chosen for each dataset, defaulting to [3, 5] and 105, respectively. If an anomaly score at any given timestamp exceeds a threshold $\sigma$, the model identifies it as an anomaly. The default value for $\sigma$ is 1. For details on parameter selection, please refer to the Appendix.
Algorithm’s Pseudo-code
Algorithm 1 PatchAD’s Python-like code
PE = PositionEmbedding(d_model) # Postional Embedding
VE_n, VE_p = Linear(dim1, d_model), Linear(dim2, d_model)
# Value Embedding of N and P views
# Patching and Embedding
x = PE(x)
x_n = rearange(x, ’b (n p) c -> b c n p’, p=patch_size)
x_p = rearange(x, ’b (p n) c -> b c p n’, p=patch_size)
x_n = VE_n(x_n)
x_p = VE_p(x_p)
# x_n’s shape (B, c_dim, n_dim, d_model)
# x_p’s shape (B, c_dim, p_dim, d_model)
Channel_Mixer = MLP(c_dim, transpose_dim=1)
Inter_Mixer = MLP(n_dim, transpose_dim=2)
Intra_Mixer = MLP(p_dim, transpose_dim=2)
MixRep_Mixer = MLP(d_model, transpose_dim=3)
# Patch Mixer Encoding
x_inter = Channel_Mixer(x_n)
x_intra = Channel_Mixer(x_p)
x_inter = Inter_Mixer(x_inter)
x_intra = Intra_Mixer(x_intra)
x_inter = MixRep_Mixer(x_inter)
x_intra = MixRep_Mixer(x_intra)
# Dual Project
x_inter = x_inter.mean(1)
x_intra = x_intra.mean(1)
Inter_Proj = MLP(d_model)
Intra_Proj = MLP(d_model)
x_inter2 = Inter_Proj(x_inter)
x_intra2 = Intra_Proj(x_intra)
# Ouput shape (B, n_dim, d_model) and (B, p_dim, d_model)
# Upsample to (B, L, d_model)
# Update
def loss_fn(a, b):
    return KL(a, b.detach()) + KL(b.detach(), a)
def inter_intra_loss_fn(n, p):
    return loss_fn(n, p) - loss_fn(p, n)
loss1 = inter_intra_loss_fn(x_inter, x_intra)
loss_n_proj = inter_intra_loss_fn(x_inter, x_intra2)
loss_p_proj = inter_intra_loss_fn(x_inter2, x_intra)
loss_proj = loss_n_proj + loss_p_proj
final_loss = (1-c) * loss1 + c * loss_proj
# backward and optimize
Full Comparison Results
In Table 5, Aff-P and Aff-R are the precision and recall of affiliation metric Huet et al. [2022], respectively. R_A_R and R_A_P are Range-AUC-ROC and Range-AUC-PR Paparrizos et al. [2022], respectively. V_ROC and V_RR are volumes under the surfaces created based on ROC curve and PR curve Paparrizos et al. [2022], respectively. Besides, we give more results in table 6, and the results from more real-world multivariate time-series datasets show that our proposed PatchAD has good robustness as well as performance.
Table 5: Multi-metrics results on all real-world multivariate datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
Dataset Method Acc F1 Aff-P Aff-R R_A_R R_A_P V_ROC V_PR
MSL AnomalyTrans 98.69 93.93 51.76 95.98 90.04 87.87 88.20 86.26
DCdetector 98.83 94.77 51.01 96.74 91.29 89.12 89.80 87.81
PatchAD 98.92 95.02 52.06 96.44 90.89 88.66 90.09 87.99
SMAP AnomalyTrans 99.05 96.41 51.39 98.68 96.32 94.07 95.52 93.37
DCdetector 99.06 96.38 50.29 98.04 93.76 92.40 92.70 91.47
PatchAD 99.15 96.75 52.30 98.70 96.05 94.23 95.12 93.42
PSM AnomalyTrans 98.68 97.37 55.35 80.28 91.83 93.03 88.71 90.71
DCdetector 98.81 97.63 52.33 76.95 89.91 91.74 85.57 88.52
PatchAD 98.95 98.11 55.58 82.72 94.85 95.65 89.26 91.67
SMD AnomalyTrans 99.11 90.06 72.63 98.27 82.12 75.90 81.86 75.69
DCdetector 98.78 86.02 51.61 93.44 76.13 69.85 73.60 67.45
PatchAD 98.84 86.62 50.81 93.32 77.55 71.16 74.09 67.88
SWaT AnomalyTrans 98.62 94.62 53.03 98.08 97.89 93.47 97.92 93.49
DCdetector 93.25 96.51 52.76 97.33 98.23 95.54 98.25 95.56
PatchAD 99.12 96.52 52.78 98.17 98.23 95.55 98.26 95.57
WADI AnomalyTrans 98.17 85.47 87.84 99.75 87.20 78.50 87.36 78.63
DCdetector 98.70 89.26 59.63 90.98 87.56 81.62 87.16 81.23
PatchAD 99.06 92.43 60.37 91.76 89.85 84.18 90.68 84.99
NIPS-TS-SWAN AnomalyTrans 84.57 62.29 58.45 9.49 86.42 93.26 84.81 92.00
DCdetector 85.99 73.31 38.73 3.13 88.02 94.78 86.02 93.49
PatchAD 86.05 73.41 51.44 4.34 88.03 94.80 86.13 93.57
NIPS-TS-GECCO AnomalyTrans 98.26 37.08 55.65 89.12 60.74 28.17 60.48 28.02
DCdetector 98.57 46.76 49.66 87.68 61.01 32.13 60.77 31.90
PatchAD 98.64 49.35 53.04 89.48 63.59 35.92 63.09 35.43
Table 6: More results on all real-world multivariate datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
Dataset MSL SMAP PSM SMD SWaT Average F1
Metric P R F1 P R F1 P R F1 P R F1 P R F1
LOF 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 56.34 39.86 46.68 72.15 65.43 68.62 60.94
OCSVM 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 44.34 76.72 56.19 45.39 49.22 47.23 60.25
U-Time 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 65.95 74.75 70.07 46.20 87.94 60.58 65.62
IForest 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 42.31 73.29 53.64 49.29 44.95 47.02 61.22
DAGMM 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 67.30 49.89 57.30 89.92 57.84 70.40 70.18
ITAD 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 86.22 73.71 79.48 63.13 52.08 57.08 70.92
VAR 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 78.35 70.26 74.08 81.59 60.29 69.34 74.66
MMPCACD 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 71.20 79.28 75.02 82.52 68.29 74.73 75.74
CL-MPPCA 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 82.36 76.07 79.09 76.78 81.50 79.07 76.66
TS-CP2 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 87.42 66.25 75.38 81.23 74.10 77.50 79.00
Deep-SVDD 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 78.54 79.67 79.10 80.42 84.45 82.39 80.97
BOCPD 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 70.90 82.04 76.07 89.46 70.75 79.01 80.33
LSTM-VAE 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 75.76 90.08 82.30 76.00 89.50 82.20 81.24
BeatGAN 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 72.90 84.09 78.10 64.01 87.46 73.92 80.24
LSTM 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 78.55 85.28 81.78 86.15 83.27 84.69 83.32
OmniAnomaly 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 83.68 86.82 85.22 81.42 84.30 82.83 84.69
InterFusion 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 87.02 85.43 86.22 80.59 85.58 83.01 85.70
THOC 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 79.76 90.95 84.99 83.94 86.36 85.13 88.01
AnomalyTrans 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 84.14 96.88 90.06 89.79 100.0 94.62 94.48
DCdetector 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 83.04 89.22 86.02 93.25 100.0 96.51 94.26
PatchAD 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 83.06 90.51 86.62 93.28 100.0 96.52 94.60
Table 7: Parameter studies on project head constraint results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
Contraint Acc P R F1 Acc P R F1 Acc P R F1
0.0 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 97.42 98.03 97.72
0.1 98.92 91.96 98.33 95.03 99.04 94.43 98.33 96.34 98.72 97.35 98.04 97.70
0.2 98.92 92.05 98.20 95.02 99.11 94.42 98.91 96.61 98.91 97.43 98.66 98.04
0.3 98.89 92.03 97.93 94.88 99.14 94.41 99.11 96.71 98.89 97.42 98.59 98.00
0.4 98.86 92.09 97.54 94.73 99.12 94.45 98.97 96.66 98.71 97.42 97.94 97.68
0.5 98.96 92.12 98.60 95.25 99.02 94.40 98.17 96.25 98.73 97.42 98.03 97.72
0.6 98.92 92.11 98.20 95.06 99.01 94.39 98.07 96.19 98.92 97.45 98.70 98.07
0.7 98.84 91.91 97.59 94.67 99.07 94.43 98.55 96.44 98.76 97.39 98.14 97.76
0.8 98.91 92.01 98.20 95.00 99.10 94.40 98.81 96.55 98.92 97.42 98.73 98.07
0.9 98.78 91.99 96.90 94.38 98.96 94.37 97.70 96.01 98.92 97.47 98.64 98.05
Table 8: Parameter studies on anomaly threshold $\sigma$ results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
AR Acc P R F1 Acc P R F1 Acc P R F1
0.5 97.93 95.39 84.39 89.55 99.15 96.46 96.91 96.68 98.45 98.61 95.77 97.17
0.6 98.27 94.70 88.54 91.52 99.15 95.84 97.58 96.70 98.77 98.33 97.20 97.76
0.7 98.17 93.76 88.54 91.07 99.09 95.22 97.77 96.48 98.70 98.07 97.20 97.64
0.8 98.45 93.19 92.04 92.61 99.15 94.73 98.86 96.75 98.95 97.94 98.29 98.11
0.9 98.45 92.36 92.93 92.64 99.01 94.07 98.49 96.23 98.92 97.68 98.45 98.06
1.0 98.84 91.80 97.73 94.67 99.01 94.05 98.49 96.22 98.70 97.30 98.03 97.66
1.1 98.73 91.22 97.36 94.19 98.85 92.76 98.69 95.63 98.68 97.06 98.20 97.63
1.2 98.64 90.45 97.36 93.78 98.76 92.19 98.69 95.33 98.62 96.85 98.20 97.52
Table 9: Parameter studies on $d_{model}$ size results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
$d_{model}$ Acc P R F1 Acc P R F1 Acc P R F1
10 98.62 91.81 95.42 93.58 99.04 94.42 98.28 96.31 98.87 97.38 98.56 97.97
20 98.92 92.08 98.20 95.04 98.98 94.33 97.88 96.07 98.73 97.37 98.05 97.71
30 98.89 92.06 97.94 94.91 99.10 94.43 98.80 96.57 98.87 97.43 98.52 97.97
40 98.91 92.03 98.20 95.02 99.13 94.49 98.95 96.67 98.73 97.42 98.03 97.72
50 98.93 92.15 98.20 95.08 99.11 94.45 98.89 96.62 98.90 97.44 98.63 98.03
60 98.91 92.03 98.20 95.02 99.13 94.44 98.98 96.66 98.94 97.48 98.74 98.10
70 98.87 92.08 97.67 94.79 99.05 94.40 98.43 96.37 98.97 97.48 98.83 98.15
80 98.77 91.99 96.70 94.29 99.03 94.38 98.24 96.27 98.89 97.44 98.58 98.00
90 98.92 92.06 98.20 95.03 99.05 94.38 98.42 96.35 98.89 97.46 98.56 98.01
Table 10: Parameter studies on encoder layers $L$ results (patch size=[3,5], window size=105). All results are in %.
Dataset MSL SMAP PSM
Encoder layer Acc P R F1 Acc P R F1 Acc P R F1
1 98.60 91.72 95.27 93.46 98.76 94.29 96.11 95.19 98.10 97.28 95.82 96.54
2 98.88 91.96 97.93 94.85 99.01 94.42 98.09 96.22 98.79 97.44 98.20 97.82
3 98.92 92.07 98.20 95.03 99.10 94.41 98.79 96.55 98.84 97.47 98.36 97.91
4 98.88 92.19 97.67 94.85 99.11 94.42 98.91 96.61 98.95 97.41 98.83 98.12
5 98.83 91.95 97.41 94.60 99.13 94.46 99.02 96.69 98.87 97.44 98.52 97.98
Table 11: Parameter studies on window size results (window size=105). All results are in %.
Dataset MSL SMAP PSM
Window size Acc P R F1 Acc P R F1 Acc P R F1
30 98.11 91.47 90.52 90.99 98.91 94.38 97.23 95.78 98.88 97.48 98.51 97.99
45 97.98 91.23 89.44 90.33 98.93 95.55 96.14 95.84 98.67 97.30 97.92 97.61
60 98.33 92.05 92.16 92.10 98.86 94.18 97.09 95.62 98.75 97.59 97.93 97.76
75 98.51 91.79 94.31 93.03 98.93 94.27 97.60 95.90 98.72 97.35 98.06 97.70
90 98.71 91.92 96.24 94.03 99.06 94.34 98.55 96.40 98.92 97.44 98.71 98.07
105 98.92 92.05 98.20 95.02 99.08 94.44 98.60 96.48 98.66 97.38 97.79 97.59
120 98.86 92.10 97.55 94.75 99.13 94.47 98.98 96.67 98.79 97.53 98.13 97.83
135 98.65 91.85 95.67 93.72 99.07 94.46 98.50 96.44 98.88 97.38 98.60 97.99
150 98.71 91.83 96.30 94.02 99.12 94.48 98.91 96.64 98.66 97.43 97.72 97.58
165 98.61 91.85 95.27 93.53 99.09 94.50 98.65 96.53 98.73 97.44 97.98 97.71
180 98.53 91.73 94.55 93.12 99.01 94.35 98.18 96.23 98.70 97.55 97.78 97.66
195 98.77 92.02 96.70 94.31 98.96 94.35 97.74 96.01 98.87 97.44 98.50 97.97
210 98.81 91.97 97.17 94.50 98.83 94.30 96.71 95.49 98.92 97.45 98.69 98.06
Table 12: Parameter studies on multi-scale patch size results (window size=105). All results are in %.
Dataset MSL SMAP PSM
Patch size Acc P R F1 Acc P R F1 Acc P R F1
[1] 98.49 91.73 94.12 92.91 99.09 94.47 98.69 96.53 98.66 97.21 97.99 97.60
[3] 98.57 91.67 95.08 93.34 99.08 94.44 98.61 96.48 98.84 97.56 98.26 97.91
[5] 98.34 91.70 92.66 92.18 99.02 94.40 98.20 96.26 98.83 97.44 98.37 97.91
[7] 98.45 91.96 93.41 92.68 99.13 94.46 99.04 96.70 98.88 97.43 98.56 97.99
[1, 3] 98.63 91.87 95.42 93.61 99.11 94.48 98.86 96.62 98.78 97.46 98.16 97.81
[1, 5] 98.63 91.88 95.42 93.61 99.07 94.43 98.58 96.46 98.72 97.37 98.04 97.70
[1, 7] 98.63 91.89 95.42 93.62 99.04 94.43 98.32 96.33 98.71 97.39 97.99 97.69
[3, 5] 98.78 91.99 96.90 94.38 99.07 94.39 98.63 96.46 98.91 97.43 98.66 98.04
[3, 7] 98.68 91.92 95.93 93.88 99.09 94.40 98.76 96.53 98.87 97.41 98.55 97.97
[5, 7] 98.61 91.86 95.27 93.53 99.08 94.44 98.63 96.49 98.85 97.40 98.46 97.93
[1, 3, 5] 98.85 92.01 97.54 94.69 99.12 94.47 98.92 96.64 98.82 97.37 98.39 97.88
[1, 3, 7] 98.74 92.00 96.46 94.18 99.09 94.40 98.71 96.50 98.79 97.43 98.21 97.82
[1, 5, 7] 98.92 92.05 98.20 95.02 99.04 94.38 98.36 96.33 98.67 97.36 97.87 97.61
[3, 5, 7] 98.56 91.76 94.89 93.30 99.08 94.42 98.61 96.46 98.87 97.45 98.52 97.98
[1, 3, 5, 7] 98.62 91.78 95.42 93.56 99.08 94.43 98.61 96.48 98.92 97.43 98.71 98.06

### v5

@@
-PatchAD: Patch-based MLP-Mixer for Time Series Anomaly Detection
+PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection
 Abstract
-Anomaly detection stands as a crucial aspect of time series analysis, aiming to identify abnormal events in time series samples.
-The central challenge of this task lies in effectively learning the representations of normal and abnormal patterns in a label-lacking scenario.
-Previous research mostly relied on reconstruction-based approaches, restricting the representational abilities of the models.
-In addition, most of the current deep learning-based methods are not lightweight enough, which prompts us to design a more efficient framework for anomaly detection.
-In this study, we introduce PatchAD, a novel multi-scale patch-based MLP-Mixer architecture that leverages contrastive learning for representational extraction and anomaly detection.
-Specifically, PatchAD is composed of four distinct MLP Mixers, exclusively utilizing the MLP architecture for high efficiency and lightweight architecture.
-Additionally, we also innovatively crafted a dual project constraint module to mitigate potential model degradation.
-Comprehensive experiments demonstrate that PatchAD achieves state-of-the-art results across multiple real-world multivariate time series datasets.
-Our code is publicly available.
+Anomaly detection in time series analysis is a pivotal task, yet it poses the challenge of discerning normal and abnormal patterns in label-deficient scenarios.
+While prior studies have largely employed reconstruction-based approaches, which limits the models’ representational capacities.
+Moreover, existing deep learning-based methods are not sufficiently lightweight.
+Addressing these issues, we present PatchAD, our novel, highly efficient multiscale patch-based MLP-Mixer architecture that utilizes contrastive learning for representation extraction and anomaly detection.
+With its four distinct MLP Mixers and innovative dual project constraint module, PatchAD mitigates potential model degradation and offers a lightweight solution, requiring only 3.2MB.
+Its efficacy is demonstrated by state-of-the-art results across 9 datasets sourced from different application scenarios, outperforming over 30 comparative algorithms.
+PatchAD significantly improves the classical F1 score by 50.5%, the Aff-F1 score by 7.8%, and the AUC by 10.0%.
+The code is publicly available.
 [^1] https://github.com/EmorZz1G/PatchAD
 1 Introduction
-Time series anomaly detection is a pivotal aspect of data analysis, focused on identifying abnormal patterns within time-series data that significantly deviate from normal behavior Belay et al.
-(2023); Choi et al.
-(2021).
-This process becomes more complex in multivariate temporal anomaly detection, where it involves multiple channel dimensions.
-Driven by the rapid progression of large-scale sensing technologies and enhanced data storage capabilities, temporal anomaly detection has found widespread application in various sectors.
-For instance, in the field of the Internet of Things (IoT), it serves to monitor abnormal events in sensor data Yang et al.
-(2021).
-In healthcare sector, it plays a vital role in monitoring a person’s physiological data, thereby assisting in the early detection of health problems Cook et al.
-(2020).
-However, extracting anomalies from extensive and complex temporal data presents significant challenges.
-First, accurately defining the representations of anomalies is difficult due to the varied and evolving nature of real-world anomalies.
-For instance, anomalies in machinery could manifest as either high or fluctuating temperatures under different scenarios.
-Second, labels for anomalous events in time-series data are probably absent or hard to acquire.
-Thirdly, the high dimensionality and temporal dependencies inherent in time-series data are intricate.
-While the typical approach involves classifying different time points as normal or anomalous, relying solely on a single time point is often not feasible.
-Due to the powerful representational capacity of neural networks, numerous deep learning-based techniques Munir et al.
-(2019); Zhang et al.
-(2022) have been proposed for time series anomaly detection, demonstrating remarkable performance and attracting significant interest from the research community.
-While supervised and semi-supervised methods grapple with the scarcity of labeled data, unsupervised techniques Zhou et al.
-(2019); Su et al.
-(2019) have gained prominence due to their ability to operate without strict labeling requirements.
-While Xu et al.
-(2022) designed six different anomaly injection methods to train the model in a self-supervised approach, Carmona et al.
-(2021) utilized outlier exposure and mixup to construct a self-supervised task.
-Reconstruction-based approaches, as typical and common unsupervised methods, involve learning representations of historical data and identifying samples that cannot be accurately reconstructed during testing as anomalies Park et al.
-(2018).
-This approach has experienced rapid development due to its clear concept and its proficiency in managing complex data representations.
-However, due to their robust reconstruction capabilities, these methods excel at recovering anomalous data, posing a challenge for reconstruction-based approaches.
-The core issue is the ineffectiveness of these methods in distinctly modeling normal and abnormal data, complicating their ability to accurately identify anomalies.
-Recently, contrastive learning has been captivated by its straightforward, yet exceptional performance, proving effective in areas such as computer vision Chen et al.
-(2020).
-In time series anomaly detection, several methods have begun incorporating contrastive learning to enhance model representational capabilities.
-Nonetheless, these approaches often overlook the design of network structures tailored for contrastive learning.
-They either do not explicitly model inter-channel dependencies Yang et al.
-(2023) or rely solely on basic MLPs or Transformers that focus on temporal information Zhou et al.
-(2022), mainly considering point-level features Audibert et al.
-(2020).
-Nevertheless, poor training and inference speeds, architectural complexity, and a lack of lightweight features are some of the drawbacks of high-performing techniques.
-It was recently shown by Zeng et al.
-(2022) that MLP structures can also result in outstanding performance on time series tasks.
-Additionally, they neglected the model degradation issue in contrastive learning, meaning models might degrade or collapse during the training process.
-In this paper, we propose PatchAD, a novel Patch-based MLP architecture designed for time series Anomaly Detection.
-PatchAD leverages contrastive learning to tackle the complexities of time series anomaly detection.
-Inspired by Yang et al.
-(2023), we hypothesize that normal time series points share similar latent patterns, indicating a strong correlation among normal time points.
-On the contrary, anomalies struggle to establish connections with other anomaly points or normal points.
-Hence, normal data exhibits consistent representations across different views, while anomalies struggle to maintain consistent representations.
-This underlying principle forms the core of our approach to anomaly detection.
-Specifically, we introduce a novel Multi-scale Patching and Embedding approach.
-This method decomposes the input data along the temporal dimension into non-overlapping patches, enabling the model to learn from features that contain richer semantic information, moving beyond simple point-level features.
-To address various time scales, different submodels are utilized, each focusing on different patch sizes.
-To capture diverse views of temporal data representation, we propose different Patch Mixer Encoders to learn the mixed relationships among inter-patches, intra-patches, and inter-channel connections.
-We utilize a shared MixRep Mixer to uniformly represent the feature space.
-To prevent the degradation, we not only construct diverse encoders but also introduce a Dual Project Constraint mechanism to prevent the model from settling on overly simplistic solutions, thus substantially boosting its representational capabilities.
-The main contributions of our proposed PatchAD can be summarized as follows:
-  • The novel PatchAD model, innovatively structured on contrastive learning framework, is adept at managing diverse anomaly types.
-This is achieved by creating Interpatch Encoders and Intra-patch Encoders, which generate both inter-patch and intra-patch perspectives.
-  • Employing a strategy named Multi-scale Patching and Embedding, PatchAD integrates multi-scale subsequence modules constructed exclusively with MLPs.
-This design is intended to facilitate the learning of semantically richer features.
-  • Dual Project Constraint mechanism is specifically designed to impose constraints on the model, reducing its susceptibility to simple, trivial solutions and thus increasing its expressive power.
-  • PatchAD’s efficacy is underscored by comprehensive experimental results on eight datasets.
-These results highlight its superior performance compared to the SOTA time series anomaly detection methods.
-Figure 1: The workflow of the proposed PatchAD framework.
+Time series anomaly detection is a pivotal aspect of data analysis, focused on identifying abnormal patterns within time-series data that significantly deviate from expected behaviour [1, 2, 3].
+With the rapid progression of large-scale sensing technologies and storage capabilities, temporal anomaly detection has found widespread application in various sectors.
+For instance, the Internet of Things (IoTs) serve to monitor abnormal events in sensor data [4, 5].
+Unsupervised algorithms, capable of operating without labels, have gained wide popularity.
+These include reconstruction-based[6, 7], density-based[8, 9], contrastive learning[10, 11], autoregression-based[12, 13, 14], large language models (LLMs)[15] etc.
+However, extracting anomalies from extensive and complex temporal data presents significant challenges for different algorithms.
+Challenge 1: Accurately defining anomaly representations is challenging due to real-world anomalies’ diverse and dynamic nature.
+Effective anomaly detection requires models that can learn the interrelationships between various data channels.
+Challenge 2: There is still significant potential for improvement in terms of training and inference speed, model complexity, and the lack of lightweight features.
+Challenge 3: In an unsupervised setting, the challenge lies in establishing a reliable basis to model normal and anomalous data.
+Specifically, several algorithms have begun incorporating contrastive learning to enhance model representational capabilities.
+Nonetheless, these approaches often overlook the design of network structures and model degradation issue of contrastive learning.
+Additionally, they either do not explicitly model inter-channel dependencies [10] or rely solely on basic MLPs or Transformers that focus on temporal information [16], mainly considering point-level features [17].
+On the other hand, we find that normal time series data points exhibit reduced information entropy, indicating higher similarity and correlation among them, whereas anomalies show increased entropy due to difficulties in establishing connections with other points.
+This difference aids our model in detecting anomalies, ensuring that normal data maintains consistent representations across different views, whereas anomalies do not.
+This fundamental principle forms the basis of our proposed anomaly detection method.
+In response to the above challenges and principles, we introduce PatchAD, a lightweight Patch-based MLP architecture for time series Anomaly Detection.
+Specifically, proposed Multi-scale Patching and Embedding decomposes the input data along the temporal dimension into non-overlapping patches, extracting semantically richer features beyond simple point-level characteristics.
+Besides, various submodels address various time scales, each focusing on different patch sizes.
+Patch Mixer Encoders are introduced to learn the complex relationships within and between patches and channels.
+A shared MixRep Mixer uniformly represents the feature space, while a Dual Project Constraint prevents the model from adopting overly simplistic solutions, thereby enhancing its representational capabilities.
+In detail, this work makes the following contributions:
+Figure 1: Overall performances under different metrics.
+Figure 2: Model sizes.
+(1) State-of-the-art Anomaly Detection (Figure2): We propose a novel contrastive learning-based time series anomaly detection method.
+PatchAD achieves superior performance on 9 benchmark datasets, surpassing over 30 baselines across various evaluation metrics, with and without prior association.
+(2) Lightweight and Efficient Design (Figure2): PatchAD employs a multi-scale patch-based MLP-Mixer architecture with different novel strategies, achieving remarkable efficiency and compactness compared to Transformer-based models.
+(3) Comprehensive Analysis: An extensive ablation study validates the effectiveness of each component of PatchAD, supported by a theoretical analysis that elucidates the model’s underlying principles and design choices.
 2 Preliminaries
-In this section, we provide an initial exposition on the formal representation of time series anomaly detection, related methods, and contrastive learning.
-These are the foundational elements of the research problem and serve as the motivation for our work.
 2.1 Problem Formulation
-In this paper, we consider a multivariate time series of length $T$ in an unsupervised setting:
-$\mathcal{X}=(x_{1},x_{2},\cdots,x_{T}),$ (1)
-where $x_{t}\in\mathbb{R}^{C}$ represents a C-dimensional channel feature, such as data from C sensors or machines.
+In this paper, we consider a multivariate time series of length $T$ in an unsupervised setting: $\mathcal{X}=(x_{1},x_{2},\cdots,x_{T})$, $x_{t}\in\mathbb{R}^{C}$ represents a C-dimensional channel feature, such as data from C sensors or machines.
 The objective is to train a model using $\mathcal{X}_{train}$ and provide predictions $\mathcal{Y}_{test}=(y_{1},y_{2},\cdots,y_{T^{\prime}})$ on another sequence $\mathcal{X}_{test}$ of length $T^{\prime}$.
@@
 2.2 Time Series Anomaly Detection
-There exist various methods proposed for time series anomaly detection, encompassing statistical approaches, classical machine learning methods, and deep learning techniques.
+Figure 3: The workflow of the proposed PatchAD framework.
+The time series anomaly detection methods encompass statistical approaches, classical machine learning methods, and deep learning techniques[18].
 Statistical methods involve moving averages and the autoregressive integrated moving average (ARIMA) model.
-Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) Schölkopf et al.
-(2001) and Support Vector Data Description (SVDD) Tax and Duin (2004).
-Deep learning methods include DAGMM, OmniAnomaly Su et al.
-(2019), USAD Audibert et al.
-(2020), LSTM-VAE Park et al.
-(2018), and DeepAnT, which are trained on normal data and detect anomalies by inference using reconstruction errors.
-Some techniques Kim et al.
-(2023),Chen et al.
-(2022) have leveraged self-attention mechanisms and have demonstrated promising detection results.
-Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like Li et al.
-(2021a) introduced sequence variational autoencoder model based on smoothness constraints, whereas Xu et al.
-(2021) proposed prior and sequence associations for temporal modeling, both achieving enhanced performance.
-2.3 MLP Mixer for Time Series
-The MLP Mixer, initially proposed as a novel architecture for computer vision Tolstikhin et al.
-(2021), has gained significant attention in recent years.
+Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) [19] and Support Vector Data Description (SVDD) [20].
+Deep learning methods include DAGMM, OmniAnomaly [7], USAD [17], LSTM-VAE [6], and DeepAnT [21], which are trained on normal data and detect anomalies by inference using reconstruction errors.
+Some techniques have leveraged self-attention mechanisms and have demonstrated promising detection results [22, 23].
+While [24] designed six different anomaly injection methods to train the model in a self-supervised approach, [25] utilized outlier exposure and mixup to construct a self-supervised task.
+Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like [26] introduced sequence variational autoencoder model based on smoothness constraints, whereas [27] proposed prior and sequence associations for temporal modeling, both achieving enhanced performance.
+Recently approaches based on diffusion models and Large language model have also been proposed [28, 15].
+2.3 MLP Mixer & Contrastive Learning for Time Series Analysis
+The MLP Mixer initially proposed as a novel architecture for computer vision [29], has gained significant attention in recent years.
 Recently, there has been analysis indicating that MLP Mixers can effectively handle sequential data and find applications in other domains, such as time series prediction.
-Specifically, works like Nie et al.
-(2023); Gong et al.
-(2023) have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results.
+Specifically, works like [30, 31, 32] have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results.
 Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time-series forecasting (LSTF).
-To the best of our knowledge, there is currently an absence of MLP Mixer application in the time series anomaly detection.
-2.4 Contrastive Learning
-Contrastive learning has garnered considerable attention across various domains for its capacity to train models by contrasting positive pairs against negative ones.
-The objective is to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs.
-Additionally, methods like BYOL Grill et al.
-(2020) and SimSiam Chen and He (2021) have demonstrated advanced performance without explicitly constructing negative pairs.
-Recently, studies such as Eldele et al.
-(2021) and Pranavan et al.
-(2022) have explored the integration of contrastive learning into time series analysis.
-Yang et al.
-(2023), on the other hand, combines self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection.
+Contrastive learning aims to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs.
+Additionally, methods like BYOL [33] and SimSiam [34] have demonstrated advanced performance without explicitly constructing negative pairs.
+Recently, studies such as [11] and [35] have explored integrating contrastive learning into time series analysis.
+[10], on the other hand, combines self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection.
 3 Proposed Method
 3.1 Overall Architecture
-Figure 1 illustrates the overall structure of PatchAD.
-PatchAD adopts patch sizes at multiple scales.
+Figure 3 illustrates the overall structure of PatchAD, which adopts patch sizes at multiple scales.
 We exemplify its architecture using a single scale.
@@
 The focal point of our design lies in the four distinct MLP Mixers.
-Inspired by Yang et al.
-(2023), we utilize various MLP Mixers to learn different representations from the input.
+We utilize various MLP Mixers to learn different representations from the input.
 Normal points share a common latent space across different views, while anomalies, being scarce and lacking specific patterns, struggle to share a coherent representation with normal points.
@@
 PatchAD leverages inter-intra discrepancy to model this relationship.
+Our anomaly detection approach, PatchAD, leverages the principle that normal time series points, characterized by low information entropy, exhibit consistent representations across diverse perspectives.
+In contrast, anomalies with higher entropy struggle to maintain such consistency (Figure 3(c)).
+PatchAD employs an Inter-Patch Encoder to capture detailed features and an Intra-Patch Encoder to learn generalized trends.
+This combined approach enables a comprehensive understanding of normal and anomalous representations, achieving effective anomaly detection.
 3.2 Multi-Scale Patching and Embedding
@@
 The input data with $C$ channels and length $T$ is transformed into patches of size $P$.
-It can be seen as dividing the time series data into $N$ non-overlapping blocks of patches, which is called $\operatorname{Patching}(\cdot)$.
+It can be seen as dividing the time series data into $N$ non-overlapping patches blocks, called $\operatorname{Patching}(\cdot)$.
 Thus, the original data of dimensions $C\times T$ is transformed into $C\times N\times P$.
@@
 This encoder incorporates four distinct MLP Mixers to extract features across different dimensions.
-Figure 1(c) illustrates the MLP block we designed.
-It consists of two fully connected layers, a normalization layer, a GELU non-linear activation layer, and utilizes a residual connection from input to output.
+The MLP Mixer consists of two fully connected layers, a normalization layer and a GELU non-linear activation layer, and utilizes a residual connection from input to output.
 This can be represented as:
@@
   1. 1.
-    The Channel Mixer enables PatchAD to capture correlations among different channels.
+    Channel Mixer captures correlations among different channels.
   2. 2.
-    The Inter Mixer empowers PatchAD to capture inter-patch representations, facilitating the learning of global expressions.
+    Inter Mixer captures inter-patch representations, improving detailed expressions.
   3. 3.
-    The Intra Mixer enables PatchAD to capture intra-patch representations, facilitating the learning of local expressions.
+    Intra Mixer captures intra-patch representations, improving generalized expressions.
   4. 4.
-    The MixRep Mixer allows PatchAD to embed two distinct views into the same representation space.
+    MixRep Mixer embeds two distinct views into the same representation space.
 The diagram depicts PatchAD comprised of an inter-patch encoder and an intra-patch encoder, generating representations denoted as $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$, representing the inter-patch and intra-patch views, respectively.
@@
 This design facilitates PatchAD in learning differences between views and prevents model degradation.
+In the end, two separate simple MLPs are utilized to reconstruct the original data, which are then combined.
+$\hat{\mathcal{X}}_{1}=\operatorname{Rec1}(\mathcal{N}),\quad\hat{\mathcal{X}}_{2}=\operatorname{Rec2}(\mathcal{P}),\quad\hat{\mathcal{X}}=\hat{\mathcal{X}}_{1}+\hat{\mathcal{X}}_{2}.$ (4)
 3.4 Dual Project Constraint
@@
 Prior research suggests that employing a similar structure effectively prevents models from converging to trivial solutions.
-For instance, SimCLR Chen et al.
-(2020) uses a projection head that maps augmented views of an image into a latent space and promotes smoother convergence during training.
+For instance, SimCLR [36] uses a projection head that maps augmented views of an image into a latent space and promotes smoother convergence during training.
 The projection head comprises two layers of MLP connections, excluding non-linear and normalization layers.
@@
 The loss function for $\mathcal{N}$ and $\mathcal{P}$ can be defined as follows:
-$\mathcal{L}_{\mathcal{N}}\{\mathcal{P},\mathcal{N}\}$ $=\sum\operatorname{KL}(\mathcal{N},\operatorname{StopGrad}(\mathcal{P}))$ (7)
-$+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{P}),\mathcal{N}),$
-$\mathcal{L}_{\mathcal{P}}\{\mathcal{P},\mathcal{N}\}$ $=\sum\operatorname{KL}(\mathcal{P},\operatorname{StopGrad}(\mathcal{N}))$ (8)
-$+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{N}),\mathcal{P}),$
+$\mathcal{L}_{\mathcal{N}}\{\mathcal{P},\mathcal{N}\}=\sum\operatorname{KL}(\mathcal{N},\operatorname{StopGrad}(\mathcal{P}))+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{P}),\mathcal{N}),$ (7)
+$\mathcal{L}_{\mathcal{P}}\{\mathcal{P},\mathcal{N}\}=\sum\operatorname{KL}(\mathcal{P},\operatorname{StopGrad}(\mathcal{N}))+\operatorname{KL}(\operatorname{StopGrad}(\mathcal{N}),\mathcal{P}),$
 where $\operatorname{KL}(\cdot||\cdot)$ represents the KL divergence distance, and $\operatorname{StopGrad}$ denotes gradient stoppage, employed for asynchronously optimizing the two branches.
@@
 $\mathcal{L}_{proj}=\frac{\mathcal{L}_{\mathcal{N}^{\prime}}-\mathcal{L}_{\mathcal{P}}}{len(\mathcal{N})}+\frac{\mathcal{L}_{\mathcal{N}}-\mathcal{L}_{\mathcal{P}^{\prime}}}{len(\mathcal{N})}.$ (9)
-The final PatchAD loss consists of the aforementioned two components, with a constraint coefficient used to control the strength of the constraint:
-$\mathcal{L}=(1-c)\cdot\mathcal{L}_{cont}+c\cdot\mathcal{L}_{proj}.$ (11)
+The final PatchAD loss comprises the two aforementioned components and the reconstruction term ($\mathcal{L}_{rec}=\operatorname{MSE}(\hat{\mathcal{X}},\mathcal{X})$), which is accompanied by a constraint coefficient to regulate the strength of the constraint.
+$\mathcal{L}=(1-c)\cdot\mathcal{L}_{cont}+c\cdot\mathcal{L}_{proj}+\mathcal{L}_{rec}.$ (10)
 3.6 Anomaly Score
@@
 Based on the aforementioned anomaly score, we set a hyperparameter $\sigma$ to determine whether a point is considered an anomaly. If it surpasses the threshold, it is identified as an outlier.
-${y}_{i}=\left\{\begin{aligned} &1\text{, }\operatorname{AnomalyScore}(\mathcal{X}_{i})\geq\sigma\\ &0\text{, }\operatorname{AnomalyScore}(\mathcal{X}_{i})<\sigma.\end{aligned}\right.$ (13)
-Table 1: Comprehensive comparative results on four datasets.
-The P, R and F1 are the precision, recall and F1-score.
-All results are in %, the best ones are in Bold, and the second ones are underlined.
+Table 1: Comprehensive comparative results on four datasets. The P, R and F1 are the precision, recall and F1-score. All results are in %, the best in Bold, and the second in underlined.
 Dataset MSL SMAP PSM SWaT Average F1
 Metric P R F1 P R F1 P R F1 P R F1
-OCSVM Schölkopf et al. (2001) 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
-IForest Liu et al. (2008) 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
-LOF Breunig et al. (2000) 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
-U-Time Perslev et al. (2019) 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
-ITAD Shin et al. (2020) 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
-DAGMM Zong et al. (2018) 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
-VAR Anderson (1976) 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
-MMPCACD Yairi et al. (2017) 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
-CL-MPPCA Tariq et al. (2019) 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
-TS-CP2 Deldari et al. (2021) 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
-BeatGAN Zhou et al. (2019) 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
-LSTM-VAE Park et al. (2018) 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
-BOCPD Adams and MacKay (2007) 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
-Deep-SVDD Ruff et al. (2018) 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
-LSTM Hundman et al. (2018) 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
-OmniAnomaly Su et al. (2019) 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
-InterFusion Li et al. (2021b) 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
-THOC Shen et al. (2020) 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
-AnomalyTrans Xu et al. (2021) 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
-DCdetector Yang et al. (2023) 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
+OCSVM [19] 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
+IForest [37] 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
+LOF [8] 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
+U-Time [38] 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
+ITAD [39] 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
+DAGMM [9] 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
+VAR [12] 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
+MMPCACD [40] 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
+CL-MPPCA [41] 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
+TS-CP2 [42] 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
+BeatGAN [43] 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
+LSTM-VAE [6] 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
+BOCPD [44] 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
+Deep-SVDD [45] 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
+LSTM [46] 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
+OmniAnomaly [7] 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
+InterFusion [47] 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
+THOC [48] 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
+AnomalyTrans [27] 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
+DCdetector [10] 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
+TimesNet [14] 89.54 75.36 81.84 90.14 56.40 69.39 98.51 96.20 97.34 90.75 95.40 93.02 85.40
+PatchTST [49] 88.34 70.96 78.70 90.64 55.46 68.82 98.84 93.47 96.08 91.10 80.94 85.72 82.33
+GPT2-Adapter [15] 82.00 82.91 82.45 90.60 60.95 72.88 98.62 95.68 97.13 92.20 96.34 94.23 86.67
+NPSR-pt [50] 10.60 100.0 19.16 24.99 96.37 39.68 27.74 100.0 43.43 12.29 100.0 21.89 31.04
+NPSR-seq [50] 17.66 100.0 30.02 12.79 100.0 22.68 37.87 99.95 54.81 86.76 88.58 87.66 48.80
 PatchAD (Ours) 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 93.28 100.0 96.52 96.60
-Table 2: Overall results on NIPS-TS datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
-Dataset NIPS-TS-SWAN NIPS-TS-GECCO Average F1
-Metric P R F1 P R F1
-MatrixProfile Yeh et al. (2016) 16.70 17.50 17.10 4.60 18.50 7.40 12.25
-GBRT Elsayed et al. (2021) 44.70 37.50 40.80 17.50 14.00 15.60 28.20
-LSTM-RNN Bontemps et al. (2016) 52.70 22.10 31.20 34.30 27.50 30.50 30.85
-Autoregression Rousseeuw and Leroy (2005) 42.10 35.40 38.50 39.20 31.40 34.90 36.70
-OCSVM Schölkopf et al. (2001) 47.40 49.80 48.50 18.50 74.30 29.60 39.05
-IForest* Liu et al. (2008) 40.60 42.50 41.60 39.20 31.50 39.00 40.30
-AutoEncoder Sakurada and Yairi (2014) 49.70 52.20 50.90 42.40 34.00 37.70 44.30
-AnomalyTrans Xu et al. (2021) 90.70 47.40 62.30 25.70 28.50 27.00 44.65
-IForest Liu et al. (2008) 56.90 59.80 58.30 43.90 35.30 39.10 48.70
-AnomalyTrans Xu et al. (2021) 90.71 47.43 62.29 29.96 48.63 37.08 49.69
-DCdetector Yang et al. (2023) 96.57 59.08 73.31 38.41 59.73 46.76 60.04
-PatchAD (Ours) 96.75 59.15 73.41 40.62 62.88 49.35 61.38
 4 Experiments
 4.1 Setups and Evaluation Metrics
-We applied eight datasets for evaluation: (1) MSL (2) SMAP  [^2] MSL & SMAP: https://github.com/ML4ITS/mtad-gat-pytorch(3) PSM (4) SMD  [^3] https://github.com/NetManAIOps/OmniAnomaly (5) SWaT  [^4] https://itrust.sutd.edu.sg/itrust-labs_datasets (6) WADI  [^5] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi (7) NIPS-TS-SWAN (8) NIPSTS-GECCO  [^6] Others: github.com/DAMO-DI-ML/KDD2023-DCdetector.
+We applied eight multivariable and one univariable datasets for evaluation: (1) MSL (2) SMAP (3) PSM (4) SMD (5) SWaT (6) WADI (7) NIPS-TS-SWAN (8) NIPSTS-GECCO (9) UCR.
 Details about these datasets are outlined in Table 4 in Appendix.
 Also, implementation and pseudo-code Algorithm 1 details in Appendix.
-We compared our model with 27 benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) Classic methods; (6) Change point detection and time-series segmentation methods; (7) Transformer-based methods; See details in Table 1 and Table 2, or in Appendix Baslines.
-Additionally, we employed various evaluation metrics for assessing, encompassing commonly used metrics in time-series anomaly detection such as accuracy, precision, recall, and F1-score Xu et al.
-(2021).
-Moreover, we adopted state-of-the-art evaluation techniques like affiliation precision and affiliation recall Huet et al.
-(2022), along with Volume under the surface (VUS) Paparrizos et al.
-(2022).
+We compared our model with several benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) Classic methods; (6) Change point detection and time-series segmentation methods; (7) Transformer-based methods; (8) Large language model; (9) Diffusion model; See details in Table 1 and Table 5, or in Appendix B.2.
+Additionally, we employed various evaluation metrics for assessing, encompassing commonly used metrics in time-series anomaly detection such as AUC, accuracy, precision, recall, F1-score with point adjustment (F1), and original F1-score (F1_ori) [27].
+Moreover, we adopted state-of-the-art evaluation techniques like affiliation precision and affiliation recall [51], along with Volume under the surface (VUS) [52].
 Although the F1-score is a widely used metric, it does not effectively evaluate anomalous events.
@@
 Affiliation precision and recall are evaluation metrics based on the proximity of predicted events to the true labels.
-VUS considers the inclusion of anomalous events in its evaluation by computing the receiver operator characteristic (ROC) curve.
+VUS considers including anomalous events in its evaluation by computing the receiver operator characteristic (ROC) curve.
 Employing various methodologies allows for a more comprehensive evaluation of our model.
-To fairly compare with other methods, we utilized the point adjustment technique, identifying an anomalous segment if the model detects any anomaly point within the segment.
-(a) Project constraint
-(b) Anomaly ratio
-(c) $d_{model}$ size
-(d) Encoder layer
-(e) Window size
-(f) Multi-scale size
-Figure 2: Parameter sensitivity studies of main hyper-parameters in PatchAD.
+To compare with other methods fairly, we utilized the point adjustment technique, identifying an anomalous segment if the model detects any anomaly point within the segment.
 4.2 Comparison Results
-Data Source Clarification: Due to substantial device variations, to ensure fair comparison with anomaly transformer (AnomalyTrans) and DCdetector, all results mentioned in the paper concerning them are obtained through our reimplementation of their codes.
+Data Source Clarification: Due to substantial device variations, to ensure a fair comparison with anomaly transformer (AnomalyTrans) and DCdetector, all results mentioned in the paper concerning them are obtained through our reimplementation of their codes.
 The results of AnomalyTrans and DCdetector in Table 1 and Table 5 are our reproductions.
@@
 To comprehensively assess the performance of our approach, additional metrics such as affiliation precision/recall and VUS have been introduced.
-In Table 2, the NIPS-TS-SWAN and NIPS-TS-GECCO datasets represent more challenging datasets, encompassing a wider array of anomaly types.
-PatchAD consistently achieves SOTA results on these two datasets compared to other baseline methods.
+In Table 5 in Appendix, the NIPS-TS-SWAN and NIPS-TS-GECCO datasets represent more challenging datasets, encompassing a wider array of anomaly types.
+Compared to other baseline methods, PatchAD consistently achieves SOTA results on these two datasets.
 Particularly noteworthy is the 9.3% increase in F1-score over the DCdetector and a 24.9% increase over the Anomaly Transformer on the NIPS-TS-GECCO dataset.
@@
 Despite PatchAD’s relatively simple network structure, its performance remains competitive.
+Table 2 demonstrates that our approach outperforms previous methods on univariate dataset UCR.
+This shows that our method can also fetch state-of-the-art performance on univariate datasets.
+Table 2: Results on UCR dataset.
+Metric Acc P R F1
+AnomalyTrans 99.49 60.41 100 73.08
+DCdetector 99.51 61.62 100 74.05
+PatchAD (Ours) 99.52 72.72 100 84.00
+Figure 2 depicts the average performance of PatchAD compared to various baselines across multiple datasets (WADI, SWAT, PSM, NIPS-TS-SWAN).
+PatchAD has demonstrated a remarkable improvement in the F1 score, with a notable increase of 50.5%.
+Additionally, there has been a 7.8% improvement in the Aff-F1 score and a commendable 10.0% boost in the AUC.
 (a) Global point
@@
 Figure 4: Visualization of generated synthetic data, PatchAD anomaly scores and ground truth.
-(a) PatchAD’s loss analysis
+(a) Loss analysis
 (b) Time cost comparison
-(c) Time cost in different $d_{model}$
-(d) Time cost in different window
+(c) Different $d_{model}$
+(d) Different window
 Figure 5: The analysis of PatchAD and DCdetector.
+4.3 Unveiling the Mechanisms
+PatchAD leverages Inter and Intra Mixers to learn detailed and generalized features respectively, as demonstrated in Figure 8.
+Utilizing multiple patch sizes, as explored in Appendix (Figures 8(a) & 8(b)), enhances feature learning and generalization.
+The interplay between these mixers, visualized in Figure 3(c), effectively separates and amplifies distinct features, improving anomaly detection.
+Empirical validation, depicted in Figure 9, reveals that increasing patch size leads to more significant features acquired by the Intra Mixer, accelerating model learning.
+Finally, Figure 10 shows that anomalies are most easily detected in the Intra Mixer view due to its enhanced distinctiveness.
+In essence, maximizing the disparity between the Inter and Intra Mixers enhances PatchAD’s detection capabilities while mitigating overfitting.
+More details are provided in Appendix.
 4.4 Model Analysis
 4.4.1 Ablation Studies
+Table 3: Ablation studies on MSL, SMAP, WADI datasets.
+Dataset MSL SMAP WADI
+Methods Acc P R F1 Acc P R F1 Acc P R F1
+wo_ch_sharing 98.78 91.93 96.90 94.35 98.84 93.50 97.75 95.58 98.71 84.99 94.25 89.38
+wo_ch_mixer 98.78 92.03 96.82 94.36 98.77 93.50 97.10 95.27 98.92 85.40 97.93 91.24
+wo_mixrep_mixer 96.41 89.54 74.66 81.43 95.25 91.52 68.55 78.70 95.76 67.55 51.99 58.76
+wo_pos_emb 97.85 91.51 87.73 89.58 96.47 92.18 79.13 85.16 98.23 83.96 85.53 84.74
+wo_constraint 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 85.26 94.25 89.53
+PatchAD (Ours) 98.92 92.05 98.20 95.02 99.10 94.45 98.75 96.55 99.06 85.92 100.00 92.43
 To study the impact of different components in our model, we conducted ablation experiments across three datasets, as presented in Table 3.
-The ’wo_ch_sharing’ method employs two independent Channel Mixer modules to model channel features, while ’wo_ch_mixer’ denotes the removal of the Channel Mixer module.
-These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modeling inter-channel information and sharing channel information are beneficial for model learning.
+The ’wo_ch_sharing’ method employs two independent Channel Mixer modules to model channel features, while ’wo_ch_mixer’ denotes removing the Channel Mixer module.
+These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modelling inter-channel information and sharing channel information benefit model learning.
 Besides, ’wo_pos_emb’ represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design.
@@
 In subsequent experiments, the default setting for the projection head constraint is 0.2.
-Table 3: Ablation experiments on MSL, SMAP, WADI datasets.
-Dataset MSL SMAP WADI
-Methods Acc P R F1 Acc P R F1 Acc P R F1
-wo_ch_sharing 98.78 91.93 96.90 94.35 98.84 93.50 97.75 95.58 98.71 84.99 94.25 89.38
-wo_ch_mixer 98.78 92.03 96.82 94.36 98.77 93.50 97.10 95.27 98.92 85.40 97.93 91.24
-wo_pos_emb 97.85 91.51 87.73 89.58 96.47 92.18 79.13 85.16 98.23 83.96 85.53 84.74
-wo_constraint 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 85.26 94.25 89.53
-Ours 98.92 92.05 98.20 95.02 99.10 94.45 98.75 96.55 99.06 85.92 100.00 92.43
 4.4.2 Parameter Sensitivity
-We conducted a sensitivity analysis on PatchAD’s parameters.
-The default parameter settings remain consistent with the basic configuration.
-Figure 1(a) demonstrates the impact of the projection head constraint on PatchAD.
-Overall, it holds more significance for smaller-scale datasets like MSL.
-However, excessively strong constraints hinder the model’s learning process.
-Therefore, setting this constraint between 0.1 to 0.4 is more advisable.
-Figure 1(b) analyzes PatchAD’s final model performance when the anomaly threshold $\sigma$ ranges from 0.5 to 1.2.
-The selection of the anomaly threshold appears more robust on the PSM and SMAP datasets compared to the MSL dataset.
-For most datasets, setting this value between 0.8 and 1 yields optimal results.
-Figure 1(c) illustrates how the model’s performance is influenced by the network dimension $d_{model}$.
-PatchAD demonstrates better results with smaller $d_{model}$ sizes.
-Therefore, to strike a balance between model performance and complexity, we set $d_{model}=40$ to ensure PatchAD operates effectively.
-Figure 1(d) indicates that PatchAD’s performance is influenced by the number of encoder layers, and given a compromise, we set it to 3 or 4.
-Figure 1(e) demonstrates PatchAD’s robustness concerning the window size, which ranges from 60 to 210.
-The time window is a critical parameter, typically set to an initial value of 105 across most datasets.
-Lastly, Figure 1(f) displays PatchAD’s performance under different multi-scale patch size combinations.
-The outcomes suggest that combining various patch sizes can enhance PatchAD’s performance.
+Figure 6: Parameter sensitivity studies of main hyper-parameters in PatchAD.
+A sensitivity analysis was conducted to investigate the impact of various PatchAD parameters on model performance (Figure 6).
+We find that a projection head constraint within the range of 0.1 to 0.4 is generally advisable, particularly for smaller datasets (Figure 6(a)).
+Optimal anomaly thresholds ($\sigma$) typically fall between 0.8 and 1.0, demonstrating robustness across various datasets (Figure 6(b)).
+A network dimension ($d_{model}$) of 40 strikes a balance between model complexity and performance (Figure 6(c)).
+PatchAD exhibits consistent performance across a range of window sizes (Figure 6(e)), and employing 3 or 4 encoder layers is found to be suitable (Figure 6(d)).
+Finally, combining multiple patch sizes demonstrates a positive impact on overall performance (Figure 6(f)).
+Detailed analysis for each parameter is provided in Appendix.
 4.4.3 Visual Analysis
 We visualized PatchAD’s response to various anomalies in Figure 4.
-Synthetic univariate time series were generated by Lai et al.
-(2021), incorporating anomalies such as global point anomalies, contextual point anomalies, and pattern anomalies such as seasonal, group, and trend-based anomalies.
+Synthetic univariate time series were generated by [53], incorporating global point anomalies, contextual point anomalies, and pattern anomalies such as seasonal, group, and trend-based anomalies.
 The first row depicts the original data, while the second row displays PatchAD’s output of anomaly scores against the anomaly threshold.
 In particular, PatchAD has the ability to detect various types of anomalies.
+Figure 7 presents the practical performance of PatchAD on two real-world datasets.
+PatchAD exhibits the capability to detect the majority of anomalies, albeit with two anomalies remaining undetected.
+Furthermore, a noticeable disparity is observed between the reconstructed features by PatchAD and the original features, indicating its proficiency in learning the underlying patterns within the data, rather than simply replicating the original features.
+Figure 7: The case studies on SWAT and WADI datasets.
 4.4.4 Time Cost
@@
 Thus, increasing the window size or enlarging $d_{model}$ has a relatively minor impact on PatchAD.
-However, the time step noticeably affects the duration as this parameter leads to increased data volume.
+However, the time step noticeably affects the duration as this parameter increases data volume.
+4.4.5 Model Size
+Figure 2 compares the sizes of different models on SWAT dataset.
+Our model has a compact size of only 3.2MB, which is smaller than both AnomTrans and DCdetector, and significantly smaller than the LLM model (GPT2-Adapter) and diffusion model (D3R).
+This observation indicates that our model is more lightweight and resource-efficient.
 5 Conclusion
 We presents PatchAD, an innovative algorithm tailored for time series anomaly detection.
-PatchAD developed a multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning.
-This architecture is unique in its use of a purely MLP-based approach to discern both inter-patch and intra-patch relationships in time series sequences.
+PatchAD developed a lightweight multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning.
+This architecture is unique in using a purely MLP-based approach to discern inter-patch and intra-patch relationships in time series sequences.
 By amplifying the differences between two data views, PatchAD enhance its capability to differentiate between normal and anomalous patterns.
@@
 Moreover, the efficacy of PatchAD, including its various components, is validated through extensive experimentation.
-The results demonstrate that proposed PatchAD outperforms other methods across eight benchmark datasets.
+The results demonstrate that proposed lightweight PatchAD outperforms over 30 methods across nine benchmark datasets from different application scenarios.
 Appendix
-This is the appendix of PatchAD: Patch-based MLP-Mixer for Time Series Anomaly Detection.
-Datasets
-Many publicly available real-world datasets are extensively used in existing methods.
-We utilize eight datasets to evaluate model performance: (1) MSL (Mars Science Laboratory dataset), (2) SMAP (Soil Moisture Active Passive dataset), (3) PSM (Pooled Server Metrics), (4) SMD (Server Machine Dataset), (5) SWaT (Secure Water Treatment Testbed), (6) WADI (Water Distribution Testbed), (7) NIPS-TS-SWAN, and (8) NIPSTS-GECCO.
-Table 4 shows the details.
+This is the appendix of PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection.
+Appendix A Theory
+In this section, our objective is to perform a theoretical analysis of PatchAD’s effectiveness.
+We want to prove that normal time series data points exhibit reduced information entropy.
+We have made the following assumptions based on the theoretical analysis presented in previous studies.
+Assumption 1: The distribution of normal data differs from that of anomalies, and both consist of two components, denoted as $\tau+\alpha$.
+Here, $\tau$ represents the underlying data pattern, while $\alpha$ represents the pattern of normal or anomalous data.
+Therefore, we denote normal data distribution by $\Pi$ and anomalies as $\Omega$.
+As a result, the normal data $\mathcal{X}_{norm}$ and the anomalous data $\mathcal{X}_{anom}$ can be expressed as follows:
+$\mathcal{X}_{norm}=\tau+\pi,\text{ and }\mathcal{X}_{anom}=\tau+\omega.$ (12)
+Assumption 2: The number of normal data samples in the dataset is significantly larger than that of anomalous data samples. Therefore, we have the following relationship:
+$|\mathcal{X}_{norm}|=\zeta\cdot|\mathcal{X}_{anom}|>|\mathcal{X}_{anom}|,$ (13)
+where ||$\cdot$|| ⋅ | represents the number of samples in the data, and $\zeta$ is a large number. We denote the entire dataset as $\mathcal{D}=\{\mathcal{X}_{norm},\mathcal{X}_{anom}\}$.
+Proof 1: The entropy of the anomalous data is higher than that of normal data.
+For a random variable $X$, its entropy is defined as, and the entropy serves as a metric for quantifying the uncertainty and complexity of the data [54]:
+$\operatorname{H}(X)=-\sum_{i}p(x_{i})\log_{2}p(x_{i})$ (14)
+[…]
+Using the addition rule of probability, we have $p(\tau+\pi)=p(\tau)\cdot p(\pi)$.
+Substituting this into the previous equation, we obtain: $\operatorname{H}(\mathcal{X}_{norm})=-\int p(\tau)\cdot p(\pi)\log(p(\tau)\cdot p(\pi))d(\tau+\pi).$ Expanding further, we can deduce:
+$\operatorname{H}(\mathcal{X}_{norm})=-\int p(\tau)\cdot p(\pi)(\log p(\tau)+\log p(\pi))d(\tau+\pi).$ (15)
+[…]
+Due to the significantly larger number of samples in normal data compared to anomalous data, i.e., $|\mathcal{X}_{norm}|=\zeta\cdot|\mathcal{X}_{anom}|>|\mathcal{X}_{anom}|$, we can extract the large number $\zeta$ and use the properties of logarithm to move it inside the $\log$ function:
+$\operatorname{H}(\mathcal{X}_{norm})=-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log P^{\prime}(x)-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log\zeta$ (16)
+The second term $-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log\zeta$ is a constant term, denoted as $C$.
+Therefore, we can express the entropy of normal data as:
+$\operatorname{H}(\mathcal{X}_{norm})=-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log P^{\prime}(x)+C.$ (17)
+Similarly, we can derive the entropy of anomalous data $\operatorname{H}(\mathcal{X}_{anom})$:
+$\operatorname{H}(\mathcal{X}_{anom})=-\sum_{x\in\mathcal{X}_{anom}}P^{\prime}(x)\log P^{\prime}(x).$ (18)
+At this point, we can conclude that $\operatorname{H}(\pi)<\operatorname{H}(\omega)$.
+Therefore, the entropy of normal data $\operatorname{H}(\mathcal{X}_{norm})$ is smaller than the entropy of anomalous data $\operatorname{H}(\mathcal{X}_{anom})$.
+Appendix B Experimental Settings
+B.1 Datasets
+We applied eight multivariable and one univariable datasets for evaluation: (1) MSL (2) SMAP  [^2] MSL & SMAP: https://github.com/ML4ITS/mtad-gat-pytorch(3) PSM (4) SMD  [^3] https://github.com/NetManAIOps/OmniAnomaly (5) SWaT  [^4] https://itrust.sutd.edu.sg/itrust-labs_datasets (6) WADI  [^5] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi (7) NIPS-TS-SWAN (8) NIPSTS-GECCO  [^6] Others:http://github.com/DAMO-DI-ML/KDD2023-DCdetector.
+(9) UCR  [^7] https://www.cs.ucr.edu/~eamonn/time_series_data_2018/UCR_TimeSeriesAnomalyDatasets2021.zip Table 4 shows the details.
 Table 4: Details of benchmark datasets.
 Dataset #Training #Test (Labeled) Dimension Anomaly ratio (%)
-SML 58317 73729 55 10.5
+MSL 58317 73729 55 10.5
 SMAP 135183 427617 25 12.8
@@
 NIPS-TS-GECCO 69260 69261 9 1.1
-Baslines
-We compared our model against 27 benchmark models, including: (1) Reconstruction-based: AutoEncoder Sakurada and Yairi [2014], LSTM-VAE Park et al.
-[2018], OmniAnomaly Su et al.
-[2019], BeatGAN Zhou et al.
-[2019], InterFusion Li et al.
-[2021b]; (2) Autoregression-based: VAR Anderson [1976], Autoregression Rousseeuw and Leroy [2005], LSTM-RNN Bontemps et al.
-[2016], LSTM Hundman et al.
-[2018], CL-MPPCA Tariq et al.
-[2019]; (3) Density-based: LOF Breunig et al.
-[2000], MMPCACD Yairi et al.
-[2017], DAGMM Zong et al.
-[2018]; (4) Clustering-based: DeepSVDD Ruff et al.
-[2018], THOC Shen et al.
-[2020], ITAD Shin et al.
-[2020]; Matrix Profile Yeh et al.
-[2016]; (5) The classic methods: OCSVM Schölkopf et al.
-[2001], OCSVM-based subsequence clustering (OCSVM*), IForest Liu et al.
-[2008], IForest-based subsequence clustering (IForest*), Gradient boosting regression (GBRT) Elsayed et al.
-[2021]; (6) Change point detection and time series segmentation methods: BOCPD Adams and MacKay [2007], U-Time Perslev et al.
-[2019], TS-CP2 Deldari et al.
-[2021]; (7) Transformer-based: Anomaly Transformer  [^7] https://github.com/thuml/Anomaly-Transformer Xu et al.
-[2021]; DCdetector  [^8] https://github.com/DAMO-DI-ML/KDD2023-DCdetector Yang et al.
-[2023]; These two are current state-of-the-art methods.
-Implementation Details
-All experiments were conducted using PyTorch implementation on a single NVIDIA GTX 1080TI GPU.
+B.2 Baslines
+We compared our model against 35 benchmark models, including: (1) Reconstruction-based: AutoEncoder [55], LSTM-VAE [6], OmniAnomaly [7], BeatGAN [43], InterFusion [47], USAD [17], COUTA[24]; (2) Autoregression-based: VAR [12], Autoregression [56], LSTM-RNN [13], LSTM [46], CL-MPPCA [41], TimesNet [14], NPSR-pt and NPSR-seq [50]; (3) Density-based: LOF [8], MMPCACD [40], DAGMM [9]; (4) Clustering-based: DeepSVDD [45], THOC [48], ITAD [39]; Matrix Profile [57]; (5) The classic methods: OCSVM [19], OCSVM-based subsequence clustering (OCSVM*), IForest [37], IForest-based subsequence clustering (IForest*), Gradient boosting regression (GBRT) [58]; (6) Change point detection and time series segmentation methods: BOCPD [44], U-Time [38], TS-CP2 [42]; (7) Transformer-based: Anomaly Transformer  [^8] https://github.com/thuml/Anomaly-Transformer [27], DCdetector  [^9] https://github.com/DAMO-DI-ML/KDD2023-DCdetector [10], TranAD [59]; (8) Large Language Model: GPT2-Adapter[15]; (9) Diffusion model: D3R [60]; These two are current state-of-the-art methods.
+B.3 Implementation Details
+All experiments used PyTorch implementation on a single NVIDIA GTX 1080TI GPU.
 We summarize the default hyperparameters of PatchAD as follows: PatchAD consists of 3 encoder layers ($L$=3).
@@
 Different patch sizes and window sizes were chosen for each dataset, defaulting to [3, 5] and 105, respectively.
-If an anomaly score at any given timestamp exceeds a threshold $\sigma$, the model identifies it as an anomaly.
+If an anomaly score of any given timestamp exceeds a threshold $\sigma$, the model identifies it as an anomaly.
 The default value for $\sigma$ is 1.
-For details on parameter selection, please refer to the Appendix.
-Algorithm’s Pseudo-code
+Appendix C Algorithm’s Pseudo-code
 Algorithm 1 PatchAD’s Python-like code
@@
 # Upsample to (B, L, d_model)
+rec1 = Rec1(x_inter)
+rec2 = Rec2(x_intra)
+rec_x = rec1 + rec2
 # Update
@@
 loss_proj = loss_n_proj + loss_p_proj
-final_loss = (1-c) * loss1 + c * loss_proj
+loss_rec = MSE(rec_x, x)
+final_loss = (1-c) * loss1 + c * loss_proj + loss_rec
 # backward and optimize
-Full Comparison Results
-In Table 5, Aff-P and Aff-R are the precision and recall of affiliation metric Huet et al.
-[2022], respectively.
-R_A_R and R_A_P are Range-AUC-ROC and Range-AUC-PR Paparrizos et al.
-[2022], respectively.
-V_ROC and V_RR are volumes under the surfaces created based on ROC curve and PR curve Paparrizos et al.
-[2022], respectively.
-Besides, we give more results in table 6, and the results from more real-world multivariate time-series datasets show that our proposed PatchAD has good robustness as well as performance.
-Table 5: Multi-metrics results on all real-world multivariate datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
+Appendix D Unveiling the Mechanisms
+Figure 8 illustrates the knowledge acquired by PatchAD in four distinct cases.
+Cases 2, 3, 4 and 5 have a patch size of 1, whereas case 1 has a patch size of $[1,2,5]$.
+The "Original" represents the original data, while "Rec 1" and "Rec 2" correspond to the features learned by the Inter Mixer and the Intra Mixer, respectively.
+It is observed that the Intra Mixer captures the generalized knowledge in the data, which encompasses the overall trend information.
+Conversely, the Inter Mixer learnsthe detailed features of the data, compensating for any aspects that may not be captured by the Intra Mixer.
+A comparison between case 1 and case 2 reveals that using three distinct patch scales facilitates the acquisition of more detailed representations and enhances the generalization of the acquired features.
+Note that the y-axis ranges for Figure 8(a) and Figure 8(b) are different.
+Moreover, these two mixers collaborate to segregate and amplify the distinctive features effectively.
+The decoupling of features assists PatchAD in amplifying the anomaly-related characteristics, as indicated in Figure 3(c), thereby enhancing the detectability of anomalies.
+The superior performance of PatchAD can also be attributed to the varying sizes of the Inter and Intra Mixer branches, which prevent overfitting.
+This is the reason why PatchAD does not require the construction of negative samples in contrastive learning, as the two different networks construct implicit positive and negative samples.
+Moreover, we empirically validate this conclusion through experiments.
+Case 1 was chosen as the baseline, and we investigated the entropy of the features acquired by the Inter and Intra Mixers under varying patch size conditions.
+Figures 9(a) and 9(b) illustrate that both the Inter Mixer and Intra Mixer display a consistent trend of decreasing entropy throughout the learning process.
+This observation suggests that the models effectively acquire knowledge and reduce uncertainty.
+Furthermore, we noticed that the Inter Mixer demonstrates the highest learning speed and exhibits the most substantial decrease in entropy when the patch size is configured as [1, 2, 5].
+Additionally, as the patch size decreases, the entropy of the Inter Mixer increases, and it actually rises when the patch size is set to 1.
+This implies that the network is required to learn an excessive number of features.
+Figure 9(c) computes the ratio of Intra to Inter entropy, and our findings indicate that as the patch size increases, the features acquired by the Intra Mixer gain greater significance, thereby accelerating model learning through the utilization of multi-scale patch sizes.
+Figure 10 illustrates the visualization of feature distributions in the SWAT dataset under three conditions.
+It is observed that anomalies are more easily detected in the Intra Mixer view (See Figure 10(b) ), as its distinctiveness is more pronounced compared to the other two views (Figure 10 (a) & (c) ).
+According to our analysis, maximizing the disparities between the Inter Mixer and Intra Mixer can enhance the model’s detection capabilities.
+During network optimization, the Inter-Intra loss compels the Intra Mixer to acquire more generalized representations, effectively mitigating model overfitting.
+Figure 8: Visualization of the features learnt by Inter and Intra Mixer.
+Figure 9: Entropy analysis.
+Figure 10: Latent distribution’s visualization by T-SNE.
+Appendix E Parameter Sensitivity
+We conducted a sensitivity analysis on PatchAD’s parameters.
+The default parameter settings remain consistent with the basic configuration.
+Figure 6(a) demonstrates the impact of the projection head constraint on PatchAD.
+Overall, it holds more significance for smaller-scale datasets like MSL.
+However, excessively strong constraints hinder the model’s learning process.
+Therefore, setting this constraint between 0.1 to 0.4 is more advisable.
+Figure 6(b) analyzes PatchAD’s final model performance when the anomaly threshold $\sigma$ ranges from 0.5 to 1.2.
+The selection of the anomaly threshold appears more robust on the PSM and SMAP datasets than the MSL dataset.
+Setting this value between 0.8 and 1 for most datasets yields optimal results.
+Figure 6(c) illustrates how the model’s performance is influenced by the network dimension $d_{model}$.
+PatchAD demonstrates better results with smaller $d_{model}$ sizes.
+Therefore, to balance model performance and complexity, we set $d_{model}=40$ to ensure PatchAD operates effectively.
+Figure 6(d) indicates that PatchAD’s performance is influenced by the number of encoder layers, and given a compromise, we set it to 3 or 4.
+Figure 6(e) demonstrates PatchAD’s robustness concerning the window size, which ranges from 60 to 210.
+The time window is a critical parameter, typically set to an initial value of 105 across most datasets.
+Lastly, Figure 6(f) displays PatchAD’s performance under different multi-scale patch size combinations.
+(e.g., a patch size of [3,5] represents using two scales in PatchAD.) The outcomes suggest that combining various patch sizes can enhance PatchAD’s performance.
+Appendix F Visualization comparison
+Figure 11 shows the performance of WindowAD algorithms in the face of different types of anomalies.
+These methods have difficulty in detecting contextual and group point anomalies and are able to detect seasonal anomalies, but there are some false positives, which indicates that it has a low precision.
+(a) Contextual
+(b) Seasonal
+(c) Group
+Figure 11: Visualization of WindowAD-based algorithm’s score.
+Appendix G Full Comparison Results
+Table 6 presents the performance of different methods on four datasets, namely WADI, SWAT, PSM, and SWAN.
+This is the original data of Figure2.
+We have selected representative algorithms for experimental comparison.
+Table 5: Overall results on NIPS-TS datasets. All results are in %, the best in Bold, and the second in underlined.
+Dataset NIPS-TS-SWAN NIPS-TS-GECCO Average F1
+Metric P R F1 P R F1
+MatrixProfile [57] 16.70 17.50 17.10 4.60 18.50 7.40 12.25
+GBRT [58] 44.70 37.50 40.80 17.50 14.00 15.60 28.20
+LSTM-RNN [13] 52.70 22.10 31.20 34.30 27.50 30.50 30.85
+Autoregression [56] 42.10 35.40 38.50 39.20 31.40 34.90 36.70
+OCSVM [19] 47.40 49.80 48.50 18.50 74.30 29.60 39.05
+IForest* [37] 40.60 42.50 41.60 39.20 31.50 39.00 40.30
+AutoEncoder [55] 49.70 52.20 50.90 42.40 34.00 37.70 44.30
+AnomalyTrans [27] 90.70 47.40 62.30 25.70 28.50 27.00 44.65
+IForest [37] 56.90 59.80 58.30 43.90 35.30 39.10 48.70
+AnomalyTrans [27] 90.71 47.43 62.29 29.96 48.63 37.08 49.69
+DCdetector [10] 96.57 59.08 73.31 38.41 59.73 46.76 60.04
+PatchAD (Ours) 96.75 59.15 73.41 40.62 62.88 49.35 61.38
+Table 6: The performance of algorithms on different evaluation metrics without point adjustment.
+Dataset WADI SWAT
+Metric P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC
+Deep SVDD 4.05 38.53 7.34 50.06 55.79 52.77 40.53 12.74 98.50 22.57 53.58 99.99 69.77 86.81
+USAD 5.74 100.00 10.86 54.92 100.00 70.90 46.39 12.17 100.00 21.70 53.00 100.00 69.28 88.66
+COUTA 98.72 15.58 26.92 99.18 23.08 37.44 49.48 94.25 6.33 11.87 79.98 33.20 46.92 75.54
+TranAD 6.49 63.89 11.78 55.91 91.96 69.54 52.60 14.70 96.40 25.50 55.65 99.66 71.42 88.90
+TimesNet 64.04 10.24 17.65 65.25 39.30 49.05 65.08 11.05 3.21 4.98 57.84 93.77 71.55 26.76
+AnomTrans 4.14 14.12 6.40 45.73 40.66 43.05 49.20 4.81 29.32 8.27 53.03 98.08 36.08 19.20
+DCdetector 5.80 14.42 8.27 59.63 90.98 69.98 49.88 15.49 5.54 8.16 52.76 97.33 68.94 50.68
+GPT2-Adapter 5.68 5.50 5.59 55.62 98.29 71.04 50.01 12.16 0.96 1.78 52.51 98.13 68.41 50.00
+D3R 6.99 80.99 12.86 56.47 99.98 72.17 51.39 32.87 75.97 45.89 61.47 78.52 68.96 79.95
+PatchAD (Ours) 89.48(2.08) 31.22(0.74) 46.28(1.09) 73.07(0.41) 86.07(0.28) 79.04(0.12) 79.89(0.55) 85.42(0.98) 70.34(0.18) 77.15(0.48) 73.62(0.52) 84.65(0.22) 78.75(0.20) 91.19(0.12)
+Dataset PSM SWAN
+Metric P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC
+Deep SVDD 28.59 99.91 44.46 53.92 99.98 70.06 78.09 41.74 85.03 55.99 56.84 89.93 69.66 75.32
+USAD 32.36 92.17 47.90 55.30 98.71 70.88 64.53 39.73 77.94 52.63 55.26 68.62 61.22 66.76
+COUTA 100.00 0.04 0.08 100.00 5.12 9.75 72.45 62.96 0.09 0.17 69.25 0.10 0.19 71.38
+TranAD 35.90 54.23 43.20 60.88 92.09 73.30 65.22 38.84 78.29 51.93 57.86 90.18 70.49 70.00
+TimesNet 64.85 1.27 2.49 77.84 67.23 72.15 58.74 38.63 48.89 43.16 60.75 81.29 69.53 53.62
+AnomTrans 26.05 84.32 39.81 55.35 80.28 50.09 52.18 31.19 79.87 44.86 58.45 9.49 0.54 45.38
+DCdetector 27.05 19.59 22.72 52.33 76.95 65.88 49.62 32.45 98.62 48.83 38.73 3.13 7.50 50.35
+GPT2-Adapter 28.31 46.64 35.24 54.33 95.16 69.17 50.74 40.00 52.05 45.23 61.17 69.55 65.09 58.77
+D3R 28.96 93.65 44.24 53.84 100.00 69.99 60.64 36.12 78.38 49.45 57.70 96.06 72.10 62.54
+PatchAD (Ours) 47.13(0.44) 49.32(0.97) 48.20(0.64) 63.52(0.82) 96.84(1.21) 76.72(0.19) 64.74(0.01) 51.76(0.18) 65.31(0.47) 57.75(0.08) 66.48(0.80) 79.39(1.54) 72.36(0.20) 72.91(0.02)
+In Table 7, Aff-P and Aff-R are the precision and recall of affiliation metric [51], respectively.
+R_A_R and R_A_P are Range-AUC-ROC and Range-AUC-PR [52], respectively.
+V_ROC and V_RR are volumes under the surfaces created based on ROC curve and PR curve [52], respectively.
+Besides, we give more results in table 8, and the results from more real-world multivariate time-series datasets show that our proposed PatchAD has good robustness and performance.
+Table 7: Multi-metrics results on all real-world multivariate datasets. All results are in %, the best in Bold, and the second in underlined.
 Dataset Method Acc F1 Aff-P Aff-R R_A_R R_A_P V_ROC V_PR
@@
 PatchAD 98.64 49.35 53.04 89.48 63.59 35.92 63.09 35.43
-Table 6: More results on all real-world multivariate datasets. All results are in %, the best ones are in Bold, and the second ones are underlined.
+Table 8: More results on all real-world multivariate datasets. All results are in %, the best in Bold, and the second in underlined.
 Dataset MSL SMAP PSM SMD SWaT Average F1
@@
 [1, 3, 5, 7] 98.62 91.78 95.42 93.56 99.08 94.43 98.61 96.48 98.92 97.43 98.71 98.06
+Appendix H Limitations
+Although PatchAD achieves optimal results on multiple datasets, some limitations still need to be addressed.
+In order to ensure the optimization of different metrics, subtle adjustments to potential parameters are necessary.
+For instance, selecting an appropriate anomaly threshold is crucial to balance precision and recall.
+Developing an automated threshold adjustment scheme is a potential avenue for future research.
+Time series anomaly detection employs a wide range of evaluation metrics, each with its own emphasis, leading to potential biases in evaluating different algorithms.
+In the future, we aim to propose a unified set of evaluation metrics that consider multiple factors, contributing to the advancement of the community.
+Appendix I Broader Impacts
+Time series anomaly detection has significant academic implications and broader impacts in various domains.
+In finance, it aids in detecting fraudulent activities like credit card fraud and insider trading, safeguarding financial systems and preventing losses.
+In healthcare, it aids in early disease detection and monitoring, enabling timely interventions and improved patient outcomes.
+Overall, time series anomaly detection improves decision-making, risk mitigation, and system performance across multiple domains.

### v6

@@
 Abstract
-Anomaly detection in time series analysis is a pivotal task, yet it poses the challenge of discerning normal and abnormal patterns in label-deficient scenarios.
-While prior studies have largely employed reconstruction-based approaches, which limits the models’ representational capacities.
+Time series anomaly detection is a pivotal task in data analysis, yet it poses the challenge of discerning normal and abnormal patterns in label-deficient scenarios.
+While prior studies have largely employed reconstruction-based approaches, which limit the models’ representational capacities.
 Moreover, existing deep learning-based methods are not sufficiently lightweight.
 Addressing these issues, we present PatchAD, our novel, highly efficient multiscale patch-based MLP-Mixer architecture that utilizes contrastive learning for representation extraction and anomaly detection.
-With its four distinct MLP Mixers and innovative dual project constraint module, PatchAD mitigates potential model degradation and offers a lightweight solution, requiring only 3.2MB.
-Its efficacy is demonstrated by state-of-the-art results across 9 datasets sourced from different application scenarios, outperforming over 30 comparative algorithms.
-PatchAD significantly improves the classical F1 score by 50.5%, the Aff-F1 score by 7.8%, and the AUC by 10.0%.
-The code is publicly available.
-[^1] https://github.com/EmorZz1G/PatchAD
-1 Introduction
-Time series anomaly detection is a pivotal aspect of data analysis, focused on identifying abnormal patterns within time-series data that significantly deviate from expected behaviour [1, 2, 3].
-With the rapid progression of large-scale sensing technologies and storage capabilities, temporal anomaly detection has found widespread application in various sectors.
+With its four distinct MLP Mixers and innovative dual project constraint module, PatchAD mitigates potential model degradation and offers a lightweight solution, requiring only 0.403M parameters.
+Its efficacy is demonstrated by state-of-the-art results across 8 datasets sourced from different application scenarios, outperforming over 30 comparative algorithms.
+PatchAD significantly improves the classical F1 score by 6.84%, the Aff-F1 score by 4.27%, and the V-ROC by 2.49%.
+Simultaneously, an in-depth analysis of the mechanisms underlying PatchAD has been conducted from both theoretical and experimental perspectives, validating the design motivations of the model.
+The code is publicly available at https://github.com/EmorZz1G/PatchAD.
+Index Terms:
+Time series, anomaly detection, data mining, lightweight, MLP.
+I Introduction
+Time series anomaly detection (TSAD) is a pivotal aspect of data mining, focused on identifying abnormal patterns within time series data that significantly deviate from expected behaviour [1, 2, 3, 4].
+In the era of big data, with the rapid advancement of large-scale sensing technologies and the continuous enhancement of storage capabilities, TSAD has been widely applied in diverse sectors.
 For instance, the Internet of Things (IoTs) serve to monitor abnormal events in sensor data [5, 6].
-Unsupervised algorithms, capable of operating without labels, have gained wide popularity.
-These include reconstruction-based[6, 7], density-based[8, 9], contrastive learning[10, 11], autoregression-based[12, 13, 14], large language models (LLMs)[15] etc.
+Unsupervised algorithms, capable of operating without labels, have gained wide popularity [7].
+These include reconstruction-based [8, 9], density-based [10], contrastive learning [11, 12], autoregression-based [13, 14, 15], large language models (LLMs) [16] etc.
 However, extracting anomalies from extensive and complex temporal data presents significant challenges for different algorithms.
 Challenge 1: Accurately defining anomaly representations is challenging due to real-world anomalies’ diverse and dynamic nature.
-Effective anomaly detection requires models that can learn the interrelationships between various data channels.
+Effective anomaly detection requires models that can learn the inter-relationships between various data channels.
 Challenge 2: There is still significant potential for improvement in terms of training and inference speed, model complexity, and the lack of lightweight features.
 Challenge 3: In an unsupervised setting, the challenge lies in establishing a reliable basis to model normal and anomalous data.
-Specifically, several algorithms have begun incorporating contrastive learning to enhance model representational capabilities.
+Specifically, several algorithms [17, 11, 18, 19] have begun incorporating contrastive learning to enhance model representational capabilities.
 Nonetheless, these approaches often overlook the design of network structures and model degradation issue of contrastive learning.
+For instance, Xu et al.
+[17] devised a contrastive learning approach that leverages prior correlation and series correlation mining to elucidate the relationship between normal and anomalous patterns.
+However, this necessitated the employment of a maximization-minimization optimization strategy to train the model and prevent its degradation.
+Similarly, Yang et al.
+[11] extended the paradigm of image contrastive learning established by BYOL [20] by designing different temporal views to circumvent the challenges associated with negative sample construction in contrastive learning.
+Nonetheless, there exists the potential for performance degradation of the model when the patch settings are not appropriately configured.
 Additionally, they either do not explicitly model inter-channel dependencies [11] or rely solely on basic MLPs or Transformers that focus on temporal information [18], mainly considering point-level features [19].
-On the other hand, we find that normal time series data points exhibit reduced information entropy, indicating higher similarity and correlation among them, whereas anomalies show increased entropy due to difficulties in establishing connections with other points.
+However, with the advancement of TSAD evaluation metrics, an increasing number of interval-based evaluation indicators [21, 22, 23, 24] have garnered research attention.
+Nevertheless, prior studies predominantly concentrated on the F1 with point ajustment [17], thereby neglecting the detection of interval anomalies.
+Consequently, there remains considerable potential for enhancing the detection capabilities for interval anomalies in current research [24].
+Furthermore, the previous TSAD algorithms based on contrastive learning [17, 11, 12, 18] lack in-depth exploration from a theoretical perspective and fail to provide more valuable intuitions for future research.
+In contrast, we find that normal time series data points exhibit reduced information entropy, indicating higher similarity and correlation among them, whereas anomalies show increased entropy due to difficulties in establishing connections with other points.
 This difference aids our model in detecting anomalies, ensuring that normal data maintains consistent representations across different views, whereas anomalies do not.
 This fundamental principle forms the basis of our proposed anomaly detection method.
+Simultaneously, to further explore this concept, we conducted both experimental and theoretical analyses to investigate the underlying principles of this approach, thereby addressing the challenges from various perspectives.
 In response to the above challenges and principles, we introduce PatchAD, a lightweight Patch-based MLP architecture for time series Anomaly Detection.
@@
 A shared MixRep Mixer uniformly represents the feature space, while a Dual Project Constraint prevents the model from adopting overly simplistic solutions, thereby enhancing its representational capabilities.
+At the same time, the lightweight model configuration enables PatchAD to better tackle Challenge 2, while the intricate encoder design enhances the model’s feature learning capacity to confront Challenge 1.
 In detail, this work makes the following contributions:
 Figure 1: Overall performances under different metrics.
-Figure 2: Model sizes.
+Figure 2: Model Params, FLOPs, and Latency.
 (1) State-of-the-art Anomaly Detection (Figure2): We propose a novel contrastive learning-based time series anomaly detection method.
-PatchAD achieves superior performance on 9 benchmark datasets, surpassing over 30 baselines across various evaluation metrics, with and without prior association.
+PatchAD demonstrates exceptional performance across eight benchmark datasets, exceeding 30 baselines across a range of evaluation metrics, including commonly used metrics and interval anomaly indicators.
 (2) Lightweight and Efficient Design (Figure2): PatchAD employs a multi-scale patch-based MLP-Mixer architecture with different novel strategies, achieving remarkable efficiency and compactness compared to Transformer-based models.
-(3) Comprehensive Analysis: An extensive ablation study validates the effectiveness of each component of PatchAD, supported by a theoretical analysis that elucidates the model’s underlying principles and design choices.
-2 Preliminaries
-2.1 Problem Formulation
+Moreover, we conducted an analysis of the model’s scalability from both experimental and theoretical perspectives.
+(3) Comprehensive Analysis: An extensive ablation study validates the effectiveness of each component of PatchAD, supported by practical and theoretical analyses that elucidate the model’s underlying principles and design choices.
+II Preliminaries
+II-A Problem Formulation
 In this paper, we consider a multivariate time series of length $T$ in an unsupervised setting: $\mathcal{X}=(x_{1},x_{2},\cdots,x_{T})$, $x_{t}\in\mathbb{R}^{C}$ represents a C-dimensional channel feature, such as data from C sensors or machines.
@@
 Here, $y_{t}\in\{0,1\}$, where 1 signifies an anomaly at that time point and 0 indicates a normal point.
-2.2 Time Series Anomaly Detection
+II-B Time Series Anomaly Detection
 Figure 3: The workflow of the proposed PatchAD framework.
@@
 Classical machine learning methods encompass classification-based approaches like One-Class Support Vector Machine (OCSVM) [26] and Support Vector Data Description (SVDD) [27].
-Deep learning methods include DAGMM, OmniAnomaly [7], USAD [17], LSTM-VAE [6], and DeepAnT [21], which are trained on normal data and detect anomalies by inference using reconstruction errors.
-Some techniques have leveraged self-attention mechanisms and have demonstrated promising detection results [22, 23].
+However, machine learning methods often struggle to extract meaningful features and detect anomalies when confronted with high-dimensional, complex, and voluminous time series data, ultimately resulting in suboptimal performance [4, 28].
+Deep learning methods include OmniAnomaly [9], USAD [19], LSTM-VAE [8], and DeepAnT [29], which are trained on normal data and detect anomalies by inference using reconstruction errors.
+They often rely on reconstruction assumptions, wherein the model’s ability to restore anomalous data diminishes after being trained on normal data.
+Previous studies [30, 31] have indicated that this prior becomes less applicable when the model’s reconstruction capability is excessively strong, leading to a decline in anomaly detection performance.
+Some techniques have leveraged self-attention mechanisms and have demonstrated promising detection results [32].
 While [33] designed six different anomaly injection methods to train the model in a self-supervised approach, [34] utilized outlier exposure and mixup to construct a self-supervised task.
+However, due to the distributional discrepancies between real and synthetic data, this results in subpar performance when confronted with real-world data, as observed in [33, 34].
 Recognizing the limitations imposed by the reconstruction-based assumption, research efforts like [35] introduced sequence variational autoencoder model based on smoothness constraints, whereas [17] proposed prior and sequence associations for temporal modeling, both achieving enhanced performance.
 Recently approaches based on diffusion models and Large language model have also been proposed [36, 16].
-2.3 MLP Mixer & Contrastive Learning for Time Series Analysis
+Although significant advancements have been made in the performance of time series anomaly detection methods, most of these approaches are based on Transformers or sophisticated attention models [17, 11, 36, 16].
+This reliance results in models that are not sufficiently lightweight and exhibit slower computational efficiency, rendering them unsuitable for certain real-world scenarios that require rapid detection.
+II-C MLP Mixer & Contrastive Learning for Time Series Analysis
 The MLP Mixer initially proposed as a novel architecture for computer vision [37], has gained significant attention in recent years.
 Recently, there has been analysis indicating that MLP Mixers can effectively handle sequential data and find applications in other domains, such as time series prediction.
-Specifically, works like [30, 31, 32] have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results.
-Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time-series forecasting (LSTF).
-Contrastive learning aims to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs.
+Specifically, works like [38, 39, 40, 41] have focused on prediction tasks, particularly for multi-step time series prediction, leveraging the capabilities of MLP Mixers to achieve improved forecasting results.
+For example, PatchTST [39] employs a single-scale patch-based MLP as its core architecture, whereas the work in [41] applies an MLP to capture frequency-domain information in time series data for predictive modeling.
+Their experiments demonstrate that the MLP Mixer surpasses existing Transformer-based and LSTM-based methods in long sequence time series forecasting (LSTF).
+Constrained by space, further detailed descriptions about Mixer are provided in Appendix -A1.
+Contrastive learning aims to drive the model to bring positive pairs closer together in the learned representation space while pushing apart negative pairs [20, 42, 12].
 Additionally, methods like BYOL [20] and SimSiam [42] have demonstrated advanced performance without explicitly constructing negative pairs.
-Recently, studies such as [11] and [35] have explored integrating contrastive learning into time series analysis.
-[10], on the other hand, combines self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection.
-3 Proposed Method
-3.1 Overall Architecture
+Recently, studies such as [12] and [43] have explored integrating contrastive learning into time series representation learning.
+TS2Vec [44] performs contrastive learning in a hierarchical way over augmentation views.
+[17, 11], on the other hand, combine self-attention with contrastive learning, achieving state-of-the-art results in time series anomaly detection.
+In consideration of space limits, technical details are provided in Appendix -A2.
+III Proposed Method
+III-A Overall Architecture
 Figure 3 illustrates the overall structure of PatchAD, which adopts patch sizes at multiple scales.
@@
 This combined approach enables a comprehensive understanding of normal and anomalous representations, achieving effective anomaly detection.
-3.2 Multi-Scale Patching and Embedding
+III-B Multi-Scale Patching and Embedding
 To enhance the temporal contextual representation in time series, we incorporated Transformer positional encoding.
 Acknowledging the significance of different sequence lengths in the time series analysis, we introduced Multi-scale Patching in PatchAD.
-This enables PatchAD to focus on features of varying sequence lengths.
+Patching divides the time series window into smaller patches [39].
+Multi-scale Patching enables PatchAD to focus on features of varying sequence lengths.
 Moreover, incorporating multi-scale information helps compensate for information loss during the patching process.
 The input data with $C$ channels and length $T$ is transformed into patches of size $P$.
-It can be seen as dividing the time series data into $N$ non-overlapping patches blocks, called $\operatorname{Patching}(\cdot)$.
+It can be seen as dividing the time series data into $N$ non-overlapping patches blocks, called $\operatorname{Patching}(\cdot)$ [39].
 Thus, the original data of dimensions $C\times T$ is transformed into $C\times N\times P$.
-$\mathcal{X}=\operatorname{Patching}(\operatorname{PE}(\mathcal{X})),\ \mathcal{X}^{C\times T}\rightarrow\mathcal{X}^{C\times N\times P},$ (1)
-where $\operatorname{PE}(\cdot)$ is positional embedding.
+$\mathcal{X}=\operatorname{Patching}(\operatorname{PE}(\mathcal{X})),\ (\mathbb{R}^{C\times T}\rightarrow\mathbb{R}^{C\times N\times P}),$ (1)
+where $\operatorname{PE}(\cdot)$ is positional embedding [17].
 PatchAD employs two value embeddings, i.e., $\operatorname{VE}(\cdot)$, facilitating the features from two distinct perspectives within the framework.
+To make feature extraction and learning easier, these are used to embed the intra-patching and inter-patching information into the $D$-dimensional space, respectively.
 And $\operatorname{VE}(\cdot)$ is a single linear layer.
-$\mathcal{X}_{inter}$ $=\operatorname{VE}(\mathcal{X}),\ \mathcal{X}^{C\times N\times P}\rightarrow\mathcal{X}^{C\times N\times D},$ (2)
-$\mathcal{X}_{intra}$ $=\operatorname{VE}(\mathcal{X}),\ \mathcal{X}^{C\times N\times P}\rightarrow\mathcal{X}^{C\times P\times D}.$
-3.3 Patch Mixer Encoder
+$\mathcal{N}$ $=\operatorname{VE}(\mathcal{X}),\ (\mathbb{R}^{C\times N\times P}\rightarrow\mathbb{R}^{C\times N\times D}),$ (2)
+$\mathcal{P}$ $=\operatorname{VE}(\mathcal{X}),\ (\mathbb{R}^{C\times N\times P}\rightarrow\mathbb{R}^{C\times P\times D}).$
+For the sake of brevity and clarity, the above describes the process of single-scale patching. In the last part of this chapter, we describe how multi-scale patching is implemented (see Sec. III-G).
+III-C Patch Mixer Encoder
 PatchAD comprises $L$ layers of Patch Mixer layers.
 This encoder incorporates four distinct MLP Mixers to extract features across different dimensions.
-The MLP Mixer consists of two fully connected layers, a normalization layer and a GELU non-linear activation layer, and utilizes a residual connection from input to output.
-This can be represented as:
-$\mathcal{X}=\operatorname{FC}(\operatorname{GELU}(\operatorname{FC}(\operatorname{Norm}(\mathcal{X}))))+\mathcal{X}.$ (3)
+The Multi-Layer Perceptron (MLP) consists of two fully connected (FC) layers, a normalization layer and a ReLU activation, and utilizes a residual connection from input to output.
+The MLP can be represented as:
+$\operatorname{MLP}(\mathcal{X})=\operatorname{FC}(\operatorname{ReLU}(\operatorname{FC}(\operatorname{Norm}(\mathcal{X}))))+\mathcal{X}.$ (3)
+The MLP Mixers of different types utilize MLP to model the feature variations along different dimensions, which can be represented as:
+$\operatorname{Mixer}(\mathcal{X},dim=i)$ (4)
+$=\operatorname{Transpose}(\operatorname{MLP}(\operatorname{Transpose}(\mathcal{X},\text{dim = i})),\text{dim = i}),$
+where $\operatorname{Transpose}(\cdot,\text{dim = i})$ denotes the operation of swapping the $i$-th dimension with the last dimension.
+The purpose of this is to enable the MLP to effectively capture the features along different dimensions.
 The following are the different types of MLP Mixers utilized in PatchAD:
@@
 This design facilitates PatchAD in learning differences between views and prevents model degradation.
-In the end, two separate simple MLPs are utilized to reconstruct the original data, which are then combined.
-$\hat{\mathcal{X}}_{1}=\operatorname{Rec1}(\mathcal{N}),\quad\hat{\mathcal{X}}_{2}=\operatorname{Rec2}(\mathcal{P}),\quad\hat{\mathcal{X}}=\hat{\mathcal{X}}_{1}+\hat{\mathcal{X}}_{2}.$ (4)
-3.4 Dual Project Constraint
+Specifically, for each layer of the Patch Mixer Encoder, there exist two types of input features: inter-patch, denoted as $\mathcal{N}$, and intra-patch, represented as $\mathcal{P}$.
+Next, we elaborate on the workflow of four types of mixers in detail. Initially, these features are processed by the shared-parameter Channel Mixer, expressed as follows:
+[…]
+[…]
+[…]
+Similarly, the Inter Mixer and Intra Mixer utilize the outputs of the channel mixer as their respective inputs, yielding:
+[…]
+[…]
+Here, $\operatorname{InterMixer}$ and $\operatorname{InraMixer}$ denote the functions of the Inter Mixer and Intra Mixer, respectively.
+It is important to note that although both operate on feature variation at dim=2, the differing shapes of the input tensors result in $\mathcal{N}\in\mathbb{R}^{C\times N\times D}$ and $\mathcal{P}\in\mathbb{R}^{C\times P\times D}$.
+Consequently, $\operatorname{InterMixer}$ processes $\mathcal{N}$ along dimension $N$ to learn detailed features of patches, whereas $\operatorname{IntraMixer}$ processes $\mathcal{P}$ along dimension $P$ to acquire global knowledge of patches.
+Secs.
+V-A and V-B provide a detailed experimental and theoretical analysis of the nature of learning in both cases.
+Subsequently, to enable the model to learn low-dimensional invariant information from different types of features and thereby enhance the generalization capability of PatchAD, inspired by [45], the MixRep Mixer is employed to embed the two distinct types of features into a unified representation space, expressed as follows:
+[…]
+[…]
+[…]
+$\mathcal{N}=\operatorname{MixRepMixer}(\operatorname{InterMixer}(\operatorname{ChannelMixer}(\mathcal{N}))),$ (8)
+$\mathcal{P}=\operatorname{MixRepMixer}(\operatorname{IntraMixer}(\operatorname{ChannelMixer}(\mathcal{N}))).$
+In summary, the four mixer workflows correspond to Figure 3(b).
+In the end, two separate simple MLPs, e.i., $\operatorname{Rec_{1}}$ and $\operatorname{Rec_{2}}$ are utilized to reconstruct the original data, which are then combined.
+$\hat{\mathcal{X}}_{1}=\operatorname{Rec_{1}}(\mathcal{N}),\quad\hat{\mathcal{X}}_{2}=\operatorname{Rec_{2}}(\mathcal{P}),\quad\hat{\mathcal{X}}=\hat{\mathcal{X}}_{1}+\hat{\mathcal{X}}_{2}.$ (9)
+Simultaneously, the structures of $\operatorname{Rec_{1}}$ and $\operatorname{Rec_{2}}$ are identical.
+Initially, the last two channels of $\mathcal{N}$ and $\mathcal{P}$ are flattened using $\operatorname{Flatten}(\cdot)$.
+Subsequently, the features are reconstructed into the original time series $\mathcal{X}$ through a MLP.
+This process can be represented as:
+$\operatorname{Rec_{1}}(\mathcal{N})$ $=\operatorname{MLP}(\operatorname{Flatten}(\mathcal{N})),$ (10)
+$\operatorname{Rec_{2}}(\mathcal{P})$ $=\operatorname{MLP}(\operatorname{Flatten}(\mathcal{P})).$
+III-D Dual Project Constraint
 To prevent the model from converging to trivial solutions, each layer’s output requires to be constrained by Dual Project Head.
@@
 $\mathcal{P}_{l}$ $=\beta_{l}\cdot\mathcal{P}_{l},\ \beta_{l}=\operatorname{Softmax}(\beta).$
+Therefore, the outputs of the encoder can be represented as:
+$\mathcal{N}=\alpha_{1}\cdot\mathcal{N}_{1}+\cdots+\alpha_{L}\cdot\mathcal{N}_{L},$ (13)
+$\mathcal{P}=\alpha_{1}\cdot\mathcal{P}_{1}+\cdots+\alpha_{L}\cdot\mathcal{P}_{L}.$
 The introduction of ReWeight allows for a better balance between outputs from different layers, resulting in improved performance.
-3.5 Objective Function
+III-E Objective Function
 The inter-patch encoder and intra-patch encoder can represent two distinct views, denoted as $\mathcal{N}$ and $\mathcal{P}$.
@@
 Given the scarcity of anomalies in the data and the abundance of normal samples sharing hidden patterns, similar inputs should yield similar representations for both.
+In Sec.
+IV-C1, the reasons for choosing the KL divergence as the basic loss are analyzed.
 The loss function for $\mathcal{N}$ and $\mathcal{P}$ can be defined as follows:
@@
 $\mathcal{L}=(1-c)\cdot\mathcal{L}_{cont}+c\cdot\mathcal{L}_{proj}+\mathcal{L}_{rec}.$ (17)
-3.6 Anomaly Score
+III-F Anomaly Score
 Based on the assumption that normal points across different views share hidden patterns, while anomalies exhibit larger discrepancies, the difference between the two views can be employed as an anomaly score.
@@
 Based on the aforementioned anomaly score, we set a hyperparameter $\sigma$ to determine whether a point is considered an anomaly. If it surpasses the threshold, it is identified as an outlier.
-Table 1: Comprehensive comparative results on four datasets. The P, R and F1 are the precision, recall and F1-score. All results are in %, the best in Bold, and the second in underlined.
-Dataset MSL SMAP PSM SWaT Average F1
-Metric P R F1 P R F1 P R F1 P R F1
-OCSVM [19] 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
-IForest [37] 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
-LOF [8] 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
-U-Time [38] 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
-ITAD [39] 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
-DAGMM [9] 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
-VAR [12] 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
-MMPCACD [40] 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
-CL-MPPCA [41] 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
-TS-CP2 [42] 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
-BeatGAN [43] 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
-LSTM-VAE [6] 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
-BOCPD [44] 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
-Deep-SVDD [45] 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
-LSTM [46] 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
-OmniAnomaly [7] 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
-InterFusion [47] 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
-THOC [48] 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
-AnomalyTrans [27] 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
-DCdetector [10] 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
-TimesNet [14] 89.54 75.36 81.84 90.14 56.40 69.39 98.51 96.20 97.34 90.75 95.40 93.02 85.40
-PatchTST [49] 88.34 70.96 78.70 90.64 55.46 68.82 98.84 93.47 96.08 91.10 80.94 85.72 82.33
-GPT2-Adapter [15] 82.00 82.91 82.45 90.60 60.95 72.88 98.62 95.68 97.13 92.20 96.34 94.23 86.67
+III-G Multi-Scale of PatchAD
+For single-scale PatchAD, it features a comprehensive Patching, Patch Mixer Encoder, and Dual Project.
+This means that, assuming the patch size for single-scale PatchAD is $P_{i}$, we obtain the final output of the Patch Mixer Encoder, $\mathcal{\hat{X}}_{i}$, as well as the final outputs of the Dual Project, $\mathcal{N}_{i}$ and $\mathcal{P}_{i}$.
+These outputs are used to update the loss function $\mathcal{L}_{i}$ (as shown in Eq.
+(17)).
+Here, $i$ denotes the $i$-th single scale.
+For a multi-scale PatchAD, there are typically $M$ branches.
+We optimize the loss functions $\{\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{M}\}$ in an integrated manner, allowing different branches to learn features at various scales.
+During the testing phase, each single-scale branch produces its own anomaly score $AS_{i}$, and the final anomaly score is the average of the scores from all branches, given by $AS=\frac{1}{M}\sum_{i=1}^{M}AS_{i}$.
+IV Experiments
+IV-A Setups and Evaluations
+IV-A1 Datasets
+We evaluated different algorithms using a total of 8 datasets, with details referred to in the appendix.
+IV-A2 Metrics
+Considering comprehensively, we adopt two groups of metrics to conduct extensive experiments. Detailed explanations for why these metrics are used can be found in Appendix -D3.
+  1. 1.
+    Metrics using point adjustment (PA): We prioritize these metrics because they are now the most commonly utilized group.
+To compare with advanced algorithms (including but not limited to [11, 17, 9, 46, 19, 15]) that used point-adjusted metrics for evaluation, we employ Acc, Pre, Rec, and PA-F1 as the main metrics.
+  2. 2.
+    Metrics without point adjustment: Meanwhile, as many current studies [28, 24] are gradually using classic metrics (AUC, F1-cls) and interval evaluation metrics (Aff-F1, VUS), in order to compare with these algorithms (including but not limited to [33, 47, 16, 48, 49, 50, 11]), we organize additional experiments for further analysis in Sec.
+IV-B2.
+IV-A3 Baselines
+We compared our model against 30+ benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) The classic methods; (6) Change point detection and time series segmentation methods; (7) Contrastive learning; (8) Large Language Model; (9) Diffusion model.
+The details regarding the selection of the algorithms are further discussed in Appendix -D4.
+IV-B Comparison Results
+IV-B1 Comparisons via Metrics with PA
+From Table I, PatchAD achieves the best average-level PA-F1 performance.
+PatchAD attains superior results across multiple datasets and multiple metrics.
+Despite PatchAD’s relatively simple network structure, its performance remains competitive.
+In the evaluation of the PA-F1 metric, DCdetector and AnomTrans emerge as the state-of-the-art(SOTA) models, with their robust capability to capture local anomalies granting them a competitive edge over other models in PA-F1.
+In contrast, PatchAD enhances its applicability to local anomalies through adaptive learning of local feature representations, achieving SOTA performance.
+Compared with AnomTrans’s limited temporal granularity, PatchAD’s multi-scale approach can better capture anomalies of different lengths, thus achieving superior PA-F1 performance.
+Simultaneously, we observe that the additional models do not perform exceptionally well on the PA-F1 metric, primarily due to two reasons: firstly, they are mainly designed for the detection of interval anomalies, a fact that becomes evident in the subsequent analysis.
+Secondly, the PA-F1 metric employs point adjustment techniques, which can result in some undetected anomalies being classified as true positives, thereby introducing a potential inflation effect in the metric [51].
+While PA-F1 is suitable for comparing the performance of different models in research contexts, it is less appropriate for assessing model performance in real-world scenarios.
+Consequently, this also accounts for the challenge faced by PatchAD in distinguishing itself from previous models.
+Therefore, in the following section, a new set of metrics will be employed to analyze the models from different perspectives.
+TABLE I: Comprehensive comparative results on four datasets. The P, R and PA-F1 are the precision, recall and F1 with PA. All results are in %, the best in Bold, and the second in underlined.
+Dataset MSL SMAP PSM SWAT Average PA-F1
+Metric P R PA-F1 P R PA-F1 P R PA-F1 P R PA-F1
+LOF 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 72.15 65.43 68.62 64.50
+OCSVM 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 45.39 49.22 47.23 61.27
+U-Time 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 46.20 87.94 60.58 64.50
+IForest 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 49.29 44.95 47.02 63.12
+DAGMM 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 89.92 57.84 70.40 73.40
+ITAD 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 63.13 52.08 57.08 68.78
+VAR 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 81.59 60.29 69.34 74.80
+MMPCACD 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 82.52 68.29 74.73 75.93
+CL-MPPCA 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 76.78 81.50 79.07 76.05
+TS-CP2 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 81.23 74.10 77.50 79.91
+Deep-SVDD 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 80.42 84.45 82.39 81.44
+BOCPD 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 89.46 70.75 79.01 81.39
+LSTM-VAE 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 76.00 89.50 82.20 80.97
+BeatGAN 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 64.01 87.46 73.92 80.78
+LSTM 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 86.15 83.27 84.69 83.71
+OmniAnomaly 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 81.42 84.30 82.83 84.56
+InterFusion 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 80.59 85.58 83.01 85.57
+THOC 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 83.94 86.36 85.13 88.76
+AnomalyTrans 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 89.79 100.0 94.62 95.58
+DCdetetor 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 93.25 100.0 96.51 96.32
+TimesNet 89.54 75.36 81.84 90.14 56.40 69.39 98.51 96.20 97.34 90.75 95.40 93.02 85.40
+PatchTST 88.34 70.96 78.70 90.64 55.46 68.82 98.84 93.47 96.08 91.10 80.94 85.72 82.33
+GPT2-Adapter 82.00 82.91 82.45 90.60 60.95 72.88 98.62 95.68 97.13 92.20 96.34 94.23 86.67
 NPSR-pt [48] 10.60 100.0 19.16 24.99 96.37 39.68 27.74 100.0 43.43 12.29 100.0 21.89 31.04
 NPSR-seq [48] 17.66 100.0 30.02 12.79 100.0 22.68 37.87 99.95 54.81 86.76 88.58 87.66 48.80
-PatchAD (Ours) 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 93.28 100.0 96.52 96.60
-4 Experiments
-4.1 Setups and Evaluation Metrics
-We applied eight multivariable and one univariable datasets for evaluation: (1) MSL (2) SMAP (3) PSM (4) SMD (5) SWaT (6) WADI (7) NIPS-TS-SWAN (8) NIPSTS-GECCO (9) UCR.
-Details about these datasets are outlined in Table 4 in Appendix.
-Also, implementation and pseudo-code Algorithm 1 details in Appendix.
-We compared our model with several benchmark models, including: (1) Reconstruction-based; (2) Autoregression-based; (3) Density-based; (4) Clustering-based; (5) Classic methods; (6) Change point detection and time-series segmentation methods; (7) Transformer-based methods; (8) Large language model; (9) Diffusion model; See details in Table 1 and Table 5, or in Appendix B.2.
-Additionally, we employed various evaluation metrics for assessing, encompassing commonly used metrics in time-series anomaly detection such as AUC, accuracy, precision, recall, F1-score with point adjustment (F1), and original F1-score (F1_ori) [27].
-Moreover, we adopted state-of-the-art evaluation techniques like affiliation precision and affiliation recall [51], along with Volume under the surface (VUS) [52].
-Although the F1-score is a widely used metric, it does not effectively evaluate anomalous events.
-Therefore, we introduce the following two new metrics.
-Affiliation precision and recall are evaluation metrics based on the proximity of predicted events to the true labels.
-VUS considers including anomalous events in its evaluation by computing the receiver operator characteristic (ROC) curve.
-Employing various methodologies allows for a more comprehensive evaluation of our model.
-To compare with other methods fairly, we utilized the point adjustment technique, identifying an anomalous segment if the model detects any anomaly point within the segment.
-4.2 Comparison Results
-Data Source Clarification: Due to substantial device variations, to ensure a fair comparison with anomaly transformer (AnomalyTrans) and DCdetector, all results mentioned in the paper concerning them are obtained through our reimplementation of their codes.
-The results of AnomalyTrans and DCdetector in Table 1 and Table 5 are our reproductions.
-From Table 1, PatchAD achieves the best average-level F1 performance.
-As far as we know, there is ongoing controversy regarding the evaluation of time series anomaly detection.
-However, precision, recall, and F1-score remain the most widely used evaluation metrics.
-To comprehensively assess the performance of our approach, additional metrics such as affiliation precision/recall and VUS have been introduced.
-In Table 5 in Appendix, the NIPS-TS-SWAN and NIPS-TS-GECCO datasets represent more challenging datasets, encompassing a wider array of anomaly types.
-Compared to other baseline methods, PatchAD consistently achieves SOTA results on these two datasets.
-Particularly noteworthy is the 9.3% increase in F1-score over the DCdetector and a 24.9% increase over the Anomaly Transformer on the NIPS-TS-GECCO dataset.
-Given the remarkable performance of AnomalyTrans and DCdetector surpassing other baselines significantly, we conducted a comprehensive comparison of PatchAD with them across all datasets using multiple metrics, as depicted in Table 7 in Appendix.
-Our proposed method outperforms or at least remains competitive compared to these two SOTA methods.
-PatchAD attains superior results across multiple datasets and multiple metrics.
-Despite PatchAD’s relatively simple network structure, its performance remains competitive.
-Table 2 demonstrates that our approach outperforms previous methods on univariate dataset UCR.
-This shows that our method can also fetch state-of-the-art performance on univariate datasets.
-Table 2: Results on UCR dataset.
-Metric Acc P R F1
-AnomalyTrans 99.49 60.41 100 73.08
-DCdetector 99.51 61.62 100 74.05
-PatchAD (Ours) 99.52 72.72 100 84.00
-Figure 2 depicts the average performance of PatchAD compared to various baselines across multiple datasets (WADI, SWAT, PSM, NIPS-TS-SWAN).
-PatchAD has demonstrated a remarkable improvement in the F1 score, with a notable increase of 50.5%.
-Additionally, there has been a 7.8% improvement in the Aff-F1 score and a commendable 10.0% boost in the AUC.
+M2N2 [52] 89.38 92.48 90.90 87.37 72.15 79.04 70.51 89.99 79.07 99.42 70.96 82.82 82.96
+AdaMemBLS [4] 88.01 93.20 90.53 88.01 93.20 90.53 89.59 95.45 92.43 93.19 82.78 87.68 90.29
+PatchAD 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 93.28 100.0 96.52 96.60
 (a) Global point
@@
 Figure 4: Visualization of generated synthetic data, PatchAD anomaly scores and ground truth.
-(a) Loss analysis
-(b) Time cost comparison
-(c) Different $d_{model}$
-(d) Different window
-Figure 5: The analysis of PatchAD and DCdetector.
-4.3 Unveiling the Mechanisms
-PatchAD leverages Inter and Intra Mixers to learn detailed and generalized features respectively, as demonstrated in Figure 8.
-Utilizing multiple patch sizes, as explored in Appendix (Figures 8(a) & 8(b)), enhances feature learning and generalization.
-The interplay between these mixers, visualized in Figure 3(c), effectively separates and amplifies distinct features, improving anomaly detection.
-Empirical validation, depicted in Figure 9, reveals that increasing patch size leads to more significant features acquired by the Intra Mixer, accelerating model learning.
-Finally, Figure 10 shows that anomalies are most easily detected in the Intra Mixer view due to its enhanced distinctiveness.
-In essence, maximizing the disparity between the Inter and Intra Mixers enhances PatchAD’s detection capabilities while mitigating overfitting.
-More details are provided in Appendix.
-4.4 Model Analysis
-4.4.1 Ablation Studies
-Table 3: Ablation studies on MSL, SMAP, WADI datasets.
-Dataset MSL SMAP WADI
-Methods Acc P R F1 Acc P R F1 Acc P R F1
-wo_ch_sharing 98.78 91.93 96.90 94.35 98.84 93.50 97.75 95.58 98.71 84.99 94.25 89.38
-wo_ch_mixer 98.78 92.03 96.82 94.36 98.77 93.50 97.10 95.27 98.92 85.40 97.93 91.24
-wo_mixrep_mixer 96.41 89.54 74.66 81.43 95.25 91.52 68.55 78.70 95.76 67.55 51.99 58.76
-wo_pos_emb 97.85 91.51 87.73 89.58 96.47 92.18 79.13 85.16 98.23 83.96 85.53 84.74
-wo_constraint 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 85.26 94.25 89.53
-PatchAD (Ours) 98.92 92.05 98.20 95.02 99.10 94.45 98.75 96.55 99.06 85.92 100.00 92.43
-To study the impact of different components in our model, we conducted ablation experiments across three datasets, as presented in Table 3.
-The ’wo_ch_sharing’ method employs two independent Channel Mixer modules to model channel features, while ’wo_ch_mixer’ denotes removing the Channel Mixer module.
+IV-B2 Comparisons via Metrics without PA
+Figure 2 depicts the average performance of PatchAD compared to various baselines across multiple datasets (MSL, SMAP, SWAT, WADI, PSM, and SWAN).
+Here, Aff-F1 is the affiliation metric [22].
+V-ROC and V-RR are volumes under the surfaces created based on ROC curve and PR curve [21], respectively.
+In terms of the F1-cls, Aff-F1, and V-ROC metrics, PatchAD demonstrates relative performance improvements of 6.84%, 4.27%, and 2.49% compared to the second-place model, respectively.
+This indicates a significant enhancement in PatchAD’s capacity for detecting interval anomalies compared to previous models.
+Notably, we observe that AnomTrans and DCdetector, which perform well on the PA-F1 metric, exhibit poor performance on interval anomaly metrics.
+This is primarily because these models are more focused on local or point anomalies, resulting in better performance on the PA-F1 metric.
+PatchAD, on the other hand, benefits from additional learning of generalized global features, enhancing its performance on interval anomalies, thus improving both PA-F1 and Aff-F1 metrics.
+Furthermore, we note that although TimesNet was proposed earlier, it currently exhibits relatively strong performance when considering the two sets of metrics.
+This can be attributed to the diversity of contemporary anomalies, which vary in type, periodicity, and length.
+TimesNet’s capability to learn temporal information across different frequencies enables it to handle various anomalies more effectively.
+PatchAD similarly adopts this approach, with multi-scale temporal feature extraction ensuring the learning of information across different temporal frequencies.
+NPSR also takes into account the utilization of two distinct feature types in its design and uniquely processes anomaly scores, thereby ensuring its excellent theoretical lower bound, which guarantees good performance in the face of real interval anomalies.
+Likewise, PatchAD benefits from the assurance provided by the lower bound of information entropy and the diverse features extracted through different Mixers, resulting in similarly outstanding performance.
+In contrast, COUTA’s training relies on anomaly injection, leading to varying adaptability across different real datasets; for instance, it performs well on the PSM and SWAN datasets but shows average performance on the WADI and SWAT datasets.
+TABLE II: Comprehensive comparative results on six datasets. All results are in %, the best in Bold, and the second in underlined.
+Datasets MSL SMAP SWAT
+Methods AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR
+LOF [2000] 55.75 19.36 66.86 60.85 17.19 62.39 23.64 58.74 57.56 18.59 84.59 58.22 5.25 66.67 41.88
+IForest [2008] 59.09 10.90 70.44 64.74 18.29 61.15 23.78 50.52 59.31 16.45 83.80 38.52 5.27 72.29 52.07
+PCA [2003] 53.25 10.33 70.72 59.42 18.08 58.70 22.96 50.72 43.90 11.69 81.85 26.05 5.25 61.87 44.75
+Deep SVDD [2018] 61.80 24.30 68.98 59.54 19.36 61.14 21.28 59.64 41.95 11.56 86.81 22.57 69.77 50.96 56.48
+USAD [2020] 57.12 22.50 67.69 58.86 19.77 62.79 16.57 50.41 33.78 11.28 88.66 21.70 69.28 84.38 62.74
+TCN-ED [2021] 51.14 19.66 67.93 58.62 18.13 58.91 22.90 67.89 43.79 11.73 89.23 21.65 69.26 61.53 57.94
+TranAD [2022] 50.03 22.81 68.97 57.25 17.55 59.74 22.74 50.51 45.15 11.98 88.90 25.50 71.42 86.25 64.19
+NCAD [2022] 60.20 22.05 67.25 67.29 21.32 53.45 23.09 68.32 48.37 13.22 82.92 68.72 73.89 76.76 44.91
+Deep IF [2023] 55.94 19.11 67.94 52.33 17.04 60.09 29.14 69.59 56.89 14.77 89.52 21.65 69.26 50.19 56.39
+AnomTrans [2021] 52.61 18.39 66.81 53.03 13.21 52.18 16.06 71.44 52.61 16.55 80.80 23.44 66.77 21.49 8.29
+TimesNet [2022] 57.18 21.24 67.64 64.63 20.67 53.73 24.37 66.19 48.83 12.91 73.24 21.66 71.55 31.10 10.29
+GPT2-Ada. [2023] 52.03 13.72 68.81 53.57 17.38 55.48 24.12 69.48 57.31 15.78 52.30 22.30 68.41 50.01 10.29
+DCdetector [2023] 50.31 11.62 67.85 52.00 14.68 58.50 26.56 71.37 60.38 16.87 52.78 23.24 68.56 50.84 14.68
+D3R [2024] 63.00 23.98 69.79 69.02 20.93 54.56 22.71 67.92 48.67 13.23 79.95 45.89 68.96 84.38 57.27
+COUTA [2024] 55.59 20.88 67.68 61.49 18.68 58.74 22.69 67.97 44.72 11.91 75.54 47.65 46.92 72.57 25.06
+NPSR [2023] 61.16 23.72 68.42 66.44 19.72 38.74 22.68 67.95 41.91 11.42 90.18 76.88 75.86 84.93 65.42
+M2N2 [2024] 59.15 21.76 67.88 65.40 21.55 54.05 22.92 67.95 48.74 14.36 67.79 38.50 73.44 61.14 26.16
+AdaMemBLS [2025] 51.78 19.11 71.41 57.46 16.06 53.66 26.93 69.49 55.57 14.26 81.61 74.10 69.76 82.00 66.12
+PatchAD 60.94 24.92 70.10 66.43 19.65 60.27 24.74 70.95 48.30 13.71 91.19 77.69 78.76 89.91 65.74
+Datasets WADI PSM NIPS-TS-Swan
+Methods AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR AUC F1-cls Aff-F1 V-ROC V-PR
+LOF [2000] 65.44 10.58 34.36 66.67 41.88 81.28 23.91 62.32 80.42 64.55 77.89 21.37 19.62 89.42 88.56
+IForest [2008] 74.84 30.88 67.90 72.29 52.07 75.12 38.52 56.77 75.19 55.50 85.12 46.18 27.81 91.27 90.98
+PCA [2003] 49.47 34.30 50.01 61.87 44.75 72.06 43.20 78.74 73.97 54.55 72.88 49.18 27.35 90.42 90.23
+Deep SVDD [2018] 59.47 7.34 52.77 50.96 56.48 78.09 44.46 70.06 53.97 66.92 75.32 55.99 69.66 80.89 91.74
+USAD [2020] 53.61 10.86 70.90 84.38 62.74 64.53 47.90 70.88 64.02 69.18 66.76 52.63 61.22 91.40 94.50
+TCN-ED [2021] 53.01 10.86 70.90 61.53 57.94 63.97 43.46 69.42 57.48 42.64 71.37 49.22 71.03 80.88 91.73
+TranAD [2022] 52.60 11.78 69.54 86.25 64.19 65.22 43.20 73.30 72.42 56.02 70.00 51.93 70.49 88.01 92.59
+NCAD [2022] 62.84 10.87 70.96 76.76 44.91 61.76 46.61 63.10 67.09 48.26 52.24 37.57 68.87 85.31 93.44
+Deep IF [2023] 51.43 10.86 70.90 50.19 56.39 69.06 43.47 69.41 67.71 45.98 73.30 49.19 71.05 80.89 91.74
+AnomTrans [2021] 52.98 11.24 66.77 21.49 8.29 52.18 39.81 68.13 49.12 31.56 54.62 44.86 69.11 83.56 81.64
+TimesNet [2022] 65.08 17.65 49.05 31.10 10.29 58.74 43.60 72.15 66.37 48.05 53.62 43.16 69.53 91.58 94.30
+GPT2-Ada. [2023] 51.21 10.68 71.04 50.01 10.29 51.23 35.24 69.17 52.45 32.79 58.77 45.23 65.09 84.31 84.23
+DCdetector [2023] 50.12 11.33 73.23 50.84 14.68 50.38 22.72 68.08 51.73 32.44 50.35 48.83 70.99 85.84 83.11
+D3R [2024] 51.39 12.86 72.17 84.38 57.27 60.64 44.24 69.99 66.92 49.38 62.54 49.45 72.10 89.27 88.29
+COUTA [2024] 50.52 26.92 37.44 72.57 25.06 69.36 48.21 73.21 64.50 43.15 71.38 47.19 66.54 82.15 92.26
+NPSR [2023] 80.19 50.67 75.74 84.93 65.42 70.67 51.02 70.56 70.55 52.12 62.61 49.30 71.07 90.90 90.35
+M2N2 [2024] 55.68 13.55 70.90 43.99 7.98 54.91 47.71 69.84 40.06 30.61 70.27 57.47 71.05 89.77 90.18
+AdaMemBLS [2025] 53.85 29.09 70.93 63.47 21.40 63.26 46.69 71.51 65.07 49.24 51.27 49.35 71.09 87.12 85.92
+PatchAD 85.27 52.82 79.04 88.29 49.60 64.74 49.33 76.73 65.11 40.72 79.53 64.52 72.36 95.62 91.52
+IV-C Model Analysis
+IV-C1 Ablation Studies
+Ablation on Components: To study the impact of different components in our model, we conducted ablation experiments across four datasets, as presented in Table III.
+In a manner analogous to the comparative experimental setup, we selected two sets of metrics to analyze the impact of components on the model from different perspectives.
+The first set comprises point-adjusted metrics, including Acc, Pre, Rec, and PA-F1.
+These metrics exhibit superior adaptability to noisy labels and demonstrate greater robustness in the presence of mislabeled data [28, 24].
+The second set consists of metrics that do not employ point adjustments, including AUC, F1-cls, V-ROC, and V-PR.
+While AUC and F1-cls exhibit poorer robustness to label noise and lower flexibility, their stricter evaluation criteria ensure a lower bound on the accuracy of the assessment [24].
+Conversely, V-ROC and V-PR are two interval-based evaluation metrics that provide more accurate assessments for longer intervals of anomalies, thereby enhancing comprehensiveness.
+The “wo_ch_sharing” method employs two independent Channel Mixer modules to model channel features, while “wo_ch_mixer” denotes removing the Channel Mixer module.
 These two variants exhibit significant effects, particularly on the MSL and SMAP datasets, indicating that modelling inter-channel information and sharing channel information benefit model learning.
-Besides, ’wo_pos_emb’ represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design.
-Furthermore, ’wo_constraint’ showcases the impact of removing the constraint on the projection head.
-In conjunction with Figure 5(a), we observed that a certain level of constraint facilitates easier model optimization, leading to improved results.
+Besides, “wo_pos_emb” represents a variant without positional encoding, and removing positional encoding significantly affects PatchAD, highlighting the crucial role of positional encoding in its design.
+Furthermore, “wo_constraint” showcases the impact of removing the constraint on the projection head.
+In conjunction with Figure 5, we observed that a certain level of constraint facilitates easier model optimization, leading to improved results.
 This constraint prevents the model from converging to trivial solutions.
 In subsequent experiments, the default setting for the projection head constraint is 0.2.
-4.4.2 Parameter Sensitivity
-Figure 6: Parameter sensitivity studies of main hyper-parameters in PatchAD.
-A sensitivity analysis was conducted to investigate the impact of various PatchAD parameters on model performance (Figure 6).
-We find that a projection head constraint within the range of 0.1 to 0.4 is generally advisable, particularly for smaller datasets (Figure 6(a)).
-Optimal anomaly thresholds ($\sigma$) typically fall between 0.8 and 1.0, demonstrating robustness across various datasets (Figure 6(b)).
-A network dimension ($d_{model}$) of 40 strikes a balance between model complexity and performance (Figure 6(c)).
-PatchAD exhibits consistent performance across a range of window sizes (Figure 6(e)), and employing 3 or 4 encoder layers is found to be suitable (Figure 6(d)).
-Finally, combining multiple patch sizes demonstrates a positive impact on overall performance (Figure 6(f)).
-Detailed analysis for each parameter is provided in Appendix.
-4.4.3 Visual Analysis
+Considering the limitation of insufficient real labels for threshold selection in practical scenarios, we validated the practicality of PatchAD using the dynamic threshold algorithm based on Extreme Value Theory (SPOT) [53], as shown in the second-to-last row of Table III.
+Results indicate that its performance gap with the optimal model is minimal, demonstrating that dynamic threshold settings can replace manual $\sigma$ tuning to achieve favorable outcomes in real-world applications.
+It is important to note that AUC, V-ROC, and V-PR are threshold-agnostic metrics, so using SPOT does not affect these results.
+Figure 5: Loss analysis of PatchAD.
+TABLE III: Ablation studies on MSL, SMAP, WADI, and PSM datasets.
+Dataset MSL SMAP WADI PSM
+With PA Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1
+wo_ch_sharing 98.78 91.93 96.9 94.35 98.84 93.5 97.75 95.58 98.71 84.99 94.25 89.38 97.69 97.62 93.96 95.75
+wo_ch_mixer 98.78 92.03 96.82 94.36 98.77 93.5 97.1 95.27 98.92 85.4 97.93 91.24 96.77 98.23 89.98 93.92
+wo_mixrep_mixer 96.41 89.54 74.66 81.43 95.25 91.52 68.55 78.7 95.76 67.55 51.99 58.76 97.35 97.87 92.44 95.08
+wo_pos_emb 97.85 91.51 87.73 89.58 96.47 92.18 79.13 85.16 98.23 83.96 85.53 84.74 96.89 97.84 90.77 94.17
+wo_constraint 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 85.26 94.25 89.53 93.95 97.22 80.45 88.04
+PatchAD with SPOT 98.49 89.94 96.46 93.08 98.93 94.2 97.65 95.89 98.53 84.18 91.53 87.7 97.75 97.79 93.99 95.85
+PatchAD (Ours) 98.92 92.05 98.2 95.02 99.1 94.49 99.13 96.75 99.06 85.92 100 92.43 98.89 97.72 98.52 98.11
+Dataset MSL SMAP WADI PSM
+Without PA AUC F1 V-ROC V-PR AUC F1 V-ROC V-PR AUC F1 V-ROC V-PR AUC F1 V-ROC V-PR
+wo_ch_sharing 53.54 20.57 60.07 18.14 59.46 22.71 43.36 11.60 85.24 50.33 88.13 49.22 64.65 48.00 65.93 40.13
+wo_ch_mixer 53.09 20.16 59.69 17.99 59.47 22.68 43.06 11.59 82.41 46.50 86.88 48.00 65.02 48.08 66.59 40.65
+wo_mixrep_mixer 53.50 20.70 60.02 18.13 59.35 22.69 43.41 11.61 81.14 50.61 87.96 47.09 64.47 48.11 65.91 40.09
+wo_pos_emb 52.91 19.55 59.68 17.95 55.95 22.68 46.87 12.15 84.63 51.02 87.36 49.96 64.76 48.04 66.17 40.25
+wo_constraint 53.53 20.79 59.97 18.11 59.93 22.71 42.85 11.52 83.56 51.29 86.78 46.09 64.93 47.95 66.54 40.53
+PatchAD with SPOT 60.94 23.45 66.43 19.65 60.27 23.77 48.30 13.71 85.27 52.75 88.29 49.60 64.74 48.58 65.11 40.72
+PatchAD (Ours) 60.94 23.92 66.43 19.65 60.27 24.74 48.30 13.71 85.27 52.82 88.29 49.60 64.74 49.33 65.11 40.72
+Ablation on Loss Functions: To facilitate the spatial separation of the two distinct views, the design in Eq.
+(14) employs the Kullback-Leibler ($\operatorname{KL}$) divergence as the foundational loss.
+To further analyze the impact of various types of loss functions on the final model performance, we selected the following three common and widely utilized distance or distributional metric functions as alternatives, enabling an examination of their effects on the ultimate outcomes.
+  1. 1.
+    L2 Distance (L2).
+  2. 2.
+    Wasserstein Distance (Wasserstein).
+  3. 3.
+    Jensen–Shannon Divergence (JSD).
+Table IV illustrates the performance of the PatchAD model after substituting the three loss functions in comparison to the model utilizing KL divergence (Ours).
+In the comparison of the first set of metrics employing PA, the models based on L2 and JSD exhibit average performance that falls short of the final model; however, the overall difference from Ours is minimal.
+Conversely, the Wasserstein distance yields a PA-F1 score of only 83.93 on the WADI dataset, indicating a performance that is significantly inferior to that of the other distance functions.
+This may be attributed to the fact that the Wasserstein distance contributes to anomaly scores in a more smooth manner, which, in turn, results in a reduced recall rate for the model and consequently leads to suboptimal performance.
+In the comparison of the second set of metrics, PatchAD consistently ranks as the best across nearly all datasets and metrics.
+Additionally, models trained with the other three foundational distances exhibit markedly lower performance in AUC, V-ROC, and V-PR compared to PatchAD.
+This suggests that these models do not possess sufficiently clear decision boundaries for anomalies, thereby complicating the model’s ability to accurately differentiate between anomalies and normal instances.
+Overall, we posit that JSD represents the best alternative to KL divergence, primarily because the construction of JSD is based on KL divergence, thereby retaining KL divergence’s property of optimizing over probability distributions.
+However, JSD is a symmetric loss function, while KL divergence is asymmetric; this discrepancy limits the ability to achieve asynchronous optimization of the parameters in different modules (the Inter Mixer and Intra Mixer).
+Ultimately, this results in JSD performing poorly in enhancing the distinction between normal and abnormal instances, a phenomenon that aligns with previous work [17].
+TABLE IV: The Influence of different loss functions as fundamental distances on PatchAD.
+Dataset MSL SMAP WADI PSM
+With PA Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1 Acc Pre Rec PA-F1
+L2 98.12 88.92 93.82 91.30 98.90 93.75 97.94 95.80 99.03 86.89 97.93 92.08 97.12 97.65 91.84 94.65
+Wasserstein 98.16 88.85 94.41 91.55 99.20 95.33 98.57 96.92 98.07 80.35 87.85 83.93 97.59 97.45 93.75 95.56
+JSD 98.32 90.79 93.51 92.13 99.16 97.81 95.55 96.66 99.01 85.37 98.57 91.50 98.41 97.88 96.33 97.10
+PatchAD (Ours) 98.92 92.05 98.20 95.02 99.10 94.49 99.13 96.75 99.06 85.92 100.00 92.43 98.89 97.72 98.52 98.11
+Dataset MSL SMAP WADI PSM
+Without PA AUC F1-cls V-ROC V-PR AUC F1-cls V-ROC V-PR AUC F1-cls V-ROC V-PR AUC F1 V-ROC V-PR
+L2 53.52 20.10 60.24 18.15 58.69 22.68 44.01 11.67 85.54 50.22 87.44 47.71 64.64 48.17 65.17 40.12
+Wasserstein 53.39 19.62 60.13 18.08 59.56 22.69 43.24 11.57 84.15 51.61 86.58 46.13 64.66 48.20 65.71 40.56
+JSD 53.64 19.57 60.34 18.13 58.78 22.68 44.02 11.67 85.08 52.62 87.92 49.24 65.66 49.11 65.69 40.36
+PatchAD (Ours) 60.94 23.92 66.43 19.65 60.27 24.74 48.30 13.71 85.27 52.82 88.29 49.60 64.74 49.33 65.11 40.72
+IV-C2 Visual Analysis
 We visualized PatchAD’s response to various anomalies in Figure 4.
@@
 Figure 6: The case studies on SWAT and WADI datasets.
-4.4.4 Time Cost
-We conducted a time consumption analysis on the SMD dataset under identical experimental conditions.
-Figure 5(b) demonstrates that under different sliding step conditions, our method outperforms DCdetector.
-Specifically, when the step is set to 1, PatchAD’s speed is 1.45x faster than DCdetector.
-Additionally, Figure 5(d) illustrates the time required for model training with a fixed window size (window size=105).
-This indicates that PatchAD exhibits higher efficiency in training compared to DCdetector.
-Figure 5(c) shows the time required for model training with a fixed sliding step (step=35) while varying the window size from 35 to 175 for different $d_{model}$ sizes.
-Both graphs show that the time consumption of PatchAD exhibits a linear growth with respect to the independent variables.
-Thus, increasing the window size or enlarging $d_{model}$ has a relatively minor impact on PatchAD.
-However, the time step noticeably affects the duration as this parameter increases data volume.
-4.4.5 Model Size
-Figure 2 compares the sizes of different models on SWAT dataset.
-Our model has a compact size of only 3.2MB, which is smaller than both AnomTrans and DCdetector, and significantly smaller than the LLM model (GPT2-Adapter) and diffusion model (D3R).
-This observation indicates that our model is more lightweight and resource-efficient.
-5 Conclusion
-We presents PatchAD, an innovative algorithm tailored for time series anomaly detection.
-PatchAD developed a lightweight multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning.
-This architecture is unique in using a purely MLP-based approach to discern inter-patch and intra-patch relationships in time series sequences.
-By amplifying the differences between two data views, PatchAD enhance its capability to differentiate between normal and anomalous patterns.
-Furthermore, PatchAD adeptly learns the relationships between different data channels and incorporates a Dual Project Constraint to avert model degradation.
-Moreover, the efficacy of PatchAD, including its various components, is validated through extensive experimentation.
-The results demonstrate that proposed lightweight PatchAD outperforms over 30 methods across nine benchmark datasets from different application scenarios.
-Appendix
-This is the appendix of PatchAD: A Lightweight Patch-Based MLP-Mixer for Time Series Anomaly Detection.
-Appendix A Theory
-In this section, our objective is to perform a theoretical analysis of PatchAD’s effectiveness.
-We want to prove that normal time series data points exhibit reduced information entropy.
-We have made the following assumptions based on the theoretical analysis presented in previous studies.
-Assumption 1: The distribution of normal data differs from that of anomalies, and both consist of two components, denoted as $\tau+\alpha$.
-Here, $\tau$ represents the underlying data pattern, while $\alpha$ represents the pattern of normal or anomalous data.
-Therefore, we denote normal data distribution by $\Pi$ and anomalies as $\Omega$.
-As a result, the normal data $\mathcal{X}_{norm}$ and the anomalous data $\mathcal{X}_{anom}$ can be expressed as follows:
-$\mathcal{X}_{norm}=\tau+\pi,\text{ and }\mathcal{X}_{anom}=\tau+\omega.$ (12)
-Assumption 2: The number of normal data samples in the dataset is significantly larger than that of anomalous data samples. Therefore, we have the following relationship:
-$|\mathcal{X}_{norm}|=\zeta\cdot|\mathcal{X}_{anom}|>|\mathcal{X}_{anom}|,$ (13)
-where ||$\cdot$|| ⋅ | represents the number of samples in the data, and $\zeta$ is a large number. We denote the entire dataset as $\mathcal{D}=\{\mathcal{X}_{norm},\mathcal{X}_{anom}\}$.
-Proof 1: The entropy of the anomalous data is higher than that of normal data.
-For a random variable $X$, its entropy is defined as, and the entropy serves as a metric for quantifying the uncertainty and complexity of the data [54]:
-$\operatorname{H}(X)=-\sum_{i}p(x_{i})\log_{2}p(x_{i})$ (14)
+IV-C3 Scalability Analysis
 […]
-Using the addition rule of probability, we have $p(\tau+\pi)=p(\tau)\cdot p(\pi)$.
-Substituting this into the previous equation, we obtain: $\operatorname{H}(\mathcal{X}_{norm})=-\int p(\tau)\cdot p(\pi)\log(p(\tau)\cdot p(\pi))d(\tau+\pi).$ Expanding further, we can deduce:
-$\operatorname{H}(\mathcal{X}_{norm})=-\int p(\tau)\cdot p(\pi)(\log p(\tau)+\log p(\pi))d(\tau+\pi).$ (15)
-[…]
-Due to the significantly larger number of samples in normal data compared to anomalous data, i.e., $|\mathcal{X}_{norm}|=\zeta\cdot|\mathcal{X}_{anom}|>|\mathcal{X}_{anom}|$, we can extract the large number $\zeta$ and use the properties of logarithm to move it inside the $\log$ function:
-$\operatorname{H}(\mathcal{X}_{norm})=-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log P^{\prime}(x)-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log\zeta$ (16)
-The second term $-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log\zeta$ is a constant term, denoted as $C$.
-Therefore, we can express the entropy of normal data as:
-$\operatorname{H}(\mathcal{X}_{norm})=-\zeta\sum_{x\in\mathcal{X}_{norm}}P^{\prime}(x)\log P^{\prime}(x)+C.$ (17)
-Similarly, we can derive the entropy of anomalous data $\operatorname{H}(\mathcal{X}_{anom})$:
-$\operatorname{H}(\mathcal{X}_{anom})=-\sum_{x\in\mathcal{X}_{anom}}P^{\prime}(x)\log P^{\prime}(x).$ (18)
-At this point, we can conclude that $\operatorname{H}(\pi)<\operatorname{H}(\omega)$.
-Therefore, the entropy of normal data $\operatorname{H}(\mathcal{X}_{norm})$ is smaller than the entropy of anomalous data $\operatorname{H}(\mathcal{X}_{anom})$.
-Appendix B Experimental Settings
-B.1 Datasets
-We applied eight multivariable and one univariable datasets for evaluation: (1) MSL (2) SMAP  [^2] MSL & SMAP: https://github.com/ML4ITS/mtad-gat-pytorch(3) PSM (4) SMD  [^3] https://github.com/NetManAIOps/OmniAnomaly (5) SWaT  [^4] https://itrust.sutd.edu.sg/itrust-labs_datasets (6) WADI  [^5] https://itrust.sutd.edu.sg/testbeds/water-distribution-wadi (7) NIPS-TS-SWAN (8) NIPSTS-GECCO  [^6] Others:http://github.com/DAMO-DI-ML/KDD2023-DCdetector.
-(9) UCR  [^7] https://www.cs.ucr.edu/~eamonn/time_series_data_2018/UCR_TimeSeriesAnomalyDatasets2021.zip Table 4 shows the details.
-Table 4: Details of benchmark datasets.
-Dataset #Training #Test (Labeled) Dimension Anomaly ratio (%)
-MSL 58317 73729 55 10.5
-SMAP 135183 427617 25 12.8
-PSM 132481 87841 25 27.8
-SMD 708377 708393 38 4.2
-SWaT 99000 89984 26 12.2
-WADI 1048571 172801 123 5.99
-NIPS-TS-SWAN 60000 60000 38 32.6
-NIPS-TS-GECCO 69260 69261 9 1.1
-B.2 Baslines
-We compared our model against 35 benchmark models, including: (1) Reconstruction-based: AutoEncoder [55], LSTM-VAE [6], OmniAnomaly [7], BeatGAN [43], InterFusion [47], USAD [17], COUTA[24]; (2) Autoregression-based: VAR [12], Autoregression [56], LSTM-RNN [13], LSTM [46], CL-MPPCA [41], TimesNet [14], NPSR-pt and NPSR-seq [50]; (3) Density-based: LOF [8], MMPCACD [40], DAGMM [9]; (4) Clustering-based: DeepSVDD [45], THOC [48], ITAD [39]; Matrix Profile [57]; (5) The classic methods: OCSVM [19], OCSVM-based subsequence clustering (OCSVM*), IForest [37], IForest-based subsequence clustering (IForest*), Gradient boosting regression (GBRT) [58]; (6) Change point detection and time series segmentation methods: BOCPD [44], U-Time [38], TS-CP2 [42]; (7) Transformer-based: Anomaly Transformer  [^8] https://github.com/thuml/Anomaly-Transformer [27], DCdetector  [^9] https://github.com/DAMO-DI-ML/KDD2023-DCdetector [10], TranAD [59]; (8) Large Language Model: GPT2-Adapter[15]; (9) Diffusion model: D3R [60]; These two are current state-of-the-art methods.
-B.3 Implementation Details
-All experiments used PyTorch implementation on a single NVIDIA GTX 1080TI GPU.
-We summarize the default hyperparameters of PatchAD as follows: PatchAD consists of 3 encoder layers ($L$=3).
-The hidden layer dimension $d_{{model}}$ is set to 40.
-The default projector constraint coefficient is $0.2$.
-The Adam optimizer’s default learning rate is $10^{-4}$, with a batch size of 128, trained for 3 epochs across all datasets.
-Different patch sizes and window sizes were chosen for each dataset, defaulting to [3, 5] and 105, respectively.
-If an anomaly score of any given timestamp exceeds a threshold $\sigma$, the model identifies it as an anomaly.
-The default value for $\sigma$ is 1.
-Appendix C Algorithm’s Pseudo-code
-Algorithm 1 PatchAD’s Python-like code
-PE = PositionEmbedding(d_model) # Postional Embedding
-VE_n, VE_p = Linear(dim1, d_model), Linear(dim2, d_model)
-# Value Embedding of N and P views
-# Patching and Embedding
-x = PE(x)
-x_n = rearange(x, ’b (n p) c -> b c n p’, p=patch_size)
-x_p = rearange(x, ’b (p n) c -> b c p n’, p=patch_size)
-x_n = VE_n(x_n)
-x_p = VE_p(x_p)
-# x_n’s shape (B, c_dim, n_dim, d_model)
-# x_p’s shape (B, c_dim, p_dim, d_model)
-Channel_Mixer = MLP(c_dim, transpose_dim=1)
-Inter_Mixer = MLP(n_dim, transpose_dim=2)
-Intra_Mixer = MLP(p_dim, transpose_dim=2)
-MixRep_Mixer = MLP(d_model, transpose_dim=3)
-# Patch Mixer Encoding
-x_inter = Channel_Mixer(x_n)
-x_intra = Channel_Mixer(x_p)
-x_inter = Inter_Mixer(x_inter)
-x_intra = Intra_Mixer(x_intra)
-x_inter = MixRep_Mixer(x_inter)
-x_intra = MixRep_Mixer(x_intra)
-# Dual Project
-x_inter = x_inter.mean(1)
-x_intra = x_intra.mean(1)
-Inter_Proj = MLP(d_model)
-Intra_Proj = MLP(d_model)
-x_inter2 = Inter_Proj(x_inter)
-x_intra2 = Intra_Proj(x_intra)
-# Ouput shape (B, n_dim, d_model) and (B, p_dim, d_model)
-# Upsample to (B, L, d_model)
-rec1 = Rec1(x_inter)
-rec2 = Rec2(x_intra)
-rec_x = rec1 + rec2
-# Update
-def loss_fn(a, b):
-return KL(a, b.detach()) + KL(b.detach(), a)
-def inter_intra_loss_fn(n, p):
-return loss_fn(n, p) - loss_fn(p, n)
-loss1 = inter_intra_loss_fn(x_inter, x_intra)
-loss_n_proj = inter_intra_loss_fn(x_inter, x_intra2)
-loss_p_proj = inter_intra_loss_fn(x_inter2, x_intra)
-loss_proj = loss_n_proj + loss_p_proj
-loss_rec = MSE(rec_x, x)
-final_loss = (1-c) * loss1 + c * loss_proj + loss_rec
-# backward and optimize
-Appendix D Unveiling the Mechanisms
+V Mechanism and Theory
+This section endeavors to analyze the mechanisms of PatchAD from both experimental and theoretical perspectives, further elucidating the principles underlying its effectiveness.
+It aims to clarify the design motivations behind PatchAD.
+V-A Unveiling the Mechanism
 Figure 7 illustrates the knowledge acquired by PatchAD in four distinct cases.
-Cases 2, 3, 4 and 5 have a patch size of 1, whereas case 1 has a patch size of $[1,2,5]$.
-The "Original" represents the original data, while "Rec 1" and "Rec 2" correspond to the features learned by the Inter Mixer and the Intra Mixer, respectively.
+Cases 2, 3, 4 and 5 have a patch size of 1, whereas case 1 has a patch size of [1,2,5][1,2,5][ 1 , 2 , 5 ].
+The Original represents the original data, while Rec 1 and Rec 2 correspond to the features learned by the Inter Mixer and the Intra Mixer, respectively.
 It is observed that the Intra Mixer captures the generalized knowledge in the data, which encompasses the overall trend information.
@@
 Case 1 was chosen as the baseline, and we investigated the entropy of the features acquired by the Inter and Intra Mixers under varying patch size conditions.
-Figures 9(a) and 9(b) illustrate that both the Inter Mixer and Intra Mixer display a consistent trend of decreasing entropy throughout the learning process.
+Figures 8(a) and 8(b) illustrate that both the Inter Mixer and Intra Mixer display a consistent trend of decreasing entropy throughout the learning process.
 This observation suggests that the models effectively acquire knowledge and reduce uncertainty.
@@
 Figure 9: Latent distribution’s visualization by T-SNE.
-Appendix E Parameter Sensitivity
-We conducted a sensitivity analysis on PatchAD’s parameters.
-The default parameter settings remain consistent with the basic configuration.
-Figure 6(a) demonstrates the impact of the projection head constraint on PatchAD.
-Overall, it holds more significance for smaller-scale datasets like MSL.
-However, excessively strong constraints hinder the model’s learning process.
-Therefore, setting this constraint between 0.1 to 0.4 is more advisable.
-Figure 6(b) analyzes PatchAD’s final model performance when the anomaly threshold $\sigma$ ranges from 0.5 to 1.2.
-The selection of the anomaly threshold appears more robust on the PSM and SMAP datasets than the MSL dataset.
-Setting this value between 0.8 and 1 for most datasets yields optimal results.
-Figure 6(c) illustrates how the model’s performance is influenced by the network dimension $d_{model}$.
-PatchAD demonstrates better results with smaller $d_{model}$ sizes.
-Therefore, to balance model performance and complexity, we set $d_{model}=40$ to ensure PatchAD operates effectively.
-Figure 6(d) indicates that PatchAD’s performance is influenced by the number of encoder layers, and given a compromise, we set it to 3 or 4.
-Figure 6(e) demonstrates PatchAD’s robustness concerning the window size, which ranges from 60 to 210.
-The time window is a critical parameter, typically set to an initial value of 105 across most datasets.
-Lastly, Figure 6(f) displays PatchAD’s performance under different multi-scale patch size combinations.
-(e.g., a patch size of [3,5] represents using two scales in PatchAD.) The outcomes suggest that combining various patch sizes can enhance PatchAD’s performance.
-Appendix F Visualization comparison
-Figure 11 shows the performance of WindowAD algorithms in the face of different types of anomalies.
-These methods have difficulty in detecting contextual and group point anomalies and are able to detect seasonal anomalies, but there are some false positives, which indicates that it has a low precision.
-(a) Contextual
-(b) Seasonal
-(c) Group
-Figure 11: Visualization of WindowAD-based algorithm’s score.
-Appendix G Full Comparison Results
-Table 6 presents the performance of different methods on four datasets, namely WADI, SWAT, PSM, and SWAN.
-This is the original data of Figure2.
-We have selected representative algorithms for experimental comparison.
-Table 5: Overall results on NIPS-TS datasets. All results are in %, the best in Bold, and the second in underlined.
-Dataset NIPS-TS-SWAN NIPS-TS-GECCO Average F1
-Metric P R F1 P R F1
-MatrixProfile [57] 16.70 17.50 17.10 4.60 18.50 7.40 12.25
-GBRT [58] 44.70 37.50 40.80 17.50 14.00 15.60 28.20
-LSTM-RNN [13] 52.70 22.10 31.20 34.30 27.50 30.50 30.85
-Autoregression [56] 42.10 35.40 38.50 39.20 31.40 34.90 36.70
-OCSVM [19] 47.40 49.80 48.50 18.50 74.30 29.60 39.05
-IForest* [37] 40.60 42.50 41.60 39.20 31.50 39.00 40.30
-AutoEncoder [55] 49.70 52.20 50.90 42.40 34.00 37.70 44.30
-AnomalyTrans [27] 90.70 47.40 62.30 25.70 28.50 27.00 44.65
-IForest [37] 56.90 59.80 58.30 43.90 35.30 39.10 48.70
-AnomalyTrans [27] 90.71 47.43 62.29 29.96 48.63 37.08 49.69
-DCdetector [10] 96.57 59.08 73.31 38.41 59.73 46.76 60.04
-PatchAD (Ours) 96.75 59.15 73.41 40.62 62.88 49.35 61.38
-Table 6: The performance of algorithms on different evaluation metrics without point adjustment.
-Dataset WADI SWAT
-Metric P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC
-Deep SVDD 4.05 38.53 7.34 50.06 55.79 52.77 40.53 12.74 98.50 22.57 53.58 99.99 69.77 86.81
-USAD 5.74 100.00 10.86 54.92 100.00 70.90 46.39 12.17 100.00 21.70 53.00 100.00 69.28 88.66
-COUTA 98.72 15.58 26.92 99.18 23.08 37.44 49.48 94.25 6.33 11.87 79.98 33.20 46.92 75.54
-TranAD 6.49 63.89 11.78 55.91 91.96 69.54 52.60 14.70 96.40 25.50 55.65 99.66 71.42 88.90
-TimesNet 64.04 10.24 17.65 65.25 39.30 49.05 65.08 11.05 3.21 4.98 57.84 93.77 71.55 26.76
-AnomTrans 4.14 14.12 6.40 45.73 40.66 43.05 49.20 4.81 29.32 8.27 53.03 98.08 36.08 19.20
-DCdetector 5.80 14.42 8.27 59.63 90.98 69.98 49.88 15.49 5.54 8.16 52.76 97.33 68.94 50.68
-GPT2-Adapter 5.68 5.50 5.59 55.62 98.29 71.04 50.01 12.16 0.96 1.78 52.51 98.13 68.41 50.00
-D3R 6.99 80.99 12.86 56.47 99.98 72.17 51.39 32.87 75.97 45.89 61.47 78.52 68.96 79.95
-PatchAD (Ours) 89.48(2.08) 31.22(0.74) 46.28(1.09) 73.07(0.41) 86.07(0.28) 79.04(0.12) 79.89(0.55) 85.42(0.98) 70.34(0.18) 77.15(0.48) 73.62(0.52) 84.65(0.22) 78.75(0.20) 91.19(0.12)
-Dataset PSM SWAN
-Metric P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC P R F1_ori Aff-Pre AFF-Rec Aff-F1 AUC
-Deep SVDD 28.59 99.91 44.46 53.92 99.98 70.06 78.09 41.74 85.03 55.99 56.84 89.93 69.66 75.32
-USAD 32.36 92.17 47.90 55.30 98.71 70.88 64.53 39.73 77.94 52.63 55.26 68.62 61.22 66.76
-COUTA 100.00 0.04 0.08 100.00 5.12 9.75 72.45 62.96 0.09 0.17 69.25 0.10 0.19 71.38
-TranAD 35.90 54.23 43.20 60.88 92.09 73.30 65.22 38.84 78.29 51.93 57.86 90.18 70.49 70.00
-TimesNet 64.85 1.27 2.49 77.84 67.23 72.15 58.74 38.63 48.89 43.16 60.75 81.29 69.53 53.62
-AnomTrans 26.05 84.32 39.81 55.35 80.28 50.09 52.18 31.19 79.87 44.86 58.45 9.49 0.54 45.38
-DCdetector 27.05 19.59 22.72 52.33 76.95 65.88 49.62 32.45 98.62 48.83 38.73 3.13 7.50 50.35
-GPT2-Adapter 28.31 46.64 35.24 54.33 95.16 69.17 50.74 40.00 52.05 45.23 61.17 69.55 65.09 58.77
-D3R 28.96 93.65 44.24 53.84 100.00 69.99 60.64 36.12 78.38 49.45 57.70 96.06 72.10 62.54
-PatchAD (Ours) 47.13(0.44) 49.32(0.97) 48.20(0.64) 63.52(0.82) 96.84(1.21) 76.72(0.19) 64.74(0.01) 51.76(0.18) 65.31(0.47) 57.75(0.08) 66.48(0.80) 79.39(1.54) 72.36(0.20) 72.91(0.02)
-In Table 7, Aff-P and Aff-R are the precision and recall of affiliation metric [51], respectively.
-R_A_R and R_A_P are Range-AUC-ROC and Range-AUC-PR [52], respectively.
-V_ROC and V_RR are volumes under the surfaces created based on ROC curve and PR curve [52], respectively.
-Besides, we give more results in table 8, and the results from more real-world multivariate time-series datasets show that our proposed PatchAD has good robustness and performance.
-Table 7: Multi-metrics results on all real-world multivariate datasets. All results are in %, the best in Bold, and the second in underlined.
-Dataset Method Acc F1 Aff-P Aff-R R_A_R R_A_P V_ROC V_PR
-MSL AnomalyTrans 98.69 93.93 51.76 95.98 90.04 87.87 88.20 86.26
-DCdetector 98.83 94.77 51.01 96.74 91.29 89.12 89.80 87.81
-PatchAD 98.92 95.02 52.06 96.44 90.89 88.66 90.09 87.99
-SMAP AnomalyTrans 99.05 96.41 51.39 98.68 96.32 94.07 95.52 93.37
-DCdetector 99.06 96.38 50.29 98.04 93.76 92.40 92.70 91.47
-PatchAD 99.15 96.75 52.30 98.70 96.05 94.23 95.12 93.42
-PSM AnomalyTrans 98.68 97.37 55.35 80.28 91.83 93.03 88.71 90.71
-DCdetector 98.81 97.63 52.33 76.95 89.91 91.74 85.57 88.52
-PatchAD 98.95 98.11 55.58 82.72 94.85 95.65 89.26 91.67
-SMD AnomalyTrans 99.11 90.06 72.63 98.27 82.12 75.90 81.86 75.69
-DCdetector 98.78 86.02 51.61 93.44 76.13 69.85 73.60 67.45
-PatchAD 98.84 86.62 50.81 93.32 77.55 71.16 74.09 67.88
-SWaT AnomalyTrans 98.62 94.62 53.03 98.08 97.89 93.47 97.92 93.49
-DCdetector 93.25 96.51 52.76 97.33 98.23 95.54 98.25 95.56
-PatchAD 99.12 96.52 52.78 98.17 98.23 95.55 98.26 95.57
-WADI AnomalyTrans 98.17 85.47 87.84 99.75 87.20 78.50 87.36 78.63
-DCdetector 98.70 89.26 59.63 90.98 87.56 81.62 87.16 81.23
-PatchAD 99.06 92.43 60.37 91.76 89.85 84.18 90.68 84.99
-NIPS-TS-SWAN AnomalyTrans 84.57 62.29 58.45 9.49 86.42 93.26 84.81 92.00
-DCdetector 85.99 73.31 38.73 3.13 88.02 94.78 86.02 93.49
-PatchAD 86.05 73.41 51.44 4.34 88.03 94.80 86.13 93.57
-NIPS-TS-GECCO AnomalyTrans 98.26 37.08 55.65 89.12 60.74 28.17 60.48 28.02
-DCdetector 98.57 46.76 49.66 87.68 61.01 32.13 60.77 31.90
-PatchAD 98.64 49.35 53.04 89.48 63.59 35.92 63.09 35.43
-Table 8: More results on all real-world multivariate datasets. All results are in %, the best in Bold, and the second in underlined.
-Dataset MSL SMAP PSM SMD SWaT Average F1
-Metric P R F1 P R F1 P R F1 P R F1 P R F1
-LOF 47.72 85.25 61.18 58.93 56.33 57.60 57.89 90.49 70.61 56.34 39.86 46.68 72.15 65.43 68.62 60.94
-OCSVM 59.78 86.87 70.82 53.85 59.07 56.34 62.75 80.89 70.67 44.34 76.72 56.19 45.39 49.22 47.23 60.25
-U-Time 57.20 71.66 63.62 49.71 56.18 52.75 82.85 79.34 81.06 65.95 74.75 70.07 46.20 87.94 60.58 65.62
-IForest 53.94 86.54 66.45 52.39 59.07 55.53 76.09 92.45 83.48 42.31 73.29 53.64 49.29 44.95 47.02 61.22
-DAGMM 89.60 63.93 74.62 86.45 56.73 68.51 93.49 70.03 80.08 67.30 49.89 57.30 89.92 57.84 70.40 70.18
-ITAD 69.44 84.09 76.07 82.42 66.89 73.85 72.80 64.02 68.13 86.22 73.71 79.48 63.13 52.08 57.08 70.92
-VAR 74.68 81.42 77.90 81.38 53.88 64.83 90.71 83.82 87.13 78.35 70.26 74.08 81.59 60.29 69.34 74.66
-MMPCACD 81.42 61.31 69.95 88.61 75.84 81.73 76.26 78.35 77.29 71.20 79.28 75.02 82.52 68.29 74.73 75.74
-CL-MPPCA 73.71 88.54 80.44 86.13 63.16 72.88 56.02 99.93 71.80 82.36 76.07 79.09 76.78 81.50 79.07 76.66
-TS-CP2 86.45 68.48 76.42 87.65 83.18 85.36 82.67 78.16 80.35 87.42 66.25 75.38 81.23 74.10 77.50 79.00
-Deep-SVDD 91.92 76.63 83.58 89.93 56.02 69.04 95.41 86.49 90.73 78.54 79.67 79.10 80.42 84.45 82.39 80.97
-BOCPD 80.32 87.20 83.62 84.65 85.85 85.24 80.22 75.33 77.70 70.90 82.04 76.07 89.46 70.75 79.01 80.33
-LSTM-VAE 85.49 79.94 82.62 92.20 67.75 78.10 73.62 89.92 80.96 75.76 90.08 82.30 76.00 89.50 82.20 81.24
-BeatGAN 89.75 85.42 87.53 92.38 55.85 69.61 90.30 93.84 92.04 72.90 84.09 78.10 64.01 87.46 73.92 80.24
-LSTM 85.45 82.50 83.95 89.41 78.13 83.39 76.93 89.64 82.80 78.55 85.28 81.78 86.15 83.27 84.69 83.32
-OmniAnomaly 89.02 86.37 87.67 92.49 81.99 86.92 88.39 74.46 80.83 83.68 86.82 85.22 81.42 84.30 82.83 84.69
-InterFusion 81.28 92.70 86.62 89.77 88.52 89.14 83.61 83.45 83.52 87.02 85.43 86.22 80.59 85.58 83.01 85.70
-THOC 88.45 90.97 89.69 92.06 89.34 90.68 88.14 90.99 89.54 79.76 90.95 84.99 83.94 86.36 85.13 88.01
-AnomalyTrans 91.92 96.03 93.93 93.59 99.41 96.41 96.94 97.81 97.37 84.14 96.88 90.06 89.79 100.0 94.62 94.48
-DCdetector 92.22 97.48 94.77 94.43 98.41 96.38 97.19 98.08 97.63 83.04 89.22 86.02 93.25 100.0 96.51 94.26
-PatchAD 92.05 98.20 95.02 94.49 99.13 96.75 97.72 98.52 98.11 83.06 90.51 86.62 93.28 100.0 96.52 94.60
-Table 9: Parameter studies on project head constraint results (patch size=[3,5], window size=105). All results are in %.
-Dataset MSL SMAP PSM
-Contraint Acc P R F1 Acc P R F1 Acc P R F1
-0.0 98.51 91.67 94.49 93.06 99.01 93.55 99.05 96.22 98.73 97.42 98.03 97.72
-0.1 98.92 91.96 98.33 95.03 99.04 94.43 98.33 96.34 98.72 97.35 98.04 97.70
-0.2 98.92 92.05 98.20 95.02 99.11 94.42 98.91 96.61 98.91 97.43 98.66 98.04
-0.3 98.89 92.03 97.93 94.88 99.14 94.41 99.11 96.71 98.89 97.42 98.59 98.00
-0.4 98.86 92.09 97.54 94.73 99.12 94.45 98.97 96.66 98.71 97.42 97.94 97.68
-0.5 98.96 92.12 98.60 95.25 99.02 94.40 98.17 96.25 98.73 97.42 98.03 97.72
-0.6 98.92 92.11 98.20 95.06 99.01 94.39 98.07 96.19 98.92 97.45 98.70 98.07
-0.7 98.84 91.91 97.59 94.67 99.07 94.43 98.55 96.44 98.76 97.39 98.14 97.76
-0.8 98.91 92.01 98.20 95.00 99.10 94.40 98.81 96.55 98.92 97.42 98.73 98.07
-0.9 98.78 91.99 96.90 94.38 98.96 94.37 97.70 96.01 98.92 97.47 98.64 98.05
-Table 10: Parameter studies on anomaly threshold $\sigma$ results (patch size=[3,5], window size=105). All results are in %.
-Dataset MSL SMAP PSM
-AR Acc P R F1 Acc P R F1 Acc P R F1
-0.5 97.93 95.39 84.39 89.55 99.15 96.46 96.91 96.68 98.45 98.61 95.77 97.17
-0.6 98.27 94.70 88.54 91.52 99.15 95.84 97.58 96.70 98.77 98.33 97.20 97.76
-0.7 98.17 93.76 88.54 91.07 99.09 95.22 97.77 96.48 98.70 98.07 97.20 97.64
-0.8 98.45 93.19 92.04 92.61 99.15 94.73 98.86 96.75 98.95 97.94 98.29 98.11
-0.9 98.45 92.36 92.93 92.64 99.01 94.07 98.49 96.23 98.92 97.68 98.45 98.06
-1.0 98.84 91.80 97.73 94.67 99.01 94.05 98.49 96.22 98.70 97.30 98.03 97.66
-1.1 98.73 91.22 97.36 94.19 98.85 92.76 98.69 95.63 98.68 97.06 98.20 97.63
-1.2 98.64 90.45 97.36 93.78 98.76 92.19 98.69 95.33 98.62 96.85 98.20 97.52
-Table 11: Parameter studies on $d_{model}$ size results (patch size=[3,5], window size=105). All results are in %.
-Dataset MSL SMAP PSM
-$d_{model}$ Acc P R F1 Acc P R F1 Acc P R F1
-10 98.62 91.81 95.42 93.58 99.04 94.42 98.28 96.31 98.87 97.38 98.56 97.97
-20 98.92 92.08 98.20 95.04 98.98 94.33 97.88 96.07 98.73 97.37 98.05 97.71
-30 98.89 92.06 97.94 94.91 99.10 94.43 98.80 96.57 98.87 97.43 98.52 97.97
-40 98.91 92.03 98.20 95.02 99.13 94.49 98.95 96.67 98.73 97.42 98.03 97.72
-50 98.93 92.15 98.20 95.08 99.11 94.45 98.89 96.62 98.90 97.44 98.63 98.03
-60 98.91 92.03 98.20 95.02 99.13 94.44 98.98 96.66 98.94 97.48 98.74 98.10
-70 98.87 92.08 97.67 94.79 99.05 94.40 98.43 96.37 98.97 97.48 98.83 98.15
-80 98.77 91.99 96.70 94.29 99.03 94.38 98.24 96.27 98.89 97.44 98.58 98.00
-90 98.92 92.06 98.20 95.03 99.05 94.38 98.42 96.35 98.89 97.46 98.56 98.01
-Table 12: Parameter studies on encoder layers $L$ results (patch size=[3,5], window size=105). All results are in %.
-Dataset MSL SMAP PSM
-Encoder layer Acc P R F1 Acc P R F1 Acc P R F1
-1 98.60 91.72 95.27 93.46 98.76 94.29 96.11 95.19 98.10 97.28 95.82 96.54
-2 98.88 91.96 97.93 94.85 99.01 94.42 98.09 96.22 98.79 97.44 98.20 97.82
-3 98.92 92.07 98.20 95.03 99.10 94.41 98.79 96.55 98.84 97.47 98.36 97.91
-4 98.88 92.19 97.67 94.85 99.11 94.42 98.91 96.61 98.95 97.41 98.83 98.12
-5 98.83 91.95 97.41 94.60 99.13 94.46 99.02 96.69 98.87 97.44 98.52 97.98
-Table 13: Parameter studies on window size results (window size=105). All results are in %.
-Dataset MSL SMAP PSM
-Window size Acc P R F1 Acc P R F1 Acc P R F1
-30 98.11 91.47 90.52 90.99 98.91 94.38 97.23 95.78 98.88 97.48 98.51 97.99
-45 97.98 91.23 89.44 90.33 98.93 95.55 96.14 95.84 98.67 97.30 97.92 97.61
-60 98.33 92.05 92.16 92.10 98.86 94.18 97.09 95.62 98.75 97.59 97.93 97.76
-75 98.51 91.79 94.31 93.03 98.93 94.27 97.60 95.90 98.72 97.35 98.06 97.70
-90 98.71 91.92 96.24 94.03 99.06 94.34 98.55 96.40 98.92 97.44 98.71 98.07
-105 98.92 92.05 98.20 95.02 99.08 94.44 98.60 96.48 98.66 97.38 97.79 97.59
-120 98.86 92.10 97.55 94.75 99.13 94.47 98.98 96.67 98.79 97.53 98.13 97.83
-135 98.65 91.85 95.67 93.72 99.07 94.46 98.50 96.44 98.88 97.38 98.60 97.99
-150 98.71 91.83 96.30 94.02 99.12 94.48 98.91 96.64 98.66 97.43 97.72 97.58
-165 98.61 91.85 95.27 93.53 99.09 94.50 98.65 96.53 98.73 97.44 97.98 97.71
-180 98.53 91.73 94.55 93.12 99.01 94.35 98.18 96.23 98.70 97.55 97.78 97.66
-195 98.77 92.02 96.70 94.31 98.96 94.35 97.74 96.01 98.87 97.44 98.50 97.97
-210 98.81 91.97 97.17 94.50 98.83 94.30 96.71 95.49 98.92 97.45 98.69 98.06
-Table 14: Parameter studies on multi-scale patch size results (window size=105). All results are in %.
-Dataset MSL SMAP PSM
-Patch size Acc P R F1 Acc P R F1 Acc P R F1
-[1] 98.49 91.73 94.12 92.91 99.09 94.47 98.69 96.53 98.66 97.21 97.99 97.60
-[3] 98.57 91.67 95.08 93.34 99.08 94.44 98.61 96.48 98.84 97.56 98.26 97.91
-[5] 98.34 91.70 92.66 92.18 99.02 94.40 98.20 96.26 98.83 97.44 98.37 97.91
-[7] 98.45 91.96 93.41 92.68 99.13 94.46 99.04 96.70 98.88 97.43 98.56 97.99
-[1, 3] 98.63 91.87 95.42 93.61 99.11 94.48 98.86 96.62 98.78 97.46 98.16 97.81
-[1, 5] 98.63 91.88 95.42 93.61 99.07 94.43 98.58 96.46 98.72 97.37 98.04 97.70
-[1, 7] 98.63 91.89 95.42 93.62 99.04 94.43 98.32 96.33 98.71 97.39 97.99 97.69
-[3, 5] 98.78 91.99 96.90 94.38 99.07 94.39 98.63 96.46 98.91 97.43 98.66 98.04
-[3, 7] 98.68 91.92 95.93 93.88 99.09 94.40 98.76 96.53 98.87 97.41 98.55 97.97
-[5, 7] 98.61 91.86 95.27 93.53 99.08 94.44 98.63 96.49 98.85 97.40 98.46 97.93
-[1, 3, 5] 98.85 92.01 97.54 94.69 99.12 94.47 98.92 96.64 98.82 97.37 98.39 97.88
-[1, 3, 7] 98.74 92.00 96.46 94.18 99.09 94.40 98.71 96.50 98.79 97.43 98.21 97.82
-[1, 5, 7] 98.92 92.05 98.20 95.02 99.04 94.38 98.36 96.33 98.67 97.36 97.87 97.61
-[3, 5, 7] 98.56 91.76 94.89 93.30 99.08 94.42 98.61 96.46 98.87 97.45 98.52 97.98
-[1, 3, 5, 7] 98.62 91.78 95.42 93.56 99.08 94.43 98.61 96.48 98.92 97.43 98.71 98.06
-Appendix H Limitations
-Although PatchAD achieves optimal results on multiple datasets, some limitations still need to be addressed.
-In order to ensure the optimization of different metrics, subtle adjustments to potential parameters are necessary.
-For instance, selecting an appropriate anomaly threshold is crucial to balance precision and recall.
-Developing an automated threshold adjustment scheme is a potential avenue for future research.
-Time series anomaly detection employs a wide range of evaluation metrics, each with its own emphasis, leading to potential biases in evaluating different algorithms.
-In the future, we aim to propose a unified set of evaluation metrics that consider multiple factors, contributing to the advancement of the community.
-Appendix I Broader Impacts
-Time series anomaly detection has significant academic implications and broader impacts in various domains.
-In finance, it aids in detecting fraudulent activities like credit card fraud and insider trading, safeguarding financial systems and preventing losses.
-In healthcare, it aids in early disease detection and monitoring, enabling timely interventions and improved patient outcomes.
-Overall, time series anomaly detection improves decision-making, risk mitigation, and system performance across multiple domains.
+V-B Theory Analysis
+Considering spatial constraints and readability, further details are elaborated in Appendix -B.
+VI Conclusion
+We presents PatchAD, an innovative algorithm tailored for time series anomaly detection.
+PatchAD developed a lightweight multi-scale, patch-based MLP Mixer architecture grounded in contrastive learning.
+This architecture is unique in using a purely MLP-based approach to discern inter-patch and intra-patch relationships in time series sequences.
+By amplifying the differences between two data views, PatchAD enhance its capability to differentiate between normal and anomalous patterns.
+Furthermore, PatchAD adeptly learns the relationships between different data channels and incorporates a Dual Project Constraint to avert model degradation.
+Moreover, the efficacy of PatchAD, including its various components, is validated through extensive experimentation.
+The results demonstrate that proposed lightweight PatchAD outperforms over 30 methods across eight benchmark datasets from different application scenarios.
+Additionally, we have conducted a comprehensive and extensive analysis of PatchAD from both theoretical and experimental perspectives, confirming the alignment between the model’s structure and its design motivations.